### v1

# quality of service 5G fifth generation

QoS
quality of service
5G
fifth generation
6G
sixth generation
ML
machine learning
MAC
medium access control
O-RAN
Open Radio Access Network
RAN
Radio Access Network
eMBB
Enhanced Mobile Broadband
URLLC
Ultra-Reliable Low Latency Communications
RU
Radio Unit
DU
Distributed Unit
CU
Central Unit
CP
control plane
UP
user plane
MANO
Management and Orchestration
NFV
Network Functions Virtualization
RIC
Radio Access Network Intelligent Controller
near-RT
near-real-time
non-RT
non-real-time
DL
deep learning
DRL
deep reinforcement learning
VAE
variational autoencoder
UE
User Equipment
PRB
Physical Resource Blocks
SINR
Signal to Interference & Noise Ratio
ESA
exhaustive search algorithm
DNN
deep neural network
MDP
Markov Decision Process
DQN
Deep Q-Network
KL
Kullback-Leibler
MSE
mean squared error
MAE
Mean Absolute Error
Semi-Supervised Learning Approach for Efficient Resource Allocation with Network Slicing in O-RAN
Salar ${}^{\dagger}$    Mojdeh Karbalaee ${}^{\dagger}$     Vahid ${}^{\dagger}$       Seyed Pooya ${}^{\dagger}$
$\dagger$ School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
{salar.nouri, mojdeh.karbalaee, vmansouri, p.shariatpanahi}@ut.ac.ir
Abstract
The Open Radio Access Network (O-RAN) technology has emerged as a promising solution for network operators, providing them with an open and favorable environment. Ensuring effective coordination of x-applications (xAPPs) is crucial to enhance flexibility and optimize network performance within the O-RAN. In this paper, we introduce an innovative approach to the resource allocation problem, aiming to coordinate multiple independent xAPPs for network slicing and resource allocation in O-RAN. Our proposed method focuses on maximizing the weighted throughput among user equipments (UE), as well as allocating physical resource blocks (PRBs). We prioritize two service types, namely enhanced Mobile Broadband and Ultra Reliable Low Latency Communication. To achieve this, we have designed two xAPPs: a power control xAPP for each UE and a PRB allocation xAPP. The proposed method consists of a two-part training phase, where the first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission as well as the user association and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model. We evaluate the performance of our proposed method by comparing its results to those obtained from an exhaustive search algorithm, deep Q-network algorithm, and by reporting performance metrics for the regression task. We also evaluate the proposed model’s performance in different scenarios among the service types. The results show that the proposed method is a more efficient and effective solution for network slicing problems compared to state-of-the-art methods.
Index Terms:
Resource allocation, Network slicing, O-RAN, Variational Autoencoder, Contrastive Learning
I INTRODUCTION
Wireless communication has evolved from a mere convenience to an absolute necessity, driven by an insatiable demand for faster and more dependable connectivity between various quality of service (QoS) requirements [1]. The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2]. However, the pursuit of wireless excellence does not conclude with 5G; it merely marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology. Poised on the horizon, 6G holds the promise to once again redefine the boundaries of connectivity.
Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3]. These advancements have the potential to not only revolutionize our modes of communication but also reshape the very fabric of our interconnected world. Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4]. These techniques aim to inject an unparalleled level of intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions.
In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal. As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly evident that the evolution of wireless communication is an unceasing journey. Each generation builds upon the accomplishments of its forerunners, establishing a robust foundation for the next phase of wireless excellence.
Resource allocation involves the meticulous distribution of finite network resources among numerous users, applications, or services [7]. In 5G, the primary objective is to support a wide spectrum of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8]. Each of these services necessitates distinct resource sets encompassing bandwidth, data rate, latency, power consumption, complexity, and reliability. As such, resource allocation within the 5G landscape must exhibit the flexibility to cater to these diverse demands while ensuring the efficient utilization of network resources.
Network slicing is an innovative concept that tailors virtual networks (slices) within a unified physical infrastructure, addressing specific service types and use cases [9]. In the context of 5G, it efficiently allocates resources for eMBB and URLLC [10] services. This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12]. Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a key technology in the relentless pursuit of wireless excellence, bridging the transition from 5G to 6G.
This concept seamlessly aligns with the architecture of the O-RAN. O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7]. Within this context, the implementation of network slicing in O-RAN not only optimizes resource allocation but also fosters multi-vendor interoperability and innovation.
At the core of O-RAN’s architecture are two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control. These components collaborate seamlessly, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5]. The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP). Adding to O-RAN’s adaptability and innovation are xApps (cross-applications), capable of spanning multiple network domains to optimize resource allocation for various services. These network segments include Fronthaul, Backhaul, and Midhaul, facilitating data transmission between components. Overseeing resource management and orchestration tasks is the Management and Orchestration (MANO) layer, ensuring the network’s efficiency. Network Functions Virtualization (NFV) decouples software from hardware, enhancing scalability. Furthermore, the Radio Access Network Intelligent Controller (RIC) plays a pivotal role in intelligent resource control, enabling dynamic allocation to meet diverse service requirements [5]. These components collectively define O-RAN’s transformative architecture, promoting openness, adaptability, and efficient resource utilization in modern wireless networks. It is worth noting that O-RAN comprises two types of RIC: near-real-time (near-RT) RIC, which manages Radio Access Network (RAN) control and optimization, and non-real-time (non-RT) RIC. Within the near-RT RIC, xAPPs facilitate RAN control and optimization, while inside the non-RT RIC, rAPPs support it by providing guidance and employing ML models. Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
Moreover, xAPPs leverage O-RAN network slicing for efficient resource utilization, allowing seamless collaboration across services. However, this introduces network management complexity, as each xAPP must adapt to multiple slice QoS needs, posing a critical coordination challenge. Additionally, cellular operators face the difficulty of establishing consistent connectivity for devices with different QoS demands [15, 16]. Therefore, in this study, we examine the resource allocation problem using network slicing by considering the xAPPs. We propose a novel method for resource allocation, with the goal of effectively coordinating numerous independent xAPPs to address the network slicing and resource allocation challenges within O-RAN.
The application of ML algorithms in wireless networks has surged due to the escalating complexity of these networks and the challenges posed by crafting custom optimization models. Traditional approaches, such as convex/non-convex optimization, are grappling with mounting complexity stemming from shifts in network architecture, evolving performance requirements, and the proliferation of devices. Present methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often neglect the capacity to learn resource distributions. Typically, these methods involve designing an agent for a specific task and training it from scratch, resulting in reduced exploration efficiency and extended convergence times. Moreover, they necessitate a substantial volume of training samples, leading to prolonged exploration phases. The limited adaptability of traditional ML methods to novel scenarios motivates us to explore an alternative approach capable of superior generalization, more efficient data utilization, and resource distribution learning. To address these objectives, we propose a novel methodology that employs a self-supervised learning technique, specifically contrastive learning [17], in conjunction with a variational autoencoder (VAE) model, known as a type of generative model [18]. These models excel in handling uncertainty by learning potential outcome distributions, thereby offering a distribution of predictions instead of a singular deterministic output. Moreover, their adeptness in handling incomplete or missing data allows for the imputation of missed values based on learned data patterns. VAEs also exhibit proficiency in semi-supervised learning setups, effectively harnessing both labeled and unlabeled data to enhance predictive accuracy. Their generative abilities facilitate the creation of new data instances resembling the training set, proving beneficial in scenarios demanding synthetic data generation or heightened resilience against noise and outliers.
Furthermore, contrastive learning involves contrasting different examples to acquire informative data representations. On the other hand, a VAE comprises an encoder and decoder that jointly learn the underlying probability distribution of data. By integrating contrastive learning with VAE, we can mitigate the common challenge of requiring a large number of training samples in ML. This approach can enhance performance metrics, such as quicker convergence and improved results, while also reducing the time and effort needed for algorithm design and training.
I-A Contribution
This paper introduces a system model within the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC, through resource allocation, with a particular emphasis on network slicing.
Our paper’s primary contributions can be summarized as follows:
  • Our research primarily aims to tackle resource allocation challenges in the O-RAN architecture, with a specific focus on optimizing weighted throughput by allocating resources like transmission power and physical resource blocks. We account for the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC. Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
  • Our solution to the resource allocation optimization problem employs a two-step algorithm. Firstly, we simplify the problem by framing it as an optimization task. Secondly, we employ a two-phase training approach. We use supervised learning with a VAE to train power transmission regression and resource allocation decisions. This is followed by unsupervised learning with a contrastive loss method to enhance the model’s generalization and robustness. In essence, our approach simultaneously handles power and radio resource allocation xAPPs. In addition, it is not limited to O-RAN but can be applied to future multi-vendor O-RAN architectures, leveraging team learning principles developed by our team.
I-B Structure of Paper
The paper is structured as follows: Section II presents a review of related works. In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem. Section IV provides a detailed description of the proposed resource allocation algorithm. In Section V, we present numerical results used to evaluate the algorithm’s performance. Finally, Section VI offers our conclusion.
II Related Works
As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements. This challenge becomes more pronounced in 5G networks due to their increased complexity and stringent QoS demands [16]. Recently, there has been a growing focus on addressing the network slicing problem in cellular networks, as evidenced by several studies [16, 19, 20, 21, 22].
Feng et al. [16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements. However, their simulation utilized only a limited number of User Equipment (UE) devices and service types, which may not accurately represent real-world 5G networks with diverse service requirements. In another study, a resource allocation approach based on the M/M/1 queuing system was presented for RAN slicing [19]. Nevertheless, practical scenarios often involve bursty traffic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services, overly optimistic. Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process. Additionally, in a study by Kazemifard et al. [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center. The authors conducted a theoretical analysis and employed a mathematical model to efficiently address minimum delay problems.
Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23]. ML algorithms, in particular, have found application in power control and resource allocation within wireless networks. These models are trained on historical resource allocation decisions and then applied to make new allocation decisions.
For instance, the SAMUS framework was employed by Bektas et al. [24] to forecast incoming traffic, enabling an exploration of various network slicing operations and a discussion of the trade-off between the data rate of eMBB and the latency of URLLC. Li et al. [25] introduced a method that utilizes DRL for proactive resource allocation to mobile UE, considering both time and frequency domains while accounting for mobility. Wu et al. [26] proposed a dynamic resource allocation method for radio and computation resources through RAN slicing, including a defined learning scheme with constraints.
Further, Ayala et al. [27] presented a method for the joint allocation of virtualized RAN’s radio and computation resources using a DRL approach based on an actor-critic network. Wang et al. [28] introduced a DRL-based self-play approach that can be integrated into RAN architecture to efficiently manage RU and DU resources, leading to cost-effective RAN operations. In another approach, Ndikumana et al. [29] combined federated learning and Deep Q-Learning with a reward function to reduce offloading, fronthaul routing, and computation delays in O-RAN.
Similarly, Samarakoon et al. [30] introduced a federated learning algorithm to optimize transmit power and radio resource allocation in vehicular networks. Additionally, Elsayed et al. [31] proposed a reinforcement learning-based method for joint power and radio resource allocation in 5G networks.
III System Model and Problem Formulation
In our study, we address network slicing and resource allocation problem within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice. The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery, while eMBB slices focus on high data rates for mobile broadband. These slices have also various QoS requirements for network resource allocation [10].
By slicing the RAN into dedicated components for URLLC and eMBB, we can tailor network optimization to distinct traffic types and services, enhancing the efficient and flexible utilization of network resources. Key resources in O-RAN allocation encompass transmission power and Physical Resource Blockss. Traffic demands, often quantified by data rate and bandwidth, reflect the volume of data requiring transmission through the network. QoS requirements delineate the desired network performance based on parameters such as data rate. Our system model for resource allocation with network slicing is illustrated in Figure 1.
Figure 1: O-RAN network system model
In our study, we investigate an O-RAN-based deployment, wherein each base station, known as RU in O-RAN, accommodates two distinct slices: eMBB and URLLC. These slices cater to UEs with comparable QoS requirements. Our analysis encompasses two critical network functions: power control and radio resource allocation, which are jointly considered. We denote the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$. Each base station $b$ serves two sets of users: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ users, where $U=U_{e,b}+U_{u,b}$. Each user is exclusively associated with a single RU. The temporal and spectral layout is partitioned into $\mathcal{M}$ segments referred to as PRBs. These PRBs constitute the smallest unit of resources for allocation [32]. Each PRB spans a time frame, defined as a time slot in the temporal domain, and covers a frequency range of $f$ Hz in the spectral domain. We also denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices. The total number of slices ($S$) will be: $S=S_{e}+S_{u}$.
The user association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if user $u$ is connected to RU $b$ in slice $s$, and 0 if not. Since each UE should be connected to only one RU, the following constraint must be met:
$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when user $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
Similarly, each PRB $m$ of RU $b$ should be associated with only one UE, as defined below:
$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
The rate achieved for a user $u$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR). The calculation for $R_{u,b,m,s}$ is determined by:
$\eta_{u,b,m,s}=\frac{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}{\sum_{j\neq b}^{N}\sum_{l\neq s}^{S}\sum_{i\neq u}^{U}\alpha_{i,j,l}h_{i,j,m,l}p_{i,j,m,l}+\sigma^{2}},$ (4)
$R_{u,b,m,s}=\log_{2}(1+\eta_{u,b,m,s}),$ (5)
$R_{u,s}=\sum_{b\in B}\sum_{m\in M}\alpha_{u,b,s}\beta_{u,b,m,s}R_{u,b,m,s},$ (6)
$R_{u}$ $=\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\alpha_{u,b,m,s}\sum_{m\in\mathcal{M}}\beta_{u,b,m,s}R_{u,b,m,s}$
$=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}\beta_{u,b,m}R_{u,b,m,s}\;\;\;\;\forall m\in\mathcal{M},\;s\in\mathcal{S}.$ (7)
The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively. The existence of quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, particularly when the original signal displays significant variations or spans a wide range of values. This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE. These values are determined as follows:
$P_{b}=\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}+\sigma_{1}^{2},$ (8)
$C_{b}=\log_{2}(1+\frac{\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{b,m,s}}}{\sigma_{1}^{2}})$
$=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}),$ (9)
where, we denote the transmission power of PRB $m$ in RU $b$ in slice s as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$. Furthermore, we define the channel gain between the resources and UE as $h_{ij}$. To represent this relationship in matrix form, we construct a matrix that accounts for $\mathcal{B}$ resources and $\mathcal{U}$ users as follows:
$H:=\begin{bmatrix}|H_{11}|^{2}&\cdots&|H_{1U}|^{2}\\ \vdots&\vdots&\vdots\\ |H_{B1}|^{2}&\cdots&|H_{BU}|^{2}\end{bmatrix}_{B\times U}$ (13)
The overall propagation delay ($D^{pro}$) comprises three distinct components: the propagation delay in the fronthaul link ($D^{fr,p}$), the midhaul link ($D^{mid,p}$), and the backhaul link ($D^{b,p}$). In each of these links, the propagation delay is determined by the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ signifies the length of the fiber link, and $C$ represents the propagation speed of the medium.
Similarly, the total transmission delay ($D^{tr}$) is the aggregate of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links. Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, represented as $D=\frac{\alpha}{R}$, where $R$ indicates the data rate of the packet, and $\alpha$ denotes the mean packet size.
$D_{total}=D^{pro}+D^{tr},$ (14)
$D_{u,s}=\frac{L}{C}+\frac{\alpha}{R_{u,s}}.$ (15)
Therefore, the delay of UE $u$ in slice $s$, as derived from Equation 15, should be constrained to be below a specific threshold relevant to its respective slice, otulined below:
$D_{u,s}\leq D^{min}_{u,s}.$ (16)
The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring compliance with QoS requirements. In this context, $w_{s}$ signifies the priority associated with individual service types and UEs. The intelligent RU is responsible for effectively managing the requirements of both network slices. Hence, we can mathematically express the problem in the following manner:
$\max_{\mathbf{p,\alpha,\beta}}$ $\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}w_{s}R_{u,s}$ (17)
subject to
$P_{b}\leq P_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (18)
$p_{u,b,m,s}\geq 0\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (19)
$p_{u,b,m,s}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (20)
$R_{u,s}\geq R_{s}^{max}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (21)
$C_{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (22)
$D_{u,s}\leq D^{min}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (23)
$\alpha_{u,b,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (24)
$\beta_{u,b,m,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (25)
$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}$ (26)
$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},u\in\mathcal{U},s\in\mathcal{S}$ (27)
$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (28)
IV Methodology
Our primary focus is to address resource allocation problem within O-RAN architecture, and we propose a novel approach based on semi-supervised learning. Our objective is to optimize the sum of rates among UEs while considering constraints, including transmission power and PRB allocation. To accomplish this, we compile a dataset comprising diverse resource allocation scenarios, encompassing various combinations of objective functions and constraints. We generate this dataset using an exhaustive search algorithm (ESA), which explores the entire feasible solution space.
Our research primarily delves into the resource allocation problem within the O-RAN architecture, and we introduce an innovative approach based on semi-supervised learning. Our primary goal is to enhance the cumulative rates for UEs while adhering to specific constraints such as transmission power, delay, and PRB allocation. To achieve this, we curate a comprehensive dataset that encompasses a wide range of resource allocation scenarios, covering diverse combinations of objective functions and constraints. This dataset is generated through the implementation of an ESA, which thoroughly explores the entire feasible solution space.
In our approach to addressing the resource allocation problem, we employ a deep neural network (DNN) model integrated with a contrastive loss function. This DNN model efficiently manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions. Through the training of our model on the dataset we’ve created, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem. This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
We aim to evaluate the performance of our learning-based approach using multiple regression-based performance metrics to assess its robustness. Additionally, we will validate the effectiveness of our model by comparing its outcomes with those of the ESA in different scenarios. Moreover, while the ESA is the optimal algorithm for resource allocation decisions, its complexity makes it impractical for deployment in real-world environments. Therefore, we will also compare the performance of our proposed algorithm with state-of-the-art methods known for their practical deployment capabilities, such as DRL. This comprehensive assessment will demonstrate the method’s reliability across diverse scenarios.
Figure 2 illustrates the overview of our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
Figure 2: The overview of proposed methodology.
IV-A Exhaustive Search Algorithm
The ESA is a well-established optimization approach for solving resource allocation problems in wireless networks. Its fundamental concept involves generating all possible resource allocation combinations and selecting the best one based on a predefined objective function [33]. ESA serves as a valuable benchmark solution, providing a ground truth against which results from other methods can be validated and compared. Its advantages include simplicity of understanding and implementation, along with the assurance of yielding an optimal solution within a finite time frame.
Nonetheless, ESA’s primary limitation lies in its computational complexity, especially when addressing resource allocation problems with extensive dimensions and intricate constraints. The computational burden grows exponentially with the size of the problem, rendering ESA impractical for large-scale resource allocation tasks [34]. Despite this drawback, ESA remains a popular choice as a baseline for comparing various optimization algorithms and offers a crucial reference for evaluating new optimization strategies.
To assess the performance, validity, and robustness of the proposed learning-based method, we will compare its results with those generated by ESA. Additionally, a subset of solutions for the resource allocation and network slicing optimization problem within O-RAN will be generated using ESA. This labeled dataset will be utilized to train the supervised learning component of the proposed method.
IV-B Deep Reinforcement learning Algorithm
As discussed in Section II, DRL algorithms have proven to be effective solutions for addressing resource allocation problems. These algorithms have been successfully deployed in various real-world wireless network environments, demonstrating their practicality and their ability to significantly improve both technical and business metrics [35]. Consequently, we utilize DRL, where an agent interacts with the environment to make resource allocation decisions and receive related rewards. These rewards are tied to the objective function outlined in Problem 17, considering its objectives and constraints. Next, we model the problem as a Markov Decision Process (MDP) and introduce the DRL algorithm to address the resource allocation challenge in network slicing. This approach serves as an extra means to evaluate the robustness and our primary proposed solution.
IV-B1 Markov Decision Process
In the specified resource allocation problem (Equation 17), an O-RAN orchestrator is responsible for determining PRB allocation decisions and determine the corresponding RU for each UE for entw. This problem can be represented as a MDP denoted by $(\mathfrak{S},\mathfrak{A},\mathfrak{P},\mathfrak{R})$, where $\mathfrak{S}$ represents the state space, $\mathfrak{A}$ is the action space, $\mathfrak{P}(\mathfrak{s}^{\prime}|\mathfrak{s},\mathfrak{a})$ signifies the transition probability from state $\mathfrak{s}$ to state $\mathfrak{s^{\prime}}$ by taking action $\mathfrak{a}$, and $\mathfrak{P}$ represents the reward function [35, 36, 37, 38], which are elaborated for our specific problem.
  1. I.
    State: The state at time $t$ ( $\mathfrak{s}(t)\in\mathfrak{S}$), is represented by $\{\mathfrak{s}_{u}(t)\}_{t=1}^{N}$, where $\mathfrak{s}_{u}(t)$ indicates the state of UE $u$ ($u=1:U$). Here, $\mathfrak{s}_{u}(t)$ is a binary indicator: it equals 1 if the data rate requirement of UE $u$ is met and 0 otherwise. It’s important to note that each data rate corresponds to a specific transmission power level.
  2. II.
    Action: An action ($\mathfrak{a}\in\mathfrak{A}$), is represented as $\mathfrak{a}=\{\alpha_{u,b,s},\{\beta_{u,b,m,s}\}_{m=1}^{\mathcal{M}}\}_{b=1}^{\mathcal{B}}$.
  3. III.
    Transition Probability: The transition probability $\mathfrak{P}(\mathfrak{s}|\mathfrak{s},\mathfrak{a})$ represents the probability of moving from one state $\mathfrak{s}$ to another one $\mathfrak{s}^{\prime}$ by taking action $\mathfrak{a}$. It is influenced by several factors, including the QoS requirements of UEs and the resource allocation decisions made for UEs.
  4. IV.
    Reward: The reward function serves as a pivotal element in assessing the gains or expenses associated with a given state-action pair. It provides guidance to the agent’s decision-making process, indicating the appeal of different actions across various states. In alignment with earlier studies [35, 36, 37], our method characterizes the reward function as a combination of the objective function and the constraints relevant to the considered problem (17). Therefore, we establish the reward function for our specific problem as follows:
    $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u,s}+\Theta_{const.}\;C_{u,b,m,s}+\Theta_{bias},$ (29)
    where, $\Theta_{r}$, $\Theta_{const.}$, and $\Theta_{bias}$ represent the respective weights assigned to the objective function (Data Rate), constraints, and the bias value.
The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP. This policy maps any state $\mathfrak{s}\in\mathfrak{S}$ to a probability distribution $\pi(\cdot|\mathfrak{s})$ over the set of possible actions $\mathfrak{A}$. The goal is to find a policy that maximizes the cumulative reward obtained through dynamic learning from acquired data. The policy $\pi$ can be influenced and controlled by decision makers.
The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state. Additionally, the action value function $Q^{\pi}(\mathfrak{s},\mathfrak{a})$ is introduced to represent the expected cumulative reward when taking action $a$ in state $\mathfrak{s}$, considering that the policy $\pi$ is being followed. This can be expressed as follows:
$V^{\pi}(\mathfrak{s})=\mathbb{E}[Q^{\pi}(\mathfrak{s},\mathfrak{A})|\mathfrak{A}\sim\pi(.|\mathfrak{s})],$ (30)
$Q^{\pi}(\mathfrak{s},\mathfrak{a})$ $=\mathbb{E}[\sum_{t=0}^{\inf}\gamma^{t}\cdot\mathfrak{R_{t}}|\mathfrak{S_{0}}=\mathfrak{s},\mathfrak{A_{0}}=\mathfrak{a}]$
$=\mathfrak{r}(\mathfrak{s},\mathfrak{a})+\gamma\;\mathbb{E}[V^{\pi}(\mathfrak{S}^{\prime})|\mathfrak{S}^{\prime}\sim\mathfrak{P}(.|\mathfrak{s,a})],$ (31)
where, $\mathfrak{R}_{t}$ represents the immediate reward obtained at time $t$, and $\gamma$ (where $10<\gamma<$ 1) denotes the discount factor applied in the computation of future rewards. The discount factor determines the weight of future rewards in the decision-making process.
Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task. Determining the value function in such scenarios presents a significant challenge. To address MDPs, two primary strategies are employed: model-based and model-free approaches. In instances where there is uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
IV-B2 Deep-Q-Network Algorithm
In our considered problem, we face the challenge of managing the states space ($\mathfrak{S}$) and actions space ($\mathfrak{A}$) that are inherently high-dimensional and complex. This complexity arises from continuous variable like transmission power, as well as the substantial number of UEs and PRB resources. Traditional Q-learning methods struggle in such scenarios, as many state-action pairs are rarely encountered, leading to inefficient learning processes [35, 38].
To overcome this challenge, we turn to the Deep Q-Network (DQN) approach to tackle our resource allocation problem and optimize parameters such as transmission power, alpha, and beta. DQN offers distinct advantages in handling these complex, high-dimensional state-action spaces by utilizing a deep neural network to approximate the $Q$-function. DQN’s capacity to generalize from prior experiences and its use of techniques such as experience replay and target networks make it an apt choice for addressing resource allocation challenges within our problem domain.
DQN employs experience replay to store into replay memory ($\mathbb{M}$) and sample transitions from the memory ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$) to train the neural network with stochastic gradient descent. This method helps acquire uncorrelated samples from a temporally correlated MDP trajectory, enhancing gradient estimation accuracy. A target network $Q_{\theta^{*}}$ is also used, where independent samples are obtained from the replay memory. These samples are used to update the $Q$-network’s parameter, improving learning stability and efficiency [38]. We compute the target value and update $\theta$ by the gradient of equations 32 33.
$Y=\mathfrak{R(s,a)}+\gamma\;\max_{\mathfrak{a}^{\prime}\in\mathfrak{A}}Q_{\theta^{*}}(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime}),$ (32)
$L(\theta)=||Y-Q_{\theta}(\mathfrak{s},\mathfrak{a}||^{2},$ (33)
The parameter $\theta^{*}$ is updated every $T_{target}$ steps by setting it equal to $\theta$. The target network remains unchanged for $T_{target}$ steps and is then updated using the current weights of the Q-network.
Our DQN training employs the epsilon-greedy policy for action selection. This means that the agent chooses the action with the highest Q-value with a probability of $1-\epsilon$, and with a probability of $\epsilon$, it explores and selects a random action. The policy function is represented as:
$\pi(S)=\begin{cases}\text{argmax}_{\mathfrak{a}}Q(\mathfrak{S,a})&\text{with probability }1-\epsilon\\ \text{random action}&\text{with probability }\epsilon\end{cases}$
It’s worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
IV-C Proposed Method
In recent times, DL techniques have found application in enhancing the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems, as demonstrated in several studies [39, 40, 41, 42]. These approaches excel in learning complex, non-linear relationships between input features like channel matrices and output decisions, such as power and resource allocation. What distinguishes DL-based resource allocation is its ability to effectively manage the dynamic and intricate nature of wireless communication systems. These approaches can adapt to varying channel conditions and UE, resulting in improved network performance compared to traditional optimization methods.
In our proposed DL-based resource allocation method, we feed the channel matrix into a DNN. This network is responsible for learning optimal resource allocation decisions that adhere to network constraints while maximizing the defined objective function (weighted aggregated data rate). We also employ a flattened vector with dimensions $UB$, derived from the channel gain matrix, as input for a machine-learning model, which includes a VAE. This autoencoder utilizes the channel matrix to generate the power vector. The key innovation in our approach is the introduction of contrastive loss as an additional term in the VAE’s loss function. This addition refines the learning of representations, significantly enhancing the accuracy of solutions to resource allocation problems. Subsequent sections will provide a detailed exploration of each component in our proposed method.
IV-C1 Variational Autoencoder
The VAE is a specific type of generative model frequently applied in unsupervised learning tasks. It functions as a DNN designed to generate new data samples resembling a provided set of training data. Utilizing a VAE confers a notable advantage: it adeptly captures the intrinsic distribution of the channel gain matrix. This acquired knowledge proves immensely valuable when confronting the resource allocation problem [43].
In a VAE, let’s denote the input data as $x$, a latent variable representing a lower-dimensional data representation as $z$, and $p_{\theta}(x)$ as a generative model representing the data distribution based on the latent variable. The primary goal of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$. This likelihood can be mathematically expressed as:
$\log p_{\theta}(x)=\mathbb{E}{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{KL}(q_{\phi}(z|x)|p(z)),$ (34)
where, $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable given the input data. The term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable.
The objective function can be enhanced by adjusting the parameters of both the generative model, denoted as $\theta$, and the encoder network, represented as $\phi$. This optimization process involves iteratively updating these parameters and computing the gradient of the objective concerning them, as outlined in [43]. Figure 3 illustrates the VAE architecture tailored for addressing the resource allocation and network slicing problem.
Figure 3: The architecture of proposed VAE.
To provide a more detailed explanation, a VAE comprises two primary components: an encoder and a decoder. The encoder accepts the channel gain matrix as input and transforms it into a lower-dimensional representation known as a latent code. Subsequently, the decoder takes this latent code as input and produces an estimation of the transmission power. These two components work in tandem during training to minimize a reconstruction loss. This loss metric quantifies the disparity between the actual transmission power and the predicted transmission power, guiding the network to generate more accurate power predictions.
Using a VAE for resource allocation offers several distinct advantages. Firstly, VAEs exhibit remarkable flexibility, enabling them to model intricate distributions. This characteristic is particularly advantageous for capturing the complex and non-linear relationship between channel gains and transmission powers. Secondly, VAEs excel at discerning the underlying structure within data, enabling them to make informed predictions even when confronted with previously unseen data [44]. Thirdly, the latent code generated by the encoder serves as a compressed representation of channel gains, which can be leveraged for diverse tasks like visualization, clustering, and anomaly detection.
Moreover, resource allocation problems are inherently non-convex and can be exceptionally challenging when using traditional optimization approaches. The introduction of VAEs provides a more versatile and effective means of tackling these intricate problems.
Our dataset comprises a total of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples. These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$).
It is important to note that the process of generating these input channel matrices is computationally less intensive compared to traditional methods, as documented in previous research [45].
To train the supervised learning component of our algorithm, we employ the ADAM optimizer [46]. This optimizer facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples. The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
$L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||VAE(H_{i})-\Gamma_{i}||^{2}.$ (35)
In our resource allocation challenge, the correlation between the xApps dealing with power estimation and the PRBs allocation emerges from their shared foundation within the VAE framework. The VAE operates as a unifying space where both xApps converge, enabling a synergistic exploration and consideration of the pertinent features relevant to both power transmission and PRB allocations. Through this shared space, the xApps gain access to a common latent representation of the channel gain matrices. This shared understanding facilitates a cohesive analysis where the interdependencies between power estimation and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the critical features that impact both domains. This integration of information aids in making more comprehensive and informed resource allocation decisions, as it ensures that both xApps leverage a unified understanding of the underlying data to optimize their respective objectives. This coordination is pivotal, as it allows for a coordinated and synchronized approach to resource allocation within the O-RAN architecture.
IV-C2 Contrastive Loss
Contrastive loss is a widely-used loss function in ML for various representation learning tasks, including image or text classification. Its fundamental aim is to construct a representation of data points that encourages similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17].
Mathematically, the contrastive loss can be expressed as follows:
$L=\frac{1}{2N}\sum_{i=1}^{N}(y_{i}\cdot d^{2}(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$
$+(1-y_{i})\cdot\max(0,m-d^{2}(\mathbf{x}_{i},\mathbf{x}_{i}^{-}))),$ (36)
where $N$ is the number of data points, $\mathbf{x}_{i}$ is the representation of the $i^{th}$ data point, $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ are the representations of the similar and dissimilar data points respectively, $y_{i}$ is a binary label indicating whether $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$ are similar or dissimilar, and $d(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$ is the distance function between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$. The parameter $m$ is a margin that separates the similar and dissimilar representations.
To use contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices. One way to create channel gain matrices that are either similar or dissimilar is by using a statistical model of the channel that captures the relevant properties of the communication environment. The Rayleigh fading model is a commonly used statistical model for wireless channels, which assumes that the channel gain is a complex Gaussian random variable with zero mean and a variance that depends on the distance between the transmitter and receiver [47, 48].
To create similar channel gain matrices, coefficients with similar magnitude and phase can be generated. Conversely, to create dissimilar channel gain matrices, coefficients with different magnitude and phase can be generated by introducing randomness into the channel model by using a Rayleigh fading model. To generate a channel gain matrix using this model, a matrix of complex Gaussian random variables with zero mean and a variance that depends on the distance between the transmitter and receiver can be generated as well as considering the interferance as noise [49, 50]. The variance can be adjusted to create similar or dissimilar channel gain matrices based on the desired level of correlation between the channels. The mathematical formulation of this approach can be expressed as:
$\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon},$ (37)
where $\mathbf{H}_{noise}$ is the channel gain matrix with added noise, $\mathbf{H}$ is the original channel gain matrix, and $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and standard deviation $\sigma$.
Once we have created similar and dissimilar channel gain matrices, we can train our contrastive loss function using these matrices as inputs. The loss function can be formulated as:
$L_{Contrastive}(\theta)=\frac{1}{\beta_{L}}$
$\sum_{i\in\beta_{L}}[-\log\frac{\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{i}})}{\tau}}}{\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{i}})}{\tau}}+\sum_{i\neq j}\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{j}})}{\tau}}}],$ (38)
where $\theta$ is the set of parameters in the variational autoencoder model, $\tau$ is a temperature hyperparameter and $L(\theta)$ is the contrastive loss function. $\underline{H_{i}}$ and $\bar{H_{i}}$ are the similar matrices to the the main channel gain matrix ($H_{i}$), which allow us to train our proposed model [51]. By minimizing this loss function, we can learn a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices and produce a power control decision that optimizes the network performance.
IV-C3 Training the proposed model
In this section, we will describe the training phase of our proposed method, which consists of two parts: the supervised learning part, where the VAE is trained to regress the power transmission, and the unsupervised learning part, where the contrastive loss is used to improve the generalization and robustness of our proposed method. There are several key parameters that should be taken into consideration in training the supervised part. The latent space dimension represents the size of the hidden representation learned by the VAE [45]. A larger latent space dimension can lead to a more expressive model, but also increases the risk of overfitting. The learning rate determines the speed at which the model learns from the data. The batch size determines the number of samples used for each iteration of the training process. The number of epochs determines the number of times the model will be trained on the data. $L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [52]. The proposed model is developed using the PyTorch [53] software framework. Both validation size and test size are $20\%$ of the dataset, allowing for a more thorough evaluation of the proposed model’s performance. Hyper-parameter tuning yields the value of hyperparameters, which can be found in the Table III.
The initial model was pre-trained through self-supervised learning, using unlabeled channel gain matrices that corresponded to the stated loss function. The pre-training was based on regression loss. After pre-training, the DL model was fine-tuned for power transmission regression using the proposed method. The performance of the trained model was evaluated on test data points using various performance metrics, while the number of labeled training channel gain matrices were varied. Additionally, a supervised baseline performance for the regression task was obtained by training the DL model with all available labeled training channel gain matrices.
IV-C4 Performance Metrics
The performance metrics for evaluating our proposed method are Mean Absolute Error (MAE), cosine similarity, and correlation. MAE measures the average absolute difference between the predicted values and the actual values. Cosine similarity measures the similarity between two vectors in a multi-dimensional space. The Pearson correlation ($R$) represents the amount of variation in the output-dependent attribute that can be predicted by the input-independent variable. It also provides insight into how accurately a model can predict outcomes based on how much of the variation in those outcomes can be explained by the input-independent variable. They can be calculated as follows:
$MAE=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|,$ (39)
[…]
$correlation(\hat{Y},Y)=\frac{\text{cov}(\hat{Y},Y)}{\sigma_{\hat{Y}}\sigma_{Y}},$ (41)
where $N$ is the total number of samples, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the predicted value, and $\cdot$ represents the dot product, and ||$\cdot$|| ⋅ | represents the magnitude of a vector. As for correlation, $\text{cov}(\hat{Y},Y)$ is the covariance between the predicted value and the true value, and $\sigma_{\hat{Y}}$ and $\sigma_{Y}$ are the standard deviations of the variables.
V SIMULATION RESULTS
In the following subsections, we will begin by analyzing the initial values to simulate the proposed algorithm. Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN. This comparative analysis aims to assess the accuracy and effectiveness of our proposed DL model in addressing the resource allocation problem.
V-A Initial points and performance of training the model
In this section, we present numerical results for the considered resource allocation problem to evaluate the performance of our proposed model (SS-VAE) compared to state-of-the-art methods, namely ESA and DQN. The initial values and parameters used to run and simulate our proposed algorithm (SS-VAE) and the benchmark algorithms (ESA and DQN) are provided in Table III, Table III, and Table III.
TABLE I: Resource Allocation Simulation Parameters
TABLE II: DQN Training Parameters
TABLE III: Proposed Model Training Parameters
Parameter Value
Noise power -174 dBm
bandwidth of each sub-carrier 180 KHz
Maximum transmit power of each O-RU 40 dBm
Maximum fronthaul capacity 46 bps
Maximum data rate for eMBB 20 bps
Maximum data rate for URLLC 2 bps
Maximum power for URLLC & eMBB 30 dBm
Number of episodes and steps 250, 50
$\gamma$, $\epsilon_{finalvalue}$, $\epsilon_{decayfactor}$ 0.9, 0.1, 0.9995
Replay memory size 400
Type Training Parameter Value
Optimizer Initial learning rate 0.001
$\beta_{1}$ 0.99
$\beta_{2}$ 0.99
Weight Decay 0.9
Model Number of encoding layer 20
Dimension of latent space 20
Contrastive loss temperature 0.25
Activation Function ReLU
Dropout Rate 0.3
Training Epoch 40
Batch Size 128
Train/Validation Split 0.2
TABLE II: DQN Training Parameters
TABLE III: Proposed Model Training Parameters
As previously discussed, our proposed model aims to estimate the transmission power corresponding to each input channel matrix. These estimated power values, derived from the weights of the VAE, are then used to predict the allocation indicators ($\alpha_{u,b,s},\beta_{u,b,m,s}$). To streamline the training process, it is crucial to convert the dimensions of each channel matrix data into a one-dimensional vector. More specifically, our model takes as input a flattened $UB$-dimensional vector obtained from the channel gain matrix.
The loss values for both the validation and training data points is depicted in Figure 4.
Figure 4: Loss values of training and validation dataset.
The decreasing trend of loss values illustrate that the proposed method (SS-VAE) could accurately estimate the target values (power of transmission, association indicators). In fact, the model can map the input channel gain matrix to the power of transmission as well as the indicators by employing the VAE’s property in learning the underlying information.
The performance metrics for the test data, as presented in Table IV, clearly indicate that our proposed method (SS-VAE), which estimates transmission power based on the input channel gain matrix, exhibits the best generalization. This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks. In addition, the pre-training method involving the contrastive loss has further enhanced our model’s capacity to learn the underlying features, ultimately reducing generalization errors.
TABLE IV: Performance metrics of the proposed model
Performance Metric MAE $R^{2}_{Score}$ Cosine Similarity
Value 0.09207532 0.76795106 0.988023
The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions. This suggests that a substantial proportion of the variability in the dependent variable (power of transmissions) can be accounted for by the independent variables (the channel gain matrix) integrated into the model. In simpler terms, the model is effective at describing and understanding the variations in the data, especially in terms of transmission power.
V-B Comparison of the algorithms
In the next phase, we will assess the performance of our proposed SS-VAE method by comparing it to the ESA algorithm. This evaluation will specifically consider the impact of the number of labeled samples used for training the supervised part of SS-VAE, as illustrated in Figure 5.
(a) Agg. throughput vs. Number of labeled data.
(b) Association Error vs. Number of labeled data.
Figure 5: Comparison of the algorithms (SS-VAE vs. ESA).
We analyze the aggregated throughput under varying numbers of labeled data points, as generated by our proposed SS-VAE algorithm and the ESA. As depicted in Figure 5a, our proposed model achieves a similar aggregated throughput when compared to the ESA. Moreover, the discrepancy between our model and the ESA decreases significantly as the number of labeled data points increases.
We further evaluate the performance of the SS-VAE model and ESA in terms of binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$). Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled channel gain matrices (data points) increases. Notably, the difference between the $L1-norm$ of binary variables obtained by the SS-VAE method significantly decreases in comparison to ESA as the number of labeled data points grows.
The results above highlight the reliable performance of our proposed method (SS-VAE) in estimating transmission powers, user association, and PRB allocation when compared to the ESA. In the subsequent section, we will assess its performance under different initial conditions, varying the number of slices, UEs, and thresholds for transmission power and rate. Since the SS-VAE exhibits lower error compared to the ESA but doesn’t perform identically in the resource allocation problem, we will also evaluate its performance alongside another benchmark algorithm, the aforementioned DQN algorithm in Section IV-B. DQN has demonstrated practical and valuable results in resource allocation problems within wireless networks and similar domains, enabling us to assess the robustness of our proposed algorithm.
V-C Performance results
In this section, we will assess the performance of the proposed model (SS-VAE) in the resource allocation problem under various scenarios. To begin, we will examine the aggregate weighted rate in relation to the number of UEs using three different algorithms: SS-VAE, ESA, and DQN, as depicted in Figure 6.
(a) Aggr. throughput vs. number of UEs.
(b) Aggr. throughput vs. number of UEs ($P_{O-RU}$).
Figure 6: Aggregated throughput of the SS-VAE vs. Number of UEs.
Figure 6a demonstrates that the proposed approach (SS-VAE) exhibits a performance that is comparable to the ESA. There are no significant differences in terms of aggregated throughput between the two. Additionally, SS-VAE achieves a superior aggregated throughput across all UE numbers when compared to the DQN algorithm. This highlights the robustness and high accuracy of our proposed method for the resource allocation problem. It’s worth noting that the sum of data rate increases initially as more UEs join the service for all three examined algorithms. However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints.
Figure 6b depicts the aggregated throughput achieved by the SS-VAE algorithm concerning different power levels of O-RU and varying numbers of UEs. As per established knowledge [15], both increasing the power of O-RU and the number of UEs lead to higher aggregated throughput. The highest aggregated throughput is observed when the power of O-RU is at its highest.
Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, in relation to the power values. This presentation aims to evaluate the performance of our proposed algorithm for these services in comparison to a benchmark algorithm.
Figure 7 displays the aggregated throughput for the two considered service types, namely eMBB and URLLC, as determined by the SS-VAE and DQN algorithms. This visualization is intended to evaluate the performance of our proposed algorithm (SS-VAE) for these services in comparison to a benchmark algorithm.
(a) Aggr. throughput vs. O-RU power.
(b) Aggr. throughput vs. Power of slice.
Figure 7: Aggregated throughput of SS-VAE vs. Power values.
Figure 7a depicts the influence of O-RU’s maximum power thresholds on the aggregate weighted rate for both eMBB and URLLC services, comparing two algorithms, SS-VAE and DQN. The results indicate a direct correlation between increased maximum power and higher weighted rates. Additionally, the SS-VAE algorithm consistently outperforms the DQN algorithm, achieving higher aggregated throughput for both service types across various power values.
Similarly, in Figure 7b, variations in total data rates are displayed concerning the maximum power for distinct cases of eMBB and URLLC services. The results reveal a direct increase in data rates as the maximum power of each UE in different service instances rises. Notably, the SS-VAE algorithm consistently outperforms the DQN algorithm, securing higher aggregated throughput for both service types across various power threshold values.
VI CONCLUSION
In this paper, a novel model is introduced for the O-RAN system’s downlink, aiming to coordinate multiple independent xAPPs to tackle the resource allocation problem with network slicing in the O-RAN architecture. This problem is addressed for two specific 5G services, namely eMBB and URLLC, with the overarching objective of optimizing the aggregate weighted rate by managing RU associations, power allocation, and PRBs.
Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we introduced a semi-supervised learning approach as an efficient solution for UE and PRB allocation. Our two-step algorithm starts by reformulating the problem into an optimization problem considering the objective function and constraints. The second step consists of a two-part training process: firstly, supervised learning employing a VAE to estimate power transmission regression and association indicators, and then unsupervised learning with a contrastive loss approach, enhancing the model’s generalization and robustness.
To evaluate our proposed algorithm’s performance, we conducted comparisons with benchmark algorithms ESA and DQN across various scenarios. The results indicate that our approach achieves a rate similar to ESA while outperforming the DQN algorithm, highlighting its effectiveness. We further explore different scenarios tailored to each service, eMBB and URLLC, considering their unique QoS requirements. Through simulations, we demonstrate the superior efficiency of our model, particularly in terms of sample efficiency, making it a compelling solution for resource allocation challenges within the O-RAN system.
In future research, we plan to explore several areas. Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and investigate the impact of deployment costs associated with network function placement tackle. Furthermore, addressing UE mobility is crucial in practical scenarios. In addition, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, providing new research avenues in this dynamic field.

### v2

@@
 PRB
-Physical Resource Blocks
+Physical Resource Block
 SINR
@@
 Deep Q-Network
+BS
+Base Station
 KL
@@
 Semi-Supervised Learning Approach for Efficient Resource Allocation with Network Slicing in O-RAN
-Salar ${}^{\dagger}$    Mojdeh Karbalaee ${}^{\dagger}$     Vahid ${}^{\dagger}$       Seyed Pooya ${}^{\dagger}$
+Salar Nouri    Mojdeh Karbalaee Motalleb     Vahid Shah-Mansouri       Seyed Pooya Shariatpanahi
 $\dagger$ School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
@@
 Abstract
-The Open Radio Access Network (O-RAN) technology has emerged as a promising solution for network operators, providing them with an open and favorable environment.
-Ensuring effective coordination of x-applications (xAPPs) is crucial to enhance flexibility and optimize network performance within the O-RAN.
-In this paper, we introduce an innovative approach to the resource allocation problem, aiming to coordinate multiple independent xAPPs for network slicing and resource allocation in O-RAN.
-Our proposed method focuses on maximizing the weighted throughput among user equipments (UE), as well as allocating physical resource blocks (PRBs).
-We prioritize two service types, namely enhanced Mobile Broadband and Ultra Reliable Low Latency Communication.
-To achieve this, we have designed two xAPPs: a power control xAPP for each UE and a PRB allocation xAPP.
-The proposed method consists of a two-part training phase, where the first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission as well as the user association and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model.
-We evaluate the performance of our proposed method by comparing its results to those obtained from an exhaustive search algorithm, deep Q-network algorithm, and by reporting performance metrics for the regression task.
-We also evaluate the proposed model’s performance in different scenarios among the service types.
-The results show that the proposed method is a more efficient and effective solution for network slicing problems compared to state-of-the-art methods.
+This paper introduces an innovative approach to the resource allocation problem, aiming to coordinate multiple independent x-applications (xAPPs) for network slicing and resource allocation in the Open Radio Access Network (O-RAN).
+Our approach maximizes the weighted throughput among user equipment (UE) and allocates physical resource blocks (PRBs).
+We prioritize two service types: enhanced Mobile Broadband and Ultra-Reliable Low-Latency Communication.
+Two xAPPs have been designed to achieve this: a power control xAPP for each UE and a PRB allocation xAPP.
+The method consists of a two-part training phase.
+The first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission, UE association, and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model.
+We evaluate the performance by comparing its results to those obtained from an exhaustive search and deep Q-network algorithms and reporting performance metrics for the regression task.
+The results demonstrate the superior efficiency of this approach in different scenarios among the service types, reaffirming its status as a more efficient and effective solution for network slicing problems compared to state-of-the-art methods.
+This innovative approach not only sets our research apart but also paves the way for exciting future advancements in resource allocation in O-RAN.
 Index Terms:
@@
 The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2].
-However, the pursuit of wireless excellence does not conclude with 5G; it merely marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology.
-Poised on the horizon, 6G holds the promise to once again redefine the boundaries of connectivity.
+However, the pursuit of wireless excellence does not conclude with 5G; it marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology.
+Poised on the horizon, 6G promises to redefine the boundaries of connectivity again.
 Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3].
-These advancements have the potential to not only revolutionize our modes of communication but also reshape the very fabric of our interconnected world.
+These advancements can potentially revolutionize our modes of communication and reshape the fabric of our interconnected world.
 Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4].
-These techniques aim to inject an unparalleled level of intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions.
+These techniques aim to inject unparalleled intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions.
 In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal.
-As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly evident that the evolution of wireless communication is an unceasing journey.
-Each generation builds upon the accomplishments of its forerunners, establishing a robust foundation for the next phase of wireless excellence.
-Resource allocation involves the meticulous distribution of finite network resources among numerous users, applications, or services [7].
-In 5G, the primary objective is to support a wide spectrum of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8].
-Each of these services necessitates distinct resource sets encompassing bandwidth, data rate, latency, power consumption, complexity, and reliability.
-As such, resource allocation within the 5G landscape must exhibit the flexibility to cater to these diverse demands while ensuring the efficient utilization of network resources.
-Network slicing is an innovative concept that tailors virtual networks (slices) within a unified physical infrastructure, addressing specific service types and use cases [9].
-In the context of 5G, it efficiently allocates resources for eMBB and URLLC [10] services.
+As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly clear that the evolution of wireless communication is an ongoing journey, establishing a robust foundation for the next era of wireless excellence.
+Resource allocation involves carefully distributing limited network resources among numerous users, applications, or services [7].
+In 5G, the main goal is to support a wide range of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8].
+Each service requires specific resource sets, including bandwidth, data rate, latency, power consumption, complexity, and reliability.
+Resource allocation within the 5G landscape needs to be flexible to meet these diverse demands while efficiently using network resources.
+Network slicing is an innovative concept that customizes virtual networks (slices) within a unified physical infrastructure to address specific service types and use cases [9].
+In 5G, it efficiently allocates resources for eMBB and URLLC [10] services.
 This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12].
-Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a key technology in the relentless pursuit of wireless excellence, bridging the transition from 5G to 6G.
-This concept seamlessly aligns with the architecture of the O-RAN.
+Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a critical technology in the transition from 5G to 6G.
+This concept aligns with the architecture of O-RAN.
 O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7].
-Within this context, the implementation of network slicing in O-RAN not only optimizes resource allocation but also fosters multi-vendor interoperability and innovation.
-At the core of O-RAN’s architecture are two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control.
-These components collaborate seamlessly, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5].
+Implementing network slicing in O-RAN optimizes resource allocation and promotes multi-vendor interoperability and innovation.
+The core of O-RAN’s architecture consists of two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control.
+These components work together, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5].
 The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP).
@@
 Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
-Moreover, xAPPs leverage O-RAN network slicing for efficient resource utilization, allowing seamless collaboration across services.
-However, this introduces network management complexity, as each xAPP must adapt to multiple slice QoS needs, posing a critical coordination challenge.
-Additionally, cellular operators face the difficulty of establishing consistent connectivity for devices with different QoS demands [15, 16].
-Therefore, in this study, we examine the resource allocation problem using network slicing by considering the xAPPs.
-We propose a novel method for resource allocation, with the goal of effectively coordinating numerous independent xAPPs to address the network slicing and resource allocation challenges within O-RAN.
-The application of ML algorithms in wireless networks has surged due to the escalating complexity of these networks and the challenges posed by crafting custom optimization models.
-Traditional approaches, such as convex/non-convex optimization, are grappling with mounting complexity stemming from shifts in network architecture, evolving performance requirements, and the proliferation of devices.
-Present methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often neglect the capacity to learn resource distributions.
-Typically, these methods involve designing an agent for a specific task and training it from scratch, resulting in reduced exploration efficiency and extended convergence times.
-Moreover, they necessitate a substantial volume of training samples, leading to prolonged exploration phases.
-The limited adaptability of traditional ML methods to novel scenarios motivates us to explore an alternative approach capable of superior generalization, more efficient data utilization, and resource distribution learning.
-To address these objectives, we propose a novel methodology that employs a self-supervised learning technique, specifically contrastive learning [17], in conjunction with a variational autoencoder (VAE) model, known as a type of generative model [18].
-These models excel in handling uncertainty by learning potential outcome distributions, thereby offering a distribution of predictions instead of a singular deterministic output.
-Moreover, their adeptness in handling incomplete or missing data allows for the imputation of missed values based on learned data patterns.
-VAEs also exhibit proficiency in semi-supervised learning setups, effectively harnessing both labeled and unlabeled data to enhance predictive accuracy.
-Their generative abilities facilitate the creation of new data instances resembling the training set, proving beneficial in scenarios demanding synthetic data generation or heightened resilience against noise and outliers.
-Furthermore, contrastive learning involves contrasting different examples to acquire informative data representations.
-On the other hand, a VAE comprises an encoder and decoder that jointly learn the underlying probability distribution of data.
-By integrating contrastive learning with VAE, we can mitigate the common challenge of requiring a large number of training samples in ML.
-This approach can enhance performance metrics, such as quicker convergence and improved results, while also reducing the time and effort needed for algorithm design and training.
+Using xAPPs leverages O-RAN network slicing to use resources efficiently and facilitate collaboration across services.
+However, this also introduces complexities in network management, as each xAPP needs to adjust the different QoS requirements of multiple services, posing a coordination challenge.
+Furthermore, cellular operators face the difficulty of establishing consistent connectivity for devices with varying QoS demands [15, 16].
+Therefore, this study addresses the resource allocation problem within O-RAN using network slicing by considering the xAPPs.
+We propose a novel resource allocation method to effectively coordinate the numerous independent xAPPs to address network slicing and resource allocation challenges.
+The use of ML algorithms in wireless networks has increased due to the growing complexity of these networks and the challenges of creating custom optimization models.
+Traditional approaches, such as convex/non-convex optimization, are struggling with the increasing complexity stemming from changes in network architecture, evolving performance requirements, and the growing number of devices.
+Current methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often overlook the ability to learn resource distributions.
+Typically, these methods involve designing an agent for a specific task and training it from scratch, which reduces exploration efficiency and extends convergence times.
+Moreover, they require many training samples, leading to prolonged exploration phases.
+The limited adaptability of traditional ML methods to new scenarios motivates us to explore an alternative approach capable of better generalization, more efficient data utilization, and resource distribution learning.
+To address these objectives, we propose a novel methodology that uses a self-supervised learning technique, specifically contrastive learning [17], in combination with a variational autoencoder (VAE) model, which is a type of generative model [18].
+These models excel in handling uncertainty by learning distributions, offering a distribution for predictions instead of a singular deterministic output.
+Moreover, their skill in handling incomplete or missing data allows them to fill in missing values based on learned patterns.
+VAEs also excel in semi-supervised learning setups, effectively using labeled and unlabeled data to improve predictive accuracy.
+Their generative abilities enable the creation of new data instances that resemble the training set, which is helpful in scenarios requiring synthetic data generation or increased resilience against noise and outliers.
+Furthermore, contrastive learning involves comparing different examples to acquire informative data representations.
+On the other hand, a VAE consists of an encoder and decoder that work together to learn the underlying probability distribution of data.
+By integrating contrastive learning with VAE, we can address the challenge of needing many training samples in ML.
+This approach can improve performance metrics, leading to faster convergence and better results while reducing the time and effort required for algorithm design and training.
+In our proposed solution, x-Apps help to implement advanced resource allocation algorithms within the O-RAN framework.
+The algorithm is deployed as an x-App on the near-real-time RIC, enabling efficient and dynamic resource allocation adjustment based on real-time network conditions and user demands.
 I-A Contribution
-This paper introduces a system model within the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC, through resource allocation, with a particular emphasis on network slicing.
-Our paper’s primary contributions can be summarized as follows:
-  • Our research primarily aims to tackle resource allocation challenges in the O-RAN architecture, with a specific focus on optimizing weighted throughput by allocating resources like transmission power and physical resource blocks.
-We account for the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC.
+This paper introduces a system model in the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC.
+It focuses on resource allocation, with a particular emphasis on network slicing.
+The primary contributions of our paper can be summarized as follows:
+  • Our research aims to address the resource allocation problem in the O-RAN architecture, explicitly focusing on optimizing weighted throughput by allocating resources such as transmission power and Physical Resource Blocks.
+We consider the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC.
 Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
-  • Our solution to the resource allocation optimization problem employs a two-step algorithm.
-Firstly, we simplify the problem by framing it as an optimization task.
-Secondly, we employ a two-phase training approach.
-We use supervised learning with a VAE to train power transmission regression and resource allocation decisions.
-This is followed by unsupervised learning with a contrastive loss method to enhance the model’s generalization and robustness.
-In essence, our approach simultaneously handles power and radio resource allocation xAPPs.
-In addition, it is not limited to O-RAN but can be applied to future multi-vendor O-RAN architectures, leveraging team learning principles developed by our team.
+  • Our approach to optimizing resource allocation involves two main steps.
+Firstly, we simplify the problem by framing it as an optimization problem.
+Then, we introduce a novel machine learning algorithm that combines supervised learning with a VAE and unsupervised learning with contrastive loss to optimize resource allocation in O-RAN.
+It significantly improves the model’s generalization ability, making it a more practical solution for real-world applications.
+  • While explicitly guaranteeing a global optimal solution is difficult, our method aims to achieve near-optimal solutions through evaluation with multiple benchmarks and significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems.
+The method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments.
+It is not limited to O-RAN and can be applied to future multi-vendor O-RAN architectures, leveraging semi-supervised generative model principles developed by our team.
 I-B Structure of Paper
-The paper is structured as follows: Section II presents a review of related works.
+The paper is structured as follows: Section II reviews related works.
 In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem.
 Section IV provides a detailed description of the proposed resource allocation algorithm.
-In Section V, we present numerical results used to evaluate the algorithm’s performance.
-Finally, Section VI offers our conclusion.
+Section V presents numerical results used to evaluate the algorithm’s performance.
+Section VI compares the computational cost of employed algorithms.
+Finally, Section VII offers our conclusion.
 II Related Works
 As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements.
-This challenge becomes more pronounced in 5G networks due to their increased complexity and stringent QoS demands [16].
-Recently, there has been a growing focus on addressing the network slicing problem in cellular networks, as evidenced by several studies [16, 19, 20, 21, 22].
+This challenge becomes more prominent in 5G and 6G networks due to their increased complexity and stringent QoS demands [16].
+Recently, a growing focus has been on addressing the network slicing problem in cellular networks [16, 19, 20, 21, 22].
 Feng et al.
 [16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements.
-However, their simulation utilized only a limited number of User Equipment (UE) devices and service types, which may not accurately represent real-world 5G networks with diverse service requirements.
-In another study, a resource allocation approach based on the M/M/1 queuing system was presented for RAN slicing [19].
-Nevertheless, practical scenarios often involve bursty traffic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services, overly optimistic.
+Another study presented a resource allocation approach based on the M/M/1 queuing system for RAN slicing [19].
+Nevertheless, practical scenarios often involve bursty traffic, overly optimistic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services.
 Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process.
-Additionally, in a study by Kazemifard et al.
-[20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center.
-The authors conducted a theoretical analysis and employed a mathematical model to efficiently address minimum delay problems.
+Additionally, in a study by Kazemifard et al., [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center.
+The authors conducted a theoretical analysis and employed a mathematical model to address minimum delay problems efficiently.
 Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23].
 ML algorithms, in particular, have found application in power control and resource allocation within wireless networks.
-These models are trained on historical resource allocation decisions and then applied to make new allocation decisions.
 For instance, the SAMUS framework was employed by Bektas et al.
@@
 Similarly, Samarakoon et al.
-[30] introduced a federated learning algorithm to optimize transmit power and radio resource allocation in vehicular networks.
+[30] introduced a federated learning algorithm to optimize power transmission and radio resource allocation in vehicular networks.
 Additionally, Elsayed et al.
@@
 III System Model and Problem Formulation
-In our study, we address network slicing and resource allocation problem within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice.
-The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery, while eMBB slices focus on high data rates for mobile broadband.
-These slices have also various QoS requirements for network resource allocation [10].
-By slicing the RAN into dedicated components for URLLC and eMBB, we can tailor network optimization to distinct traffic types and services, enhancing the efficient and flexible utilization of network resources.
-Key resources in O-RAN allocation encompass transmission power and Physical Resource Blockss.
+In our study, we address the resource allocation problem using network slicing within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice.
+The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery.
+In contrast, eMBB slices focus on high data rates for mobile broadband.
+Both slices have different QoS requirements for network resource allocation [10].
+Slicing the RAN into separate components for URLLC and eMBB optimizes the network for different traffic types and services, improving the efficient and flexible utilization of network resources.
+Key resources in O-RAN allocation include transmission power and PRBs.
 Traffic demands, often quantified by data rate and bandwidth, reflect the volume of data requiring transmission through the network.
-QoS requirements delineate the desired network performance based on parameters such as data rate.
-Our system model for resource allocation with network slicing is illustrated in Figure 1.
+QoS requirements define the desired network performance based on factors such as data rate.
+The system model for our resource allocation problem with network slicing is illustrated in Figure 1.
+Table I describes all variables and parameters used in this study.
 Figure 1: O-RAN network system model
-In our study, we investigate an O-RAN-based deployment, wherein each base station, known as RU in O-RAN, accommodates two distinct slices: eMBB and URLLC.
-These slices cater to UEs with comparable QoS requirements.
-Our analysis encompasses two critical network functions: power control and radio resource allocation, which are jointly considered.
-We denote the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$.
-Each base station $b$ serves two sets of users: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ users, where $U=U_{e,b}+U_{u,b}$.
-Each user is exclusively associated with a single RU.
-The temporal and spectral layout is partitioned into $\mathcal{M}$ segments referred to as PRBs.
-These PRBs constitute the smallest unit of resources for allocation [32].
-Each PRB spans a time frame, defined as a time slot in the temporal domain, and covers a frequency range of $f$ Hz in the spectral domain.
-We also denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices.
-The total number of slices ($S$) will be: $S=S_{e}+S_{u}$.
-The user association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if user $u$ is connected to RU $b$ in slice $s$, and 0 if not.
+In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC.
+These slices are designed to meet the QoS requirements of User Equipments.
+Our analysis focuses on two key network functions, power control, and radio resource allocation, which are jointly considered.
+We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$.
+Each BS $b$ serves two sets of UEs: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e,b}+U_{u,b}$.
+Each UE is exclusively associated with a single RU.
+The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [32].
+Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain.
+We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices.
+The total number of slices ($S$) will be $S=S_{e}+S_{u}$.
+The UE association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if UE $u$ is connected to RU $b$ in slice $s$, and 0 if not.
 Since each UE should be connected to only one RU, the following constraint must be met:
 $\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
-We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when user $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
+We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
 $\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
@@
 $\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
-The rate achieved for a user $u$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR).
+We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$.
+The rate achieved for a UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR).
 The calculation for $R_{u,b,m,s}$ is determined by:
@@
 The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively.
-The existence of quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, particularly when the original signal displays significant variations or spans a wide range of values.
+Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values.
 This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE.
@@
 $=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}),$ (9)
-where, we denote the transmission power of PRB $m$ in RU $b$ in slice s as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$.
-Furthermore, we define the channel gain between the resources and UE as $h_{ij}$.
-To represent this relationship in matrix form, we construct a matrix that accounts for $\mathcal{B}$ resources and $\mathcal{U}$ users as follows:
-$H:=\begin{bmatrix}|H_{11}|^{2}&\cdots&|H_{1U}|^{2}\\ \vdots&\vdots&\vdots\\ |H_{B1}|^{2}&\cdots&|H_{BU}|^{2}\end{bmatrix}_{B\times U}$ (13)
-The overall propagation delay ($D^{pro}$) comprises three distinct components: the propagation delay in the fronthaul link ($D^{fr,p}$), the midhaul link ($D^{mid,p}$), and the backhaul link ($D^{b,p}$).
-In each of these links, the propagation delay is determined by the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ signifies the length of the fiber link, and $C$ represents the propagation speed of the medium.
-Similarly, the total transmission delay ($D^{tr}$) is the aggregate of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links.
-Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, represented as $D=\frac{\alpha}{R}$, where $R$ indicates the data rate of the packet, and $\alpha$ denotes the mean packet size.
-$D_{total}=D^{pro}+D^{tr},$ (14)
-$D_{u,s}=\frac{L}{C}+\frac{\alpha}{R_{u,s}}.$ (15)
-Therefore, the delay of UE $u$ in slice $s$, as derived from Equation 15, should be constrained to be below a specific threshold relevant to its respective slice, otulined below:
-$D_{u,s}\leq D^{min}_{u,s}.$ (16)
-The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring compliance with QoS requirements.
-In this context, $w_{s}$ signifies the priority associated with individual service types and UEs.
+The channel gain matrices in our model are generated using a Rayleigh fading process to simulate variable channel conditions typical in non-line-of-sight environments.
+The channel vector from the PRB $m$ in RU $b$ to the $u$ UE in the $s$ slice, denoted by $\mathbf{h}_{u,b,m,s}$, is modeled as $\mathbf{h}_{u,b,m,s}=\sqrt{\beta_{u,b,m,s}}\mathbf{g}_{u,b,m,s}$, where $\beta_{u,b,m,s}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u,b,m,s}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable.
+To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
+The total delay experienced by each UE is the sum of propagation and transmission delays.
+The overall propagation delay ($D^{pro}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{b,p}$) links.
+In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ is the link length and $C$ is the propagation speed.
+Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links.
+Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u,s}}$, where $R_{u,s}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size.
+The total delay for UE $u$ in slice $s$ is $D_{u,s}=\frac{L}{C}+\frac{\mu}{R_{u,s}}$, and should be below a specific threshold ($D^{max}_{u,s}$), expressed as $D_{u,s}\leq D^{max}_{u,s}$.
+The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring satisfying the QoS requirements.
+In this context, $w_{s}$ indicates the priority associated with individual service types.
 The intelligent RU is responsible for effectively managing the requirements of both network slices.
@@
 $p_{u,b,m,s}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (13)
-$R_{u,s}\geq R_{s}^{max}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (21)
+$R_{u,s}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
 $C_{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (15)
-$D_{u,s}\leq D^{min}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (23)
+$D_{u,s}\leq D^{max}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (16)
 $\alpha_{u,b,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (17)
@@
 $\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (21)
+This formulation can be seen as a well-known NP-hard problem, a multi-dimensional knapsack problem.
+Since we can reduce it to our resource allocation problem in polynomial time, it follows that it is also NP-hard.
+Given the NP-hard nature of the resource allocation problem, finding an exact solution in polynomial time is computationally infeasible for large-scale networks.
+The system model is constrained by limited resources such as UE or RU and slice power and fronthaul capacity, preventing the aggregate throughput from exceeding its optimal value.
+This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely.
+As a result, the algorithm used to solve the system will converge to an optimal solution.
+If the objective function is a strictly ascending function concerning the number of iterations, it will converge to its global optimum.
+The algorithm will converge to a local optimum if it is a non-monotonically ascending function.
+This guarantees the stability and convergence of the system model within the feasible region.
+Notation Definition
+$b,\mathcal{B},B$ Index, set, and the number of RUs
+$u,\mathcal{U},U$ Index, set, and the number of UEs
+$m,\mathcal{M},M$ Index, set, and the number of PRBs per RU
+$s,\mathcal{S},S$ Index, set, and the number of slices
+$h,H$ Index, and channel gain matrix between RUs and UEs
+$\alpha_{u,b,s}$ UE $u$ association to RU $b$ in slice $s$
+$\beta_{u,b,m,s}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
+$p_{u,b,m,s}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
+$R_{u,b,m,s}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
+$R_{u,s}$ Rate achieved for UE $u$ in slice $s$
+$R_{u}$ Rate achieved for UE $u$
+$w_{s}$ Priority weights for each service type
+$P_{b}$ Transmitted power RU in the fronthaul link
+$C_{b}$ Data rate in the fronthaul link
+$\sigma_{1}^{2}$ Quantization Noise
+$D^{pro}$ Overall propagation delay
+$D^{fr,p}$ Propagation delay in the fronthaul link
+$D^{mid,p}$ Propagation delay in the midhaul link
+$D^{b,p}$ Propagation delay in the backhaul link
+$L$ Length of the fiber link
+$C$ Propagation speed of the medium
+$D^{tr}$ Total transmission delay
+$D^{fr,t}$ Transmission delay in the fronthaul link
+$D^{mid,t}$ Transmission delay in the midhaul link
+$D^{b,t}$ Transmission delay in the backhaul link
+$R$ Data rate of the packet
+$\mu$ Mean packet size
+TABLE I: Notations with definitions for the system model parameters.
 IV Methodology
-Our primary focus is to address resource allocation problem within O-RAN architecture, and we propose a novel approach based on semi-supervised learning.
-Our objective is to optimize the sum of rates among UEs while considering constraints, including transmission power and PRB allocation.
-To accomplish this, we compile a dataset comprising diverse resource allocation scenarios, encompassing various combinations of objective functions and constraints.
-We generate this dataset using an exhaustive search algorithm (ESA), which explores the entire feasible solution space.
-Our research primarily delves into the resource allocation problem within the O-RAN architecture, and we introduce an innovative approach based on semi-supervised learning.
-Our primary goal is to enhance the cumulative rates for UEs while adhering to specific constraints such as transmission power, delay, and PRB allocation.
-To achieve this, we curate a comprehensive dataset that encompasses a wide range of resource allocation scenarios, covering diverse combinations of objective functions and constraints.
-This dataset is generated through the implementation of an ESA, which thoroughly explores the entire feasible solution space.
-In our approach to addressing the resource allocation problem, we employ a deep neural network (DNN) model integrated with a contrastive loss function.
-This DNN model efficiently manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions.
-Through the training of our model on the dataset we’ve created, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem.
+Our research addresses the resource allocation problem within the O-RAN architecture by introducing a semi-supervised learning algorithm.
+To achieve this, we have compiled a comprehensive dataset that includes various resource allocation scenarios, covering different objective functions and constraint combinations.
+We have created labeled data (computed $\alpha$, $\beta$, and transmission powers) using exhaustive search algorithm (ESA), thoroughly exploring the feasible solution space to find the global optimum solution as input for our algorithm’s supervised learning part and as a benchmark for models evaluation.
+We use a deep neural network (DNN) model with a contrastive loss function to solve the resource allocation problem.
+This DNN model effectively manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions.
+By training our model, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem (QoS).
 This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
-We aim to evaluate the performance of our learning-based approach using multiple regression-based performance metrics to assess its robustness.
-Additionally, we will validate the effectiveness of our model by comparing its outcomes with those of the ESA in different scenarios.
-Moreover, while the ESA is the optimal algorithm for resource allocation decisions, its complexity makes it impractical for deployment in real-world environments.
-Therefore, we will also compare the performance of our proposed algorithm with state-of-the-art methods known for their practical deployment capabilities, such as DRL.
-This comprehensive assessment will demonstrate the method’s reliability across diverse scenarios.
-Figure 2 illustrates the overview of our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
-Figure 2: The overview of proposed methodology.
+We aim to evaluate the performance of our SS-VAE method using multiple regression-based performance metrics to ensure its robustness.
+We will also validate the effectiveness of our model by comparing its results with those of the ESA across various scenarios.
+While ESA is an algorithm for resource allocation decisions that find global optimum solutions, its complexity makes it impractical for real-world deployment.
+Therefore, we will also compare its performance with state-of-the-art methods known for their practical deployment capabilities, such as DRL.
+This thorough evaluation will demonstrate the reliability of our method across different scenarios.
+Figure 2 illustrates our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
+Figure 2: Overview of the proposed methodology.
 IV-A Exhaustive Search Algorithm
-The ESA is a well-established optimization approach for solving resource allocation problems in wireless networks.
-Its fundamental concept involves generating all possible resource allocation combinations and selecting the best one based on a predefined objective function [33].
-ESA serves as a valuable benchmark solution, providing a ground truth against which results from other methods can be validated and compared.
-Its advantages include simplicity of understanding and implementation, along with the assurance of yielding an optimal solution within a finite time frame.
-Nonetheless, ESA’s primary limitation lies in its computational complexity, especially when addressing resource allocation problems with extensive dimensions and intricate constraints.
-The computational burden grows exponentially with the size of the problem, rendering ESA impractical for large-scale resource allocation tasks [34].
-Despite this drawback, ESA remains a popular choice as a baseline for comparing various optimization algorithms and offers a crucial reference for evaluating new optimization strategies.
-To assess the performance, validity, and robustness of the proposed learning-based method, we will compare its results with those generated by ESA.
-Additionally, a subset of solutions for the resource allocation and network slicing optimization problem within O-RAN will be generated using ESA.
-This labeled dataset will be utilized to train the supervised learning component of the proposed method.
+The ESA is a fundamental approach used to find the global optimal solution to resource allocation problems by evaluating all possible configurations and selecting the best one based on a predefined objective function [33].
+The ESA is a valuable benchmark solution that provides a ground truth against comparing results from other methods.
+Its advantages include simplicity of understanding and implementation and the assurance of yielding an optimal solution within a finite time frame.
+However, its guarantee of optimality comes at the cost of high computational effort, making it suitable primarily for small or well-defined problems where the solution space can be thoroughly explored.
+Nonetheless, ESA’s main limitation lies in its computational complexity, especially when addressing resource allocation problems with high dimensions and complex constraints.
+The computational cost increases exponentially as the problem size grows, making ESA impractical for large-scale resource allocation tasks [34].
+Despite this limitation, ESA is still widely used as a benchmark for comparing different optimization algorithms and as a reference for evaluating new optimization techniques.
+To evaluate the performance, accuracy, and reliability of the SS-VAE method, we will compare its results with those obtained by ESA.
+We will use the ESA to obtain a subset of solutions for the objective variables ($\alpha$, $\beta$, $p$) in the resource allocation optimization problem with network slicing.
+This labeled dataset will be used to train the supervised learning part of our SS-VAE method.
+The detailed steps of ESA are described in Algorithm 1.
+Algorithm 1 ESA for Resource Allocation in O-RAN
+1:$B$, $U$, $S$, $M$, $w_{s}$ , $P^{max}_{b}$, $P^{max}_{s}$, $R^{min}_{s}$, $C_{b}^{max}$, $D_{u,s}^{min}$, and $H$
+2:Optimal resource allocation $best\_solution$ and its objective value $best\_value$
+3:Initialize dataset $\mathcal{X}\leftarrow\emptyset$
+4:$best\_value\leftarrow-\infty$
+5:$best\_solution\leftarrow\emptyset$
+6:for each possible combination of $(\alpha,\beta,p)$ do
+7:     Calculate $R_{u,s}$ using Eq. (III)
+8:     if all constraints are satisfied then
+9:         $current\_value\leftarrow\sum_{s\in S}\sum_{u\in U}w_{s}R_{u,s}$
+10:         if $current\_value>best\_value$ then
+11:              $best\_value\leftarrow current\_value$
+12:              $best\_solution\leftarrow(\alpha,\beta,p)$
+13:              Add $(\alpha,\beta,p,R_{u,s})$ to dataset $\mathcal{X}$
+14:         end if
+15:     end if
+16:end for
+17:return $best\_solution$, $best\_value$, $\mathcal{X}$
 IV-B Deep Reinforcement learning Algorithm
-As discussed in Section II, DRL algorithms have proven to be effective solutions for addressing resource allocation problems.
-These algorithms have been successfully deployed in various real-world wireless network environments, demonstrating their practicality and their ability to significantly improve both technical and business metrics [35].
-Consequently, we utilize DRL, where an agent interacts with the environment to make resource allocation decisions and receive related rewards.
-These rewards are tied to the objective function outlined in Problem 17, considering its objectives and constraints.
-Next, we model the problem as a Markov Decision Process (MDP) and introduce the DRL algorithm to address the resource allocation challenge in network slicing.
-This approach serves as an extra means to evaluate the robustness and our primary proposed solution.
-IV-B1 Markov Decision Process
-In the specified resource allocation problem (Equation 17), an O-RAN orchestrator is responsible for determining PRB allocation decisions and determine the corresponding RU for each UE for entw.
-This problem can be represented as a MDP denoted by $(\mathfrak{S},\mathfrak{A},\mathfrak{P},\mathfrak{R})$, where $\mathfrak{S}$ represents the state space, $\mathfrak{A}$ is the action space, $\mathfrak{P}(\mathfrak{s}^{\prime}|\mathfrak{s},\mathfrak{a})$ signifies the transition probability from state $\mathfrak{s}$ to state $\mathfrak{s^{\prime}}$ by taking action $\mathfrak{a}$, and $\mathfrak{P}$ represents the reward function [35, 36, 37, 38], which are elaborated for our specific problem.
+As discussed in Section II, DRL algorithms effectively address resource allocation, proving their practicality in wireless networks and enhancing key metrics [35].
+In this study, we use DRL, where an agent allocates resources based on rewards linked to the objective in Problem 10.
+Hence, we model the problem as a Markov Decision Process (MDP) and use a DRL algorithm as a benchmark for evaluating our SS-VAE method.
+In the resource allocation problem (Equation 10), an O-RAN orchestrator is responsible for determining the PRB allocation decisions and identifying the corresponding RU for each UE within each slice.
+This problem can be represented as a MDP [35, 36, 37, 38].
   1. I.
     State: The state at time $t$ ( $\mathfrak{s}(t)\in\mathfrak{S}$), is represented by $\{\mathfrak{s}_{u}(t)\}_{t=1}^{N}$, where $\mathfrak{s}_{u}(t)$ indicates the state of UE $u$ ($u=1:U$).
-Here, $\mathfrak{s}_{u}(t)$ is a binary indicator: it equals 1 if the data rate requirement of UE $u$ is met and 0 otherwise.
-It’s important to note that each data rate corresponds to a specific transmission power level.
+Here, $\mathfrak{s}_{u}(t)$ is a binary indicator: it equals one if the data rate requirement of UE $u$ is satisfied and 0 otherwise.
+Each data rate also corresponds to a quantized transmission power level.
   2. II.
@@
   4. IV.
-    Reward: The reward function serves as a pivotal element in assessing the gains or expenses associated with a given state-action pair.
-It provides guidance to the agent’s decision-making process, indicating the appeal of different actions across various states.
-In alignment with earlier studies [35, 36, 37], our method characterizes the reward function as a combination of the objective function and the constraints relevant to the considered problem (17).
-Therefore, we establish the reward function for our specific problem as follows:
+    Reward: The reward function evaluates the gains or costs of a given state-action pair.
+It guides the agent’s decision-making process by indicating the desirability of various actions across different states.
+Consistent with previous studies [35, 36, 37], our approach defines the reward function as a combination of the objective function and the constraints pertinent to the considered problem (Equation 10).
+Thus, we specify the reward function for our specific scenario as follows:
     $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u,s}+\Theta_{const.}\;C_{u,b,m,s}+\Theta_{bias},$ (22)
     where, $\Theta_{r}$, $\Theta_{const.}$, and $\Theta_{bias}$ represent the respective weights assigned to the objective function (Data Rate), constraints, and the bias value.
-The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP.
-This policy maps any state $\mathfrak{s}\in\mathfrak{S}$ to a probability distribution $\pi(\cdot|\mathfrak{s})$ over the set of possible actions $\mathfrak{A}$.
-The goal is to find a policy that maximizes the cumulative reward obtained through dynamic learning from acquired data.
-The policy $\pi$ can be influenced and controlled by decision makers.
+The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP, that maximizes the cumulative reward obtained through dynamic learning from acquired data.
 The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state.
-Additionally, the action value function $Q^{\pi}(\mathfrak{s},\mathfrak{a})$ is introduced to represent the expected cumulative reward when taking action $a$ in state $\mathfrak{s}$, considering that the policy $\pi$ is being followed.
-This can be expressed as follows:
-$V^{\pi}(\mathfrak{s})=\mathbb{E}[Q^{\pi}(\mathfrak{s},\mathfrak{A})|\mathfrak{A}\sim\pi(.|\mathfrak{s})],$ (30)
-$Q^{\pi}(\mathfrak{s},\mathfrak{a})$ $=\mathbb{E}[\sum_{t=0}^{\inf}\gamma^{t}\cdot\mathfrak{R_{t}}|\mathfrak{S_{0}}=\mathfrak{s},\mathfrak{A_{0}}=\mathfrak{a}]$
-$=\mathfrak{r}(\mathfrak{s},\mathfrak{a})+\gamma\;\mathbb{E}[V^{\pi}(\mathfrak{S}^{\prime})|\mathfrak{S}^{\prime}\sim\mathfrak{P}(.|\mathfrak{s,a})],$ (31)
-where, $\mathfrak{R}_{t}$ represents the immediate reward obtained at time $t$, and $\gamma$ (where $10<\gamma<$ 1) denotes the discount factor applied in the computation of future rewards.
-The discount factor determines the weight of future rewards in the decision-making process.
 Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task.
 Determining the value function in such scenarios presents a significant challenge.
-To address MDPs, two primary strategies are employed: model-based and model-free approaches.
-In instances where there is uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
-IV-B2 Deep-Q-Network Algorithm
-In our considered problem, we face the challenge of managing the states space ($\mathfrak{S}$) and actions space ($\mathfrak{A}$) that are inherently high-dimensional and complex.
-This complexity arises from continuous variable like transmission power, as well as the substantial number of UEs and PRB resources.
-Traditional Q-learning methods struggle in such scenarios, as many state-action pairs are rarely encountered, leading to inefficient learning processes [35, 38].
-To overcome this challenge, we turn to the Deep Q-Network (DQN) approach to tackle our resource allocation problem and optimize parameters such as transmission power, alpha, and beta.
-DQN offers distinct advantages in handling these complex, high-dimensional state-action spaces by utilizing a deep neural network to approximate the $Q$-function.
-DQN’s capacity to generalize from prior experiences and its use of techniques such as experience replay and target networks make it an apt choice for addressing resource allocation challenges within our problem domain.
-DQN employs experience replay to store into replay memory ($\mathbb{M}$) and sample transitions from the memory ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$) to train the neural network with stochastic gradient descent.
-This method helps acquire uncorrelated samples from a temporally correlated MDP trajectory, enhancing gradient estimation accuracy.
-A target network $Q_{\theta^{*}}$ is also used, where independent samples are obtained from the replay memory.
-These samples are used to update the $Q$-network’s parameter, improving learning stability and efficiency [38].
-We compute the target value and update $\theta$ by the gradient of equations 32 33.
-$Y=\mathfrak{R(s,a)}+\gamma\;\max_{\mathfrak{a}^{\prime}\in\mathfrak{A}}Q_{\theta^{*}}(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime}),$ (32)
-$L(\theta)=||Y-Q_{\theta}(\mathfrak{s},\mathfrak{a}||^{2},$ (33)
-The parameter $\theta^{*}$ is updated every $T_{target}$ steps by setting it equal to $\theta$.
-The target network remains unchanged for $T_{target}$ steps and is then updated using the current weights of the Q-network.
+Two primary strategies are employed to address MDPs: model-based and model-free approaches.
+In uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
+Our resource allocation problem involves managing high-dimensional and complex state and action spaces due to continuous variables like transmission power and numerous UEs and PRB resources.
+Traditional Q-learning methods struggle in this context because of inefficient learning with rarely encountered state-action pairs [35, 38].
+To solve the problem of finding parameters such as transmission power ($p$), $\alpha$, and $\beta$, we use the Deep Q-Network (DQN) approach, which handles complex state-action spaces by approximating the $Q$ function with a deep neural network.
+DQN benefits from experience replay and target networks, enhancing learning stability and efficiency [38].
+Experience replay stores transition in memory ($\mathbb{M}$) and samples them to train the neural network ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$), reducing correlation in the data and improving gradient estimation.
+A target network stabilizes learning by updating the $Q$-network’s parameters with independent samples [38].
 Our DQN training employs the epsilon-greedy policy for action selection.
@@
 $\pi(S)=\begin{cases}\text{argmax}_{\mathfrak{a}}Q(\mathfrak{S,a})&\text{with probability }1-\epsilon\\ \text{random action}&\text{with probability }\epsilon\end{cases}$
-It’s worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
-IV-C Proposed Method
-In recent times, DL techniques have found application in enhancing the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems, as demonstrated in several studies [39, 40, 41, 42].
-These approaches excel in learning complex, non-linear relationships between input features like channel matrices and output decisions, such as power and resource allocation.
-What distinguishes DL-based resource allocation is its ability to effectively manage the dynamic and intricate nature of wireless communication systems.
-These approaches can adapt to varying channel conditions and UE, resulting in improved network performance compared to traditional optimization methods.
-In our proposed DL-based resource allocation method, we feed the channel matrix into a DNN.
-This network is responsible for learning optimal resource allocation decisions that adhere to network constraints while maximizing the defined objective function (weighted aggregated data rate).
-We also employ a flattened vector with dimensions $UB$, derived from the channel gain matrix, as input for a machine-learning model, which includes a VAE.
-This autoencoder utilizes the channel matrix to generate the power vector.
-The key innovation in our approach is the introduction of contrastive loss as an additional term in the VAE’s loss function.
-This addition refines the learning of representations, significantly enhancing the accuracy of solutions to resource allocation problems.
-Subsequent sections will provide a detailed exploration of each component in our proposed method.
+It is worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
+IV-C Proposed Method (SS-VAE)
+Traditional resource allocation methods in wireless communication systems, such as heuristic and meta-heuristic algorithms, often rely on predefined rules or optimization algorithms that do not adapt well to changing network conditions.
+In recent times, DL techniques have been used to improve the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems [39, 40, 41, 42].
+These approaches are practical in learning complex, non-linear relationships between input features (e.g., channel matrices) and output decisions (e.g., power and resource allocation).
+The distinctive feature of DL-based resource allocation is its capacity to manage wireless communication systems’ dynamic and complex nature effectively.
+DL-based approaches can adjust to changes in network conditions, user demands, and interference patterns, providing more flexible and efficient resource management than traditional methods.
+Traditional methods often face challenges in real-time adaptation due to predefined rules, complexity in optimization, and a lack of real-time learning.
+Our SS-VAE method inputs the channel gain matrix into a DNN.
+This network learns optimal resource allocation decisions that satisfy network constraints and maximize the weighted aggregated data rate.
+We use a flattened vector, with dimensions $UB$, derived from the channel gain matrix as the input for the machine learning model.
+Our main contribution involves integrating a VAE with contrastive loss to solve the resource allocation problem.
+By adding contrastive loss into the VAE‘s loss function, we enhance the encoder’s ability to learn more refined representations, significantly improving the accuracy of resource allocation decisions.
+The SS-VAE method allows the encoder layers of the VAE to be explicitly utilized for the resource allocation task.
+In our problem, the connection between the xApps that handle power transmissions and PRB allocation emerges from their shared foundation within the VAE framework.
+The VAE operates as a shared space where both xApps come together, enabling a collaborative exploration and consideration of the relevant features for both power transmission and PRB allocations.
+The xApps access a common latent representation of the channel gain matrices through this shared space.
+This shared understanding facilitates a cohesive analysis where the relationships between power transmissions and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the key features that impact both domains.
+This integration of information helps make more comprehensive and informed decisions about resource allocation, as it ensures that both xApps use a unified understanding of the underlying data to optimize their respective objectives.
+This coordination allows for a synchronized approach to resource allocation within the O-RAN architecture.
 IV-C1 Variational Autoencoder
-The VAE is a specific type of generative model frequently applied in unsupervised learning tasks.
-It functions as a DNN designed to generate new data samples resembling a provided set of training data.
-Utilizing a VAE confers a notable advantage: it adeptly captures the intrinsic distribution of the channel gain matrix.
-This acquired knowledge proves immensely valuable when confronting the resource allocation problem [43].
-In a VAE, let’s denote the input data as $x$, a latent variable representing a lower-dimensional data representation as $z$, and $p_{\theta}(x)$ as a generative model representing the data distribution based on the latent variable.
-The primary goal of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$.
-This likelihood can be mathematically expressed as:
+The VAE is a type of DNN generative model used in unsupervised learning tasks.
+It is designed to generate new data samples that resemble a provided set of training data.
+One notable advantage of using a VAE is its ability to adeptly capture the intrinsic distribution of the channel gain matrix (input data), which is valuable when addressing the resource allocation problem [43].
+In the VAE part of our method, the input data (flattened channel gain matrix) is represented by $x$, while a lower-dimensional data representation (latent space) is denoted as $z$.
+The generative model based on the latent variable is represented as $p_{\theta}(x)$.
+The main objective of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$.
+This likelihood can be expressed mathematically as:
 $\log p_{\theta}(x)=\mathbb{E}{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{KL}(q_{\phi}(z|x)|p(z)),$ (23)
-where, $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable given the input data.
-The term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable.
-The objective function can be enhanced by adjusting the parameters of both the generative model, denoted as $\theta$, and the encoder network, represented as $\phi$.
-This optimization process involves iteratively updating these parameters and computing the gradient of the objective concerning them, as outlined in [43].
-Figure 3 illustrates the VAE architecture tailored for addressing the resource allocation and network slicing problem.
-Figure 3: The architecture of proposed VAE.
-To provide a more detailed explanation, a VAE comprises two primary components: an encoder and a decoder.
-The encoder accepts the channel gain matrix as input and transforms it into a lower-dimensional representation known as a latent code.
-Subsequently, the decoder takes this latent code as input and produces an estimation of the transmission power.
-These two components work in tandem during training to minimize a reconstruction loss.
-This loss metric quantifies the disparity between the actual transmission power and the predicted transmission power, guiding the network to generate more accurate power predictions.
-Using a VAE for resource allocation offers several distinct advantages.
-Firstly, VAEs exhibit remarkable flexibility, enabling them to model intricate distributions.
-This characteristic is particularly advantageous for capturing the complex and non-linear relationship between channel gains and transmission powers.
-Secondly, VAEs excel at discerning the underlying structure within data, enabling them to make informed predictions even when confronted with previously unseen data [44].
-Thirdly, the latent code generated by the encoder serves as a compressed representation of channel gains, which can be leveraged for diverse tasks like visualization, clustering, and anomaly detection.
-Moreover, resource allocation problems are inherently non-convex and can be exceptionally challenging when using traditional optimization approaches.
-The introduction of VAEs provides a more versatile and effective means of tackling these intricate problems.
-Our dataset comprises a total of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples.
+where $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable based on the input data, the term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable.
+The loss function (Equation 23) can be optimized by adjusting the parameters of the generative model ($\theta$) and the encoder network represented ($\phi$).
+This optimization process involves iteratively updating these parameters and computing the gradient of the objective with respect to them [43].
+The architecture of the VAE for addressing the resource allocation problem with network slicing is illustrated Figure 3.
+The VAE comprises an encoder and a decoder.
+The encoder compresses the channel gain matrix into a lower-dimensional latent code while the decoder reconstructs the channel gain matrix from this code.
+During training, the VAE minimizes reconstruction loss, improving the accuracy of predicted transmission power.
+Figure 3: Architecture of the VAE.
+Several key advantages can be observed when using the VAE for resource allocation.
+The model can represent complex distributions, capture the non-linear relationship between channel gains and resource allocation decisions, and provide a compressed latent space helpful for visualization, clustering, and anomaly detection tasks.
+Moreover, resource allocation problems are inherently non-convex and challenging for traditional optimization methods, where the VAEs can propose a more versatile and practical solution to addressing these complexities.
+The dataset used to train the SS-VAE model consists of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples.
 These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$).
-It is important to note that the process of generating these input channel matrices is computationally less intensive compared to traditional methods, as documented in previous research [45].
+The labeled dataset results from the ESA applied to the same training dataset.
+Specifically, we have used power transmission and allocation decisions ($p,\alpha,\beta$) obtained by ESA as labels for training our SS-VAE model.
 To train the supervised learning component of our algorithm, we employ the ADAM optimizer [44].
-This optimizer facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples.
+It facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples.
 The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
-$L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||VAE(H_{i})-\Gamma_{i}||^{2}.$ (35)
-In our resource allocation challenge, the correlation between the xApps dealing with power estimation and the PRBs allocation emerges from their shared foundation within the VAE framework.
-The VAE operates as a unifying space where both xApps converge, enabling a synergistic exploration and consideration of the pertinent features relevant to both power transmission and PRB allocations.
-Through this shared space, the xApps gain access to a common latent representation of the channel gain matrices.
-This shared understanding facilitates a cohesive analysis where the interdependencies between power estimation and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the critical features that impact both domains.
-This integration of information aids in making more comprehensive and informed resource allocation decisions, as it ensures that both xApps leverage a unified understanding of the underlying data to optimize their respective objectives.
-This coordination is pivotal, as it allows for a coordinated and synchronized approach to resource allocation within the O-RAN architecture.
+$L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||encoder(VAE(H_{i}))-\Gamma_{i}||^{2}.$ (24)
+The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u,b,s}$ and $\beta_{u,b,m,s}$).
+Our model takes a flattened vector of dimension $UB$ as input, obtained from the channel gain matrix.
 IV-C2 Contrastive Loss
-Contrastive loss is a widely-used loss function in ML for various representation learning tasks, including image or text classification.
-Its fundamental aim is to construct a representation of data points that encourages similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17].
+Contrastive loss is a widely-used loss function in DL for various representation learning tasks, including image or text classification.
+Its fundamental aim is to construct a representation of data points that provokes similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17].
 Mathematically, the contrastive loss can be expressed as follows:
@@
 The parameter $m$ is a margin that separates the similar and dissimilar representations.
-To use contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices.
-One way to create channel gain matrices that are either similar or dissimilar is by using a statistical model of the channel that captures the relevant properties of the communication environment.
-The Rayleigh fading model is a commonly used statistical model for wireless channels, which assumes that the channel gain is a complex Gaussian random variable with zero mean and a variance that depends on the distance between the transmitter and receiver [47, 48].
-To create similar channel gain matrices, coefficients with similar magnitude and phase can be generated.
-Conversely, to create dissimilar channel gain matrices, coefficients with different magnitude and phase can be generated by introducing randomness into the channel model by using a Rayleigh fading model.
-To generate a channel gain matrix using this model, a matrix of complex Gaussian random variables with zero mean and a variance that depends on the distance between the transmitter and receiver can be generated as well as considering the interferance as noise [49, 50].
-The variance can be adjusted to create similar or dissimilar channel gain matrices based on the desired level of correlation between the channels.
-The mathematical formulation of this approach can be expressed as:
-$\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon},$ (37)
-where $\mathbf{H}_{noise}$ is the channel gain matrix with added noise, $\mathbf{H}$ is the original channel gain matrix, and $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and standard deviation $\sigma$.
-Once we have created similar and dissimilar channel gain matrices, we can train our contrastive loss function using these matrices as inputs. The loss function can be formulated as:
+To apply contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices.
+This can be achieved using a channel’s statistical model that reflects the relevant properties of the communication environment.
+By introducing randomness through a Rayleigh fading model, we can produce channel gain matrices with varying coefficients in magnitude and phase, thus creating either similar or dissimilar matrices.
+To generate a channel gain matrix using this approach, we create a matrix of complex Gaussian random variables with zero mean, where the variance is influenced by the distance between the transmitter and receiver while also accounting for interference as noise [45, 46].
+The variance can be adjusted to control the level of correlation between the channels, thereby generating similar or dissimilar channel gain matrices as needed.
+We can train our model with contrastive loss function once we have created similar and dissimilar channel gain matrices.
+The loss function can be formulated as:
 $L_{Contrastive}(\theta)=\frac{1}{\beta_{L}}$
 $\sum_{i\in\beta_{L}}[-\log\frac{\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{i}})}{\tau}}}{\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{i}})}{\tau}}+\sum_{i\neq j}\exp^{\frac{f(\bar{H_{i}})^{T}\times f(\underline{H_{j}})}{\tau}}}],$ (26)
-where $\theta$ is the set of parameters in the variational autoencoder model, $\tau$ is a temperature hyperparameter and $L(\theta)$ is the contrastive loss function.
-$\underline{H_{i}}$ and $\bar{H_{i}}$ are the similar matrices to the the main channel gain matrix ($H_{i}$), which allow us to train our proposed model [51].
-By minimizing this loss function, we can learn a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices and produce a power control decision that optimizes the network performance.
-IV-C3 Training the proposed model
-In this section, we will describe the training phase of our proposed method, which consists of two parts: the supervised learning part, where the VAE is trained to regress the power transmission, and the unsupervised learning part, where the contrastive loss is used to improve the generalization and robustness of our proposed method.
-There are several key parameters that should be taken into consideration in training the supervised part.
+where $\theta$ is the set of parameters in the VAE model, $\tau$ is a temperature hyperparameter, and $L(\theta)$ is the contrastive loss function.
+$\underline{H_{i}}$ and $\bar{H_{i}}$ are similar matrices to the main channel gain matrix ($H_{i}$) and are used in training the SS-VAE model [47].
+By minimizing this loss function, we can find a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices; therefore, it is used to solve the resource allocation problem.
+The connection between the VAE and the additional contrastive loss is essential for an optimal solution.
+The initial step establishes a strong base by using supervised learning to acquire a good representation of the resource allocation problem.
+The subsequent step further improves these representations, enhancing the model’s capacity to generalize and effectively handle new, unforeseen scenarios.
+IV-C3 Training the SS-VAE model
+The training phase of SS-VAE method consists of two parts.
+First, supervised learning involves training the VAE and using the VAE’s Encoder to predict power transmission.
+The second part is unsupervised learning, which uses contrastive loss to improve the generalization and robustness of our method.
 The latent space dimension represents the size of the hidden representation learned by the VAE [48].
-A larger latent space dimension can lead to a more expressive model, but also increases the risk of overfitting.
+A larger latent space dimension can lead to a more expressive model but also increases the risk of overfitting.
 The learning rate determines the speed at which the model learns from the data.
@@
 $L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [49].
-The proposed model is developed using the PyTorch [53] software framework.
-Both validation size and test size are $20\%$ of the dataset, allowing for a more thorough evaluation of the proposed model’s performance.
-Hyper-parameter tuning yields the value of hyperparameters, which can be found in the Table III.
-The initial model was pre-trained through self-supervised learning, using unlabeled channel gain matrices that corresponded to the stated loss function.
-The pre-training was based on regression loss.
-After pre-training, the DL model was fine-tuned for power transmission regression using the proposed method.
-The performance of the trained model was evaluated on test data points using various performance metrics, while the number of labeled training channel gain matrices were varied.
-Additionally, a supervised baseline performance for the regression task was obtained by training the DL model with all available labeled training channel gain matrices.
+The model is developed using the PyTorch [50] software framework.
+The dataset’s validation and test sizes are $20\%$, allowing for a more thorough evaluation of the model’s performance.
+Hyperparameter tuning yields the value of hyperparameters, which can be found in Table IV.
 IV-C4 Performance Metrics
-The performance metrics for evaluating our proposed method are Mean Absolute Error (MAE), cosine similarity, and correlation.
-MAE measures the average absolute difference between the predicted values and the actual values.
-Cosine similarity measures the similarity between two vectors in a multi-dimensional space.
-The Pearson correlation ($R$) represents the amount of variation in the output-dependent attribute that can be predicted by the input-independent variable.
-It also provides insight into how accurately a model can predict outcomes based on how much of the variation in those outcomes can be explained by the input-independent variable.
-They can be calculated as follows:
-$MAE=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|,$ (39)
 […]
-$correlation(\hat{Y},Y)=\frac{\text{cov}(\hat{Y},Y)}{\sigma_{\hat{Y}}\sigma_{Y}},$ (41)
-where $N$ is the total number of samples, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the predicted value, and $\cdot$ represents the dot product, and ||$\cdot$|| ⋅ | represents the magnitude of a vector.
-As for correlation, $\text{cov}(\hat{Y},Y)$ is the covariance between the predicted value and the true value, and $\sigma_{\hat{Y}}$ and $\sigma_{Y}$ are the standard deviations of the variables.
 V SIMULATION RESULTS
-In the following subsections, we will begin by analyzing the initial values to simulate the proposed algorithm.
-Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN.
-This comparative analysis aims to assess the accuracy and effectiveness of our proposed DL model in addressing the resource allocation problem.
+We analyzed the initial values in the following subsections to simulate the SS-VAE algorithm. Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN.
 V-A Initial points and performance of training the model
-In this section, we present numerical results for the considered resource allocation problem to evaluate the performance of our proposed model (SS-VAE) compared to state-of-the-art methods, namely ESA and DQN.
-The initial values and parameters used to run and simulate our proposed algorithm (SS-VAE) and the benchmark algorithms (ESA and DQN) are provided in Table III, Table III, and Table III.
-TABLE I: Resource Allocation Simulation Parameters
-TABLE II: DQN Training Parameters
-TABLE III: Proposed Model Training Parameters
+This section presents the numerical results for the resource allocation problem to evaluate the performance of the SS-VAE model compared to ESA and DQN.
+In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna.
+We consider a network with two slices, eMBB and URLLC services.
+These UEs are randomly distributed within a single-cell environment, mostly 200 to 500 meters from the RU.
+The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table IV, Table IV, and Table IV.
+TABLE II: Resource Allocation Simulation Parameters
 Parameter Value
@@
 Maximum power for URLLC & eMBB 30 dBm
+TABLE III: DQN Training Parameters
 Number of episodes and steps 250, 50
@@
 Replay memory size 400
+TABLE IV: SS-VAE Model Training Parameters
 Type Training Parameter Value
@@
 Train/Validation Split 0.2
-TABLE II: DQN Training Parameters
-TABLE III: Proposed Model Training Parameters
-As previously discussed, our proposed model aims to estimate the transmission power corresponding to each input channel matrix.
-These estimated power values, derived from the weights of the VAE, are then used to predict the allocation indicators ($\alpha_{u,b,s},\beta_{u,b,m,s}$).
-To streamline the training process, it is crucial to convert the dimensions of each channel matrix data into a one-dimensional vector.
-More specifically, our model takes as input a flattened $UB$-dimensional vector obtained from the channel gain matrix.
-The loss values for both the validation and training data points is depicted in Figure 4.
 Figure 4: Loss values of training and validation dataset.
-The decreasing trend of loss values illustrate that the proposed method (SS-VAE) could accurately estimate the target values (power of transmission, association indicators).
-In fact, the model can map the input channel gain matrix to the power of transmission as well as the indicators by employing the VAE’s property in learning the underlying information.
-The performance metrics for the test data, as presented in Table IV, clearly indicate that our proposed method (SS-VAE), which estimates transmission power based on the input channel gain matrix, exhibits the best generalization.
+Figure 4 shows the training and validation loss values.
+The decreasing trend of loss values indicates that the SS-VAE method could accurately estimate the target values (power of transmission, association indicators).
+The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE’s to learn the underlying information.
+The performance metrics for the test data, as given in Table V, clearly indicate that SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability.
+The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions.
 This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks.
-In addition, the pre-training method involving the contrastive loss has further enhanced our model’s capacity to learn the underlying features, ultimately reducing generalization errors.
-TABLE IV: Performance metrics of the proposed model
+Furthermore, the pre-training method involving contrastive loss has improved our model’s ability to learn the underlying features, reducing generalization errors.
+TABLE V: Performance metrics of the SS-VAE model
 Performance Metric MAE $R^{2}_{Score}$ Cosine Similarity
 Value 0.09207532 0.76795106 0.988023
-The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions.
-This suggests that a substantial proportion of the variability in the dependent variable (power of transmissions) can be accounted for by the independent variables (the channel gain matrix) integrated into the model.
-In simpler terms, the model is effective at describing and understanding the variations in the data, especially in terms of transmission power.
 V-B Comparison of the algorithms
-In the next phase, we will assess the performance of our proposed SS-VAE method by comparing it to the ESA algorithm.
-This evaluation will specifically consider the impact of the number of labeled samples used for training the supervised part of SS-VAE, as illustrated in Figure 5.
+We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 5.
+We analyze the aggregated throughput under varying numbers of labeled data points obtained by SS-VAE algorithm and the ESA.
+As depicted in Figure 5a, SS-VAE achieves a similar aggregated throughput compared to the ESA.
+Moreover, the discrepancy between SS-VAE and the ESA decreases significantly as the number of labeled data points increases.
+The difference in aggregated throughput between these methods is $2.1$ for $2000$ labeled samples, which will be decreased using more labeled samples.
 (a) Agg. throughput vs. Number of labeled data.
@@
 Figure 5: Comparison of the algorithms (SS-VAE vs. ESA).
-We analyze the aggregated throughput under varying numbers of labeled data points, as generated by our proposed SS-VAE algorithm and the ESA.
-As depicted in Figure 5a, our proposed model achieves a similar aggregated throughput when compared to the ESA.
-Moreover, the discrepancy between our model and the ESA decreases significantly as the number of labeled data points increases.
-We further evaluate the performance of the SS-VAE model and ESA in terms of binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$).
-Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled channel gain matrices (data points) increases.
-Notably, the difference between the $L1-norm$ of binary variables obtained by the SS-VAE method significantly decreases in comparison to ESA as the number of labeled data points grows.
-The results above highlight the reliable performance of our proposed method (SS-VAE) in estimating transmission powers, user association, and PRB allocation when compared to the ESA.
-In the subsequent section, we will assess its performance under different initial conditions, varying the number of slices, UEs, and thresholds for transmission power and rate.
-Since the SS-VAE exhibits lower error compared to the ESA but doesn’t perform identically in the resource allocation problem, we will also evaluate its performance alongside another benchmark algorithm, the aforementioned DQN algorithm in Section IV-B.
-DQN has demonstrated practical and valuable results in resource allocation problems within wireless networks and similar domains, enabling us to assess the robustness of our proposed algorithm.
+We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$).
+Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled samples increases.
+The difference between them reaches its minimum at 2000 samples (4.7% error) and is expected to decrease significantly as more labeled samples are used.
+These results demonstrate the reliable performance of our SS-VAE method in addressing the resource allocation problem.
+The following section will evaluate its performance under varying initial conditions, including different slice numbers, UEs, and transmission power and rate thresholds.
+Although SS-VAE shows lower error than ESA, it does not perform identically in resource allocation.
+Therefore, we will also compare it with the DQN algorithm, which has shown effective results in similar resource allocation problems within wireless networks [31, 35, 51].
 V-C Performance results
-In this section, we will assess the performance of the proposed model (SS-VAE) in the resource allocation problem under various scenarios.
-To begin, we will examine the aggregate weighted rate in relation to the number of UEs using three different algorithms: SS-VAE, ESA, and DQN, as depicted in Figure 6.
 (a) Aggr. throughput vs. number of UEs.
@@
 Figure 6: Aggregated throughput of the SS-VAE vs. Number of UEs.
-Figure 6a demonstrates that the proposed approach (SS-VAE) exhibits a performance that is comparable to the ESA.
-There are no significant differences in terms of aggregated throughput between the two.
-Additionally, SS-VAE achieves a superior aggregated throughput across all UE numbers when compared to the DQN algorithm.
-This highlights the robustness and high accuracy of our proposed method for the resource allocation problem.
-It’s worth noting that the sum of data rate increases initially as more UEs join the service for all three examined algorithms.
+Figure 6 illustrates the aggregate weighted rate across different numbers of UEs as achieved by three algorithms: SS-VAE, ESA, and DQN.
+Figure 6a demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs.
+The difference in performance between SS-VAE and ESA is consistent across all UEs.
+Moreover, SS-VAE outperforms DQN in aggregated throughput across all UE.
+In addition, the aggregated throughput increases initially as more UEs join the service for all three examined algorithms.
 However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints.
-Figure 6b depicts the aggregated throughput achieved by the SS-VAE algorithm concerning different power levels of O-RU and varying numbers of UEs.
-As per established knowledge [15], both increasing the power of O-RU and the number of UEs lead to higher aggregated throughput.
-The highest aggregated throughput is observed when the power of O-RU is at its highest.
-Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, in relation to the power values.
-This presentation aims to evaluate the performance of our proposed algorithm for these services in comparison to a benchmark algorithm.
-Figure 7 displays the aggregated throughput for the two considered service types, namely eMBB and URLLC, as determined by the SS-VAE and DQN algorithms.
-This visualization is intended to evaluate the performance of our proposed algorithm (SS-VAE) for these services in comparison to a benchmark algorithm.
+Figure 6b shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs.
+In resonation with [15], increasing the maximum power of O-RU and the number of UEs leads to higher aggregated throughput.
+The highest aggregated throughput is observed when the maximum power of O-RU is at its highest.
 (a) Aggr. throughput vs. O-RU power.
@@
 Figure 7: Aggregated throughput of SS-VAE vs. Power values.
-Figure 7a depicts the influence of O-RU’s maximum power thresholds on the aggregate weighted rate for both eMBB and URLLC services, comparing two algorithms, SS-VAE and DQN.
-The results indicate a direct correlation between increased maximum power and higher weighted rates.
-Additionally, the SS-VAE algorithm consistently outperforms the DQN algorithm, achieving higher aggregated throughput for both service types across various power values.
-Similarly, in Figure 7b, variations in total data rates are displayed concerning the maximum power for distinct cases of eMBB and URLLC services.
-The results reveal a direct increase in data rates as the maximum power of each UE in different service instances rises.
-Notably, the SS-VAE algorithm consistently outperforms the DQN algorithm, securing higher aggregated throughput for both service types across various power threshold values.
-VI CONCLUSION
-In this paper, a novel model is introduced for the O-RAN system’s downlink, aiming to coordinate multiple independent xAPPs to tackle the resource allocation problem with network slicing in the O-RAN architecture.
-This problem is addressed for two specific 5G services, namely eMBB and URLLC, with the overarching objective of optimizing the aggregate weighted rate by managing RU associations, power allocation, and PRBs.
-Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we introduced a semi-supervised learning approach as an efficient solution for UE and PRB allocation.
-Our two-step algorithm starts by reformulating the problem into an optimization problem considering the objective function and constraints.
-The second step consists of a two-part training process: firstly, supervised learning employing a VAE to estimate power transmission regression and association indicators, and then unsupervised learning with a contrastive loss approach, enhancing the model’s generalization and robustness.
-To evaluate our proposed algorithm’s performance, we conducted comparisons with benchmark algorithms ESA and DQN across various scenarios.
-The results indicate that our approach achieves a rate similar to ESA while outperforming the DQN algorithm, highlighting its effectiveness.
-We further explore different scenarios tailored to each service, eMBB and URLLC, considering their unique QoS requirements.
-Through simulations, we demonstrate the superior efficiency of our model, particularly in terms of sample efficiency, making it a compelling solution for resource allocation challenges within the O-RAN system.
-In future research, we plan to explore several areas.
-Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and investigate the impact of deployment costs associated with network function placement tackle.
-Furthermore, addressing UE mobility is crucial in practical scenarios.
-In addition, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, providing new research avenues in this dynamic field.
+Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, about the maximum power of slice and RU.
+Figure 7a and Figure 7b both illustrate the aggregate weighted throughput achieved by SS-VAE and DQN under differetn maximum power levels for RU and slices, respectively, in eMBB and URLLC services.
+Both figures demonstrate a direct correlation between increased maximum power levels and higher weighted throughput.
+Across all scenarios, the SS-VAE algorithm consistently outperforms the DQN algorithm, providing higher aggregate throughput for both service types, regardless of whether the power levels pertain to the RU or the individual slices.
+Due to their need for higher data rates, eMBB service type exhibits higher aggregate throughput than URLLC services.
+This consistent performance advantage underscores the effectiveness of the SS-VAE algorithm in optimizing resource allocation across different power settings.
+Although ESA ensures a globally optimal solution, its computational complexity makes it impractical for real-world applications.
+DQN, while powerful, faces challenges related to convergence time, exploration efficiency, and sample requirements.
+Our SS-VAE method employs the strengths of supervised and unsupervised learning to provide a scalable and efficient solution for resource allocation in O-RAN, outperforming these benchmarks in real-world scenarios.
+This contrasts to DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency.
+Additionally, DQN often requires significant training time to converge to an optimal solution, especially in environments with high variability and complex constraints.
+It may also face challenges with exploration efficiency, often needing many iterations to fully explore the solution space, resulting in suboptimal performance in dynamic conditions.
+VI COMPUTATIONAL COMPLEXITY ANALYSIS
+This section analyzes the computational costs of the ESA, DQN, and the SS-VAE method and discusses how each algorithm scales with increasing network size and complexity.
+VI-1 ESA
+The ESA evaluates all possible combinations of resource allocations to find the global optimal solution.
+If there are $|U|$ UEs, $|B|$ RUs, $M$ PRBs, and $P_{\text{levels}}$ number of discrete power levels, the total cost is given by:
+Total $\text{Total Cost}=(|U|\times|B|)\times(M+1)!\times P_{\text{levels}}$
+Where $(M+1)!$ represents the worst case of assigning UE to PRB or not, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements.
+Despite its guarantee of finding the global optimal solution, it is impractical for real-time applications and large-scale networks.
+VI-2 DQN
+The complexity of training a DQN is expressed as $O(E\cdot(T\cdot(|S|\cdot|A|+F)))$, where $E$ is the number of episodes, $T$ is the number of steps per episode, $|S|$ is the size of the state space, $|A|$ is the size of the action space, and $F$ denotes the complexity of the neural network forward pass and backpropagation.
+The complexity of making a decision using a trained DQN is $O(F)$, where $F$ is the neural network forward pass complexity.
+The total computational cost for training and inference is dominated by the training phase, which can be expressed as:
+Total $\text{Total Cost}=O(E\cdot(T\cdot(|S|\cdot|A|+F)))$
+DQN can become computationally expensive as the state and action spaces grow due to the need for extensive exploration and training to converge to an optimal policy.
+This makes DQN less scalable for large, complex networks.
+VI-3 SS-VAE
+The complexity of training a VAE is $O(E\cdot(D\cdot L+L^{2}))$, where $E$ is the number of epochs, $D$ is the dimensionality of the input data, and $L$ is the number of latent dimensions.
+The complexity of inferring resource allocations using a trained VAE is $O(D\cdot L+L^{2})$.
+In addition, the complexity of training with contrastive loss is $O(N\cdot D\cdot L)$, where $N$ is the number of samples.
+Combining both phases, the total computational cost of the SS-VAE method is:
+Total $\text{Total Cost}=O(E\cdot(D\cdot L+L^{2}))+O(N\cdot D\cdot L)$
+The SS-VAE method scales efficiently with the number of features and samples due to its reliance on learning from data.
+While the training phase is computationally intensive, it is performed offline, and the inference phase is relatively lightweight, making it suitable for real-time applications in dynamic O-RAN environments.
+VI-4 Comparison and Implications
+The ESA guarantees finding the global optimal solution, but its computational infeasibility makes it impractical for large-scale networks.
+DQN provides a scalable solution; however, it requires extensive training and help with efficiency in large, complex networks.
+The SS-VAE method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments.
+It achieves near-optimal performance with significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems.
+VII CONCLUSION
+This paper presents a novel model for the downlink of the O-RAN system, designed to coordinate multiple independent xAPPs to address the resource allocation problem with network slicing in the O-RAN architecture.
+It focuses on optimizing the aggregate weighted throughput by managing RU associations, power allocation, and PRBs to support two central 5G services, eMBB, and URLLC.
+Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we proposed a semi-supervised learning approach to solve the resource allocation problem efficiently.
+Our algorithm begins by reformulating the problem as an optimization problem.
+The training phase involves supervised learning using a VAE to estimate power transmission and association indicators, followed by unsupervised learning using a contrastive loss approach to improve the model’s generalization and robustness.
+We compared the SS-VAE method with benchmark algorithms ESA and DQN in various scenarios and discussed their computational complexity.
+The results indicate that our method performs similarly to ESA while outperforming DQN.
+We also examined scenarios for eMBB and URLLC services, considering their specific QoS requirements.
+Through simulations, we demonstrated the efficiency of our model, making it a robust candidate for resource allocation in O-RAN systems.
+Given the dynamic and complex nature of O-RAN environments, the sample efficiency of our model allows it to adapt to changing conditions without extensive retraining, ensuring high performance with fewer labeled data samples and enabling faster deployment and response times.
+The SS-VAE model provides robustness, reliability, and generalization from limited data.
+In our future research, we aim to explore several areas.
+Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and study the impact of deployment costs associated with network function placement.
+Addressing UE mobility is crucial in practical scenarios.
+Furthermore, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, paving the way for new research opportunities in this dynamic field.

### v3

@@
-quality of service 5G fifth generation
+Generative Artificial Intelligence xApp eXtended Application
+GAI
+Generative Artificial Intelligence
+xApp
+eXtended Application
 QoS
@@
 Sixth Generation
+MIMO
+Multiple-Input Multiple-Output
 ML
 Machine Learning
+AI
+Artificial Intelligence
+CSI
+Channel State Information
 MAC
@@
 Radio Access Network
+C-RAN
+Cloud Radio Access Network
+vRAN
+Virtualized Radio Access Network
+PDCP
+Packet Data Convergence Protocol
+SDAP
+Service Data Adaptation Protocol
 eMBB
@@
 Deep Learning
+RL
+Reinforcement Learning
 DRL
@@
 Mean Absolute Error
-Semi-Supervised Learning Approach for Efficient Resource Allocation with Network Slicing in O-RAN
+mMTC
+Massive Machine-Type Communications
+FL
+Federated Learning
+TL
+Transfer Learning
+NLOS
+non-line-of-sight
+RB
+Resource Block
+AMF
+Access and Mobility Management Function
+UPF
+User Plane Function
+OPEX
+Operational Expenditures
+LLM
+Large Language Model
+MARL
+Multi-Agent Reinforcement Learning
+FDRL
+Federated Deep Reinforcement Learning
+GAN
+Generative Adversarial Network
+DM
+Diffusion Model
+Generative AI for O-RAN Slicing: A Semi-Supervised Approach with VAE and Contrastive Learning
 Salar Nouri    Mojdeh Karbalaee Motalleb     Vahid Shah-Mansouri       Seyed Pooya Shariatpanahi
-$\dagger$ School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
+School of Electrical and Computer Engineering, College of Engineering, University of Tehran, Tehran, Iran
 {salar.nouri, mojdeh.karbalaee, vmansouri, p.shariatpanahi}@ut.ac.ir
 Abstract
-This paper introduces an innovative approach to the resource allocation problem, aiming to coordinate multiple independent x-applications (xAPPs) for network slicing and resource allocation in the Open Radio Access Network (O-RAN).
-Our approach maximizes the weighted throughput among user equipment (UE) and allocates physical resource blocks (PRBs).
-We prioritize two service types: enhanced Mobile Broadband and Ultra-Reliable Low-Latency Communication.
-Two xAPPs have been designed to achieve this: a power control xAPP for each UE and a PRB allocation xAPP.
-The method consists of a two-part training phase.
-The first part uses supervised learning with a Variational Autoencoder trained to regress the power transmission, UE association, and PRB allocation decisions, and the second part uses unsupervised learning with a contrastive loss approach to improve the generalization and robustness of the model.
-We evaluate the performance by comparing its results to those obtained from an exhaustive search and deep Q-network algorithms and reporting performance metrics for the regression task.
-The results demonstrate the superior efficiency of this approach in different scenarios among the service types, reaffirming its status as a more efficient and effective solution for network slicing problems compared to state-of-the-art methods.
-This innovative approach not only sets our research apart but also paves the way for exciting future advancements in resource allocation in O-RAN.
+This paper introduces a novel generative AI (GAI)-driven, unified semi-supervised learning architecture for optimizing resource allocation and network slicing in O-RAN.
+Termed Generative Semi-Supervised VAE-Contrastive Learning, our approach maximizes the weighted user equipment (UE) throughput and allocates physical resource blocks (PRBs) to enhance the quality of service for eMBB and URLLC services.
+The GAI framework utilizes a dedicated xApp for intelligent power control and PRB allocation.
+This integrated GAI model synergistically combines the generative power of a VAE with contrastive learning to achieve robustness in an end-to-end trainable system.
+It is a semi-supervised training approach that concurrently optimizes supervised regression of resource allocation decisions (i.e., power, UE association, PRB) and unsupervised contrastive objectives.
+This intrinsic fusion improves the precision of resource management and model generalization in dynamic mobile networks.
+We evaluated our GAI methodology against exhaustive search and deep Q-Network algorithms using key performance metrics.
+Results show our integrated GAI approach offers superior efficiency and effectiveness in various scenarios, presenting a compelling GAI-based solution for critical network slicing and resource management challenges in next-generation O-RAN systems.
 Index Terms:
-Resource allocation, Network slicing, O-RAN, Variational Autoencoder, Contrastive Learning
-I INTRODUCTION
-Wireless communication has evolved from a mere convenience to an absolute necessity, driven by an insatiable demand for faster and more dependable connectivity between various quality of service (QoS) requirements [1].
-The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2].
-However, the pursuit of wireless excellence does not conclude with 5G; it marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology.
-Poised on the horizon, 6G promises to redefine the boundaries of connectivity again.
-Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3].
-These advancements can potentially revolutionize our modes of communication and reshape the fabric of our interconnected world.
-Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4].
-These techniques aim to inject unparalleled intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions.
-In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal.
-As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly clear that the evolution of wireless communication is an ongoing journey, establishing a robust foundation for the next era of wireless excellence.
-Resource allocation involves carefully distributing limited network resources among numerous users, applications, or services [7].
-In 5G, the main goal is to support a wide range of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8].
-Each service requires specific resource sets, including bandwidth, data rate, latency, power consumption, complexity, and reliability.
-Resource allocation within the 5G landscape needs to be flexible to meet these diverse demands while efficiently using network resources.
-Network slicing is an innovative concept that customizes virtual networks (slices) within a unified physical infrastructure to address specific service types and use cases [9].
-In 5G, it efficiently allocates resources for eMBB and URLLC [10] services.
-This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12].
-Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a critical technology in the transition from 5G to 6G.
-This concept aligns with the architecture of O-RAN.
-O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7].
-Implementing network slicing in O-RAN optimizes resource allocation and promotes multi-vendor interoperability and innovation.
-The core of O-RAN’s architecture consists of two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control.
-These components work together, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5].
-The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP).
-Adding to O-RAN’s adaptability and innovation are xApps (cross-applications), capable of spanning multiple network domains to optimize resource allocation for various services.
-These network segments include Fronthaul, Backhaul, and Midhaul, facilitating data transmission between components.
-Overseeing resource management and orchestration tasks is the Management and Orchestration (MANO) layer, ensuring the network’s efficiency.
-Network Functions Virtualization (NFV) decouples software from hardware, enhancing scalability.
-Furthermore, the Radio Access Network Intelligent Controller (RIC) plays a pivotal role in intelligent resource control, enabling dynamic allocation to meet diverse service requirements [5].
-These components collectively define O-RAN’s transformative architecture, promoting openness, adaptability, and efficient resource utilization in modern wireless networks.
-It is worth noting that O-RAN comprises two types of RIC: near-real-time (near-RT) RIC, which manages Radio Access Network (RAN) control and optimization, and non-real-time (non-RT) RIC.
-Within the near-RT RIC, xAPPs facilitate RAN control and optimization, while inside the non-RT RIC, rAPPs support it by providing guidance and employing ML models.
-Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
-Using xAPPs leverages O-RAN network slicing to use resources efficiently and facilitate collaboration across services.
-However, this also introduces complexities in network management, as each xAPP needs to adjust the different QoS requirements of multiple services, posing a coordination challenge.
-Furthermore, cellular operators face the difficulty of establishing consistent connectivity for devices with varying QoS demands [15, 16].
-Therefore, this study addresses the resource allocation problem within O-RAN using network slicing by considering the xAPPs.
-We propose a novel resource allocation method to effectively coordinate the numerous independent xAPPs to address network slicing and resource allocation challenges.
-The use of ML algorithms in wireless networks has increased due to the growing complexity of these networks and the challenges of creating custom optimization models.
-Traditional approaches, such as convex/non-convex optimization, are struggling with the increasing complexity stemming from changes in network architecture, evolving performance requirements, and the growing number of devices.
-Current methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often overlook the ability to learn resource distributions.
-Typically, these methods involve designing an agent for a specific task and training it from scratch, which reduces exploration efficiency and extends convergence times.
-Moreover, they require many training samples, leading to prolonged exploration phases.
-The limited adaptability of traditional ML methods to new scenarios motivates us to explore an alternative approach capable of better generalization, more efficient data utilization, and resource distribution learning.
-To address these objectives, we propose a novel methodology that uses a self-supervised learning technique, specifically contrastive learning [17], in combination with a variational autoencoder (VAE) model, which is a type of generative model [18].
-These models excel in handling uncertainty by learning distributions, offering a distribution for predictions instead of a singular deterministic output.
-Moreover, their skill in handling incomplete or missing data allows them to fill in missing values based on learned patterns.
-VAEs also excel in semi-supervised learning setups, effectively using labeled and unlabeled data to improve predictive accuracy.
-Their generative abilities enable the creation of new data instances that resemble the training set, which is helpful in scenarios requiring synthetic data generation or increased resilience against noise and outliers.
-Furthermore, contrastive learning involves comparing different examples to acquire informative data representations.
-On the other hand, a VAE consists of an encoder and decoder that work together to learn the underlying probability distribution of data.
-By integrating contrastive learning with VAE, we can address the challenge of needing many training samples in ML.
-This approach can improve performance metrics, leading to faster convergence and better results while reducing the time and effort required for algorithm design and training.
-In our proposed solution, x-Apps help to implement advanced resource allocation algorithms within the O-RAN framework.
-The algorithm is deployed as an x-App on the near-real-time RIC, enabling efficient and dynamic resource allocation adjustment based on real-time network conditions and user demands.
-I-A Contribution
-This paper introduces a system model in the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC.
-It focuses on resource allocation, with a particular emphasis on network slicing.
-The primary contributions of our paper can be summarized as follows:
-  • Our research aims to address the resource allocation problem in the O-RAN architecture, explicitly focusing on optimizing weighted throughput by allocating resources such as transmission power and Physical Resource Blocks.
-We consider the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC.
-Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
-  • Our approach to optimizing resource allocation involves two main steps.
-Firstly, we simplify the problem by framing it as an optimization problem.
-Then, we introduce a novel machine learning algorithm that combines supervised learning with a VAE and unsupervised learning with contrastive loss to optimize resource allocation in O-RAN.
-It significantly improves the model’s generalization ability, making it a more practical solution for real-world applications.
-  • While explicitly guaranteeing a global optimal solution is difficult, our method aims to achieve near-optimal solutions through evaluation with multiple benchmarks and significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems.
-The method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments.
-It is not limited to O-RAN and can be applied to future multi-vendor O-RAN architectures, leveraging semi-supervised generative model principles developed by our team.
-I-B Structure of Paper
-The paper is structured as follows: Section II reviews related works.
-In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem.
-Section IV provides a detailed description of the proposed resource allocation algorithm.
-Section V presents numerical results used to evaluate the algorithm’s performance.
-Section VI compares the computational cost of employed algorithms.
-Finally, Section VII offers our conclusion.
-II Related Works
-As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements.
-This challenge becomes more prominent in 5G and 6G networks due to their increased complexity and stringent QoS demands [16].
-Recently, a growing focus has been on addressing the network slicing problem in cellular networks [16, 19, 20, 21, 22].
-Feng et al.
-[16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements.
-Another study presented a resource allocation approach based on the M/M/1 queuing system for RAN slicing [19].
-Nevertheless, practical scenarios often involve bursty traffic, overly optimistic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services.
-Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process.
-Additionally, in a study by Kazemifard et al., [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center.
-The authors conducted a theoretical analysis and employed a mathematical model to address minimum delay problems efficiently.
-Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23].
-ML algorithms, in particular, have found application in power control and resource allocation within wireless networks.
-For instance, the SAMUS framework was employed by Bektas et al.
-[24] to forecast incoming traffic, enabling an exploration of various network slicing operations and a discussion of the trade-off between the data rate of eMBB and the latency of URLLC.
+Generative AI, Variational Autoencoder, Contrastive Learning, Open Radio Access Network (O-RAN) Slicing, Resource Allocation
+1 Introduction
+Wireless communication has evolved from a convenience to a necessity, driven by the increasing demand for faster and more reliable connectivity across diverse Quality of Service (QoS) requirements [1].
+The advent of Fifth Generation (5G) marks a significant leap, enabling applications such as augmented reality and autonomous vehicles [2], while laying the foundation for Sixth Generation (6G), which aims to further redefine global connectivity.
+To meet the growing complexity of future networks, 3GPP begins integrating Artificial Intelligence (AI) into core Radio Access Network (RAN) functions such as beam management, positioning, and Channel State Information (CSI) feedback, initiating a shift towards intelligent and adaptive systems.
+As technologies like extreme Multiple-Input Multiple-Output (MIMO), deep receivers, and deep schedulers emerge, AI becomes essential to handle real-time data-driven tasks such as Physical Resource Blocks allocation and power control in both uplink and downlink [3, 4, 5].
+Network slicing and O-RAN are critical enablers in both 5G and 6G, offering the flexibility to support services like Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and Massive Machine-Type Communications (mMTC), each with distinct performance demands [6, 4].
+In 6G, network slicing will extend to advanced use cases such as holographic communication and tactile Internet, ensuring scalable and adaptable resource management [7, 8].
+At the same time, O-RAN—with its open interfaces, disaggregation of hardware and software—provides a modular, cost-effective platform for embedding AI-based control and enabling multi-vendor interoperability [9].
+These architectures are foundational for realizing sophisticated AI-driven resource allocation strategies.
+As network demands grow and QoS requirements diversify, highly adaptive intelligent solutions for dynamic scheduling, power control, and PRB management become increasingly paramount.
+Effectively integrating such advanced AI within flexible frameworks such as network slicing and O-RAN is crucial to ensure future wireless systems are not only efficient and responsive but also proactively service-aware, making sophisticated AI capabilities indispensable for unlocking the full potential of 6G [10, 11].
+The inherent complexity and dynamism of these emerging wireless environments have spurred the adoption of AI algorithms, as conventional optimization methods often lack the agility to cope with rapidly evolving network states and diverse performance criteria.
+Although learning-based paradigms, such as Deep Learning (DL), have demonstrated potential for specific RAN optimization tasks, their application to multifaceted resource allocation in O-RAN slicing often encounters significant hurdles.
+These typically include the need for extensive, task-specific agent engineering, computationally intensive training regimens that usually start from scratch, and practical limitations related to sample inefficiency and suboptimal generalization to novel network conditions or previously unseen service demands.
+Such drawbacks critically hinder the development of truly autonomous systems capable of generating robust, adaptive, and near-optimal resource management strategies in real time.
+The previously discussed limitations of conventional AI methods in achieving adaptable and data-efficient resource allocation for dynamic networks underscore the need for novel Generative Artificial Intelligence (GAI) paradigms.
+To this end, we propose an innovative GAI framework, termed Generative Semi-Supervised Variational Autoencoder (VAE)-Contrastive Learning (SS-VAE), designed explicitly for intelligent O-RAN network slicing.
+This framework is realized through a unified architecture that intrinsically integrates the distributional learning and generative capabilities of a VAE [12] with the potent representation learning power of contrastive learning principles [13].
+Trained end-to-end using a semi-supervised strategy, our SS-VAE model is designed to learn complex resource distributions and generate optimized allocation decisions with superior generalization and data efficiency.
+The synergistic design of our Generative SS-VAE framework offers distinct advantages for GAI-driven resource optimization in O-RAN.
+The VAE core excels at modeling the underlying probability distributions of network states and resource utilization, which is crucial for managing the inherent uncertainty and can be leveraged to handle incomplete data by inferring missing values from learned patterns.
+Its generative capability, central to our GAI approach, allows the model to proactively construct or refine resource allocation strategies, finely tuned to diverse QoS demands within O-RAN slices.
+This is powerfully augmented by the integrated contrastive learning mechanism, which compels the model to learn highly informative and robust latent representations from both limited labeled data and abundant unlabeled data.
+Such a focus on representation quality, integral to our semi-supervised methodology, significantly enhances sample efficiency and fortifies the model’s ability to generalize to unseen network conditions and resist noise or outliers, making it exceptionally suited for the dynamic O-RAN environment.
+Our proposed SS-VAE algorithm is designed for practical implementation within the O-RAN architecture.
+It is envisioned for deployment as an eXtended Application (xApp) operating on the near-real-time Radio Access Network Intelligent Controller (RIC).
+This strategic placement enables the SS-VAE xApp to dynamically adjust resource allocation by leveraging real-time network telemetry and user demand information, thereby facilitating efficient, adaptive, and QoS-aware management of network slices.
+1.1 Contribution
+This paper makes significant contributions to the field of intelligent resource allocation and network slicing within the O-RAN architecture, explicitly targeting the distinct demands of eMBB and URLLC services.
+We introduce and rigorously evaluate a novel GAI driven solution.
+The primary contributions of this work are as follows.
+  • We present a detailed system model and formulate the resource allocation problem in O-RAN environments, aiming to optimize weighted throughput via efficient assignment of transmission power and PRBs in the downlink.
+This formulation explicitly considers the unique QoS requirements and service priorities of eMBB and URLLC services.
+It is designed for implementation within an xApp-based architecture to enhance the flexibility of AI-driven network management.
+  • We propose, design and implement a novel GAI framework to optimize resource allocation in O-RAN.
+This framework employs a unified deep learning architecture that combines a VAE—which captures complex data patterns and can generate resource allocation solutions—with contrastive learning to learn robust and meaningful feature representations.
+The entire model is trained end-to-end using a semi-supervised approach, which improves both data efficiency and overall performance.
+This enables effective generalization across various O-RAN environment scenarios.
+This makes the framework highly adaptable and practical for real-world, dynamic O-RAN systems.
+  • We demonstrate through extensive simulations and comparative analysis against multiple benchmarks—including the exhaustive search algorithm (ESA) and deep reinforcement learning—that our proposed AI-driven approach achieves near-optimal resource allocation performance with substantially lower computational complexity than state-of-the-art methods.
+The results validate the robustness, computational efficiency, and scalability of our AI solution, affirming its suitability for real-time resource management in dynamic O-RAN systems.
+Furthermore, the principles of our semi-supervised generative model pave the way for its application in future multi-vendor O-RAN architectures and evolving service paradigms.
+1.2 Structure of the Paper
+The paper is structured as follows: Section 2 provides a brief overview of the O-RAN architecture and its components relevant to intelligent control.
+Section 3 reviews related works in resource allocation, particularly focusing on AI-based approaches in O-RAN.
+In Section 4, we introduce the system model for the resource allocation challenge in the O-RAN architecture and formally define the problem as an optimization task.
+Section 5 elaborates on our novel Generative AI-driven resource allocation framework, detailing its underlying model architecture and training methodology.
+Section 6 presents the numerical results from our performance evaluation, comparing our approach against benchmark algorithms.
+Section 7 provides an analysis of the computational complexity of the employed algorithms.
+Finally, Section 8 offers concluding remarks and outlines potential future research directions.
+For clarity, the following section lists the main acronyms used throughout the paper.
+List of Acronyms and Abbreviations
+5G
+Fifth Generation
+6G
+Sixth Generation
+AI
+Artificial Intelligence
+AMF
+Access and Mobility Management Function
+BS
+Base Station
+C-RAN
+Cloud Radio Access Network
+CP
+Control Plane
+CSI
+Channel State Information
+CU
+Central Unit
+DL
+Deep Learning
+DM
+Diffusion Model
+DNN
+Deep Neural Network
+DQN
+Deep Q-Network
+DRL
+Deep Reinforcement Learning
+DU
+Distributed Unit
+eMBB
+Enhanced Mobile Broadband
+ESA
+Exhaustive Search Algorithm
+FDRL
+Federated Deep Reinforcement Learning
+FL
+Federated Learning
+GAI
+Generative Artificial Intelligence
+GAN
+Generative Adversarial Network
+KL
+Kullback-Leibler
+LLM
+Large Language Model
+MAC
+Medium Access Control
+MAE
+Mean Absolute Error
+MANO
+Management and Orchestration
+MARL
+Multi-Agent Reinforcement Learning
+MDP
+Markov Decision Process
+MIMO
+Multiple-Input Multiple-Output
+ML
+Machine Learning
+mMTC
+Massive Machine-Type Communications
+MSE
+Mean Squared Error
+near-RT
+Near-Real-Time
+NFV
+Network Functions Virtualization
+NLOS
+Non-Line-of-Sight
+non-RT
+Non-Real-Time
+O-RAN
+Open Radio Access Network
+OPEX
+Operational Expenditures
+PDCP
+Packet Data Convergence Protocol
+PRB
+Physical Resource Block
+QoS
+Quality of Service
+RAN
+Radio Access Network
+RB
+Resource Block
+RIC
+Radio Access Network Intelligent Controller
+RL
+Reinforcement Learning
+RU
+Radio Unit
+SDAP
+Service Data Adaptation Protocol
+SINR
+Signal to Interference & Noise Ratio
+TL
+Transfer Learning
+UE
+User Equipment
+UP
+User Plane
+UPF
+User Plane Function
+URLLC
+Ultra-Reliable Low Latency Communications
+VAE
+Variational Autoencoder
+vRAN
+Virtualized Radio Access Network
+xApp
+eXtended Application
+2 Background
+Figure 1: O-RAN network slicing
+The O-RAN Alliance has created an innovative RAN framework designed to enable an open, intelligent, virtualized, and interoperable RAN, which is crucial for affordable, next-generation wireless networks [14].
+The O-RAN architecture combines Cloud Radio Access Network (C-RAN) and Virtualized Radio Access Network (vRAN) to create open, interoperable, and flexible mobile networks that support both 5G and future 6G requirements.
+The core components of O-RAN include the Open Radio Unit (O-RU), the Open Distributed Unit (O-DU), and the Open Centralized Unit (O-CU), each responsible for specific layers of the protocol stack.
+The O-RU handles radio frequency (RF) and lower physical (PHY) functions.
+The O-DU processes the medium access control (MAC), radio link control (RLC), and higher PHY layers.
+Meanwhile, the O-CU manages the Packet Data Convergence Protocol (PDCP) and Service Data Adaptation Protocol (SDAP) and is further divided into the Control Plane (CP) and User Plane (UP) for efficient signaling and traffic routing [15, 14].
+Beyond these disaggregated data plane components, key O-RAN architectural pillars include Orchestration and Automation, the O-Cloud, and notably, the RAN RIC.
+The RIC is pivotal in enabling intelligent network operations, facilitating near-real-time and non-real-time optimization and learning-based control over the RAN.
+In the subsequent section, we provide a brief overview of the key methods and characteristics employed in the O-RAN system, which improve its adaptability and efficiency.
+2.1 Network Slicing in O-RAN
+A cornerstone of 5G/6G models, network slicing facilitates the creation of customized virtual networks on demand that operate on shared infrastructure.
+This end-to-end capability is achieved by logically separating and managing resources across three key domains: the RAN, the core, and the transport network.
+In the RAN domain, slicing involves partitioning radio resources, such as PRBs, and instantiating dedicated virtual network functions for different services, including MAC/RLC in the O-DU and PDCP/SDAP in the O-CU, as illustrated in Figure 1.
+In the core network, key nodes, such as the User Plane Function (UPF) and Access and Mobility Management Function (AMF), are virtualized and isolated to create slices that accommodate specific service-level agreements.
+The transport network is sliced to create dedicated data pathways, ensuring that the performance and QoS for these diverse connections are guaranteed.
+The O-RAN framework, with its inherent intelligence and virtualization, serves as a crucial enabler for sophisticated RAN slicing, which is indispensable to perform comprehensive end-to-end network slicing [16, 17].
+2.2 Radio Intelligent Controller (RIC)
+The RIC is functionally divided into two layers: the near-real-time (near-RT) and the non-real-time (non-RT) RIC.
+The near-RT RIC, designed for control loops operating within 10 milliseconds to 1 second latency, hosts modular xApps that enable fast and responsive control over RAN functions.
+These xApps handle tasks such as interference mitigation, handover control, and dynamic resource allocation, interacting with the RAN in real time.
+They are fundamental to enabling service-level customization, particularly through the use of network slicing.
+In contrast, the non-RT RIC, typically situated within the broader Orchestration and Automation framework, manages operations with latency requirements exceeding 1 second.
+It supports functions like policy management, network analytics, and AI model training by hosting rApps.
+These applications analyze historical data, orchestrate the training of AI models, and generate strategic policies or guidance for the near-RT RIC and its xApps [18, 19].
+Together, the non-RT RIC and near-RT RIC establish a hierarchical intelligent control loop.
+Typically, AI models are trained or designed in the non-RT RIC using historical data and network policies and then deployed or used by xApps in the near-RT RIC for inference and real-time action.
+This synergy enables predictive and adaptive RAN behavior, significantly enhancing automation in dynamic network environments.
+Both xApps and rApps are designed to be modular and leverage open interfaces, which encourages third-party innovation and accelerates the deployment of AI-driven functionalities, thereby promoting multi-vendor interoperability within the O-RAN ecosystem.
+2.3 Artificial Intelligence (AI) in O-RAN
+The integration of AI is a foundational tenet of the O-RAN architecture, designed to introduce intelligence and autonomy into network operations.
+This approach is crucial for enhancing QoS through dynamic optimization and reducing Operational Expenditures (OPEX) through automated management.
+The application of AI is envisioned across a spectrum of RAN use cases, from real-time resource allocation and traffic steering to anomaly detection and cybersecurity.
+To address these challenges, the O-RAN framework is equipped to leverage a diverse range of AI methodologies.
+Supervised learning enables predictive capabilities, such as traffic forecasting, by training on labeled historical data.
+In contrast, unsupervised learning identifies latent patterns in unlabeled data for tasks such as user clustering or threat detection.
+For autonomous control, Reinforcement Learning (RL) and its deep-learning variant ( Deep Reinforcement Learning (DRL)) allow agents to derive optimal policies through direct interaction with the network environment.
+To address privacy concerns in multi-agent environments, Federated Learning (FL) facilitates collaborative model training on distributed data without centralizing sensitive information.
+Further advancing these capabilities, GAI models like VAEs can learn data distributions to generate synthetic training data or even propose new resource management solutions directly.
+These core techniques can be augmented by Large Language Models, which provides a semantic understanding of unstructured data to enhance the overall intelligence of the system’s decision-making.
+These AI functionalities are operationally hosted within the O-RAN Intelligent Controllers (RIC).
+The non-RT RIC typically manages functions with relaxed latency constraints, such as offline model training or acting as the central server in a Federated Learning setup.
+Conversely, the near-RT RIC executes tasks requiring rapid response, including model inference for sub-second control loops and the distributed training aspects of RL and FL.
+This architectural division allows O-RAN to effectively support the entire AI lifecycle, from model creation in the non-RT RIC to real-time deployment and execution in the near-RT RIC.
+3 Related Works
+As highlighted in Section 1, ensuring reliable connectivity for devices with diverse QoS requirements presents significant challenges for cellular operators, especially in the complex and demanding landscape of 5G and emerging 6G networks [20].
+As a result, network slicing has become a key focus of research in cellular networks.
+With the increasing complexity and diversity of services, methodologies for resource allocation have undergone significant evolution.
+They have shifted from traditional mathematical optimization and heuristics to more adaptive and intelligent AI-driven approaches [20, 21, 22, 23, 24].
+3.1 Conventional Optimization and Heuristic Approaches
+Resource allocation in wireless networks has been addressed through formal mathematical optimization and heuristic algorithms [20, 21, 22, 25].
+Algorithms such as Linear and Non-Linear Programming provide a rigorous framework for defining these problems, and for specific cases like convex problems, can even yield globally optimal solutions efficiently.
+When exact solutions are computationally infeasible due to the NP-hard nature of many practical scenarios involving discrete decisions, a wide range of heuristic and meta-heuristic algorithms, including greedy methods, simulated annealing, and genetic algorithms, have been employed.
+These methods aim to find good, albeit often suboptimal, solutions quickly, providing a practical alternative to complex optimization.
+Despite their utility, these conventional approaches have significant limitations in the context of dynamic, sliced O-RAN environments.
+Their primary failing is the computational complexity that makes them too slow for the near-real-time control loops required by the O-RAN RIC.
+Furthermore, these methods typically rely on static models and lack the agility to adapt to the highly dynamic nature of wireless channels, traffic loads, and user mobility.
+Their performance is tied to the accuracy of underlying network models, and heuristics can be brittle, performing poorly when conditions deviate from their design assumptions.
+Crucially, they do not learn from operational data to improve future decisions, which makes them ill-suited for the data-rich, learning-driven paradigm enabled by O-RAN’s open interfaces.
+3.2 Discriminative AI and Reinforcement Learning Paradigms
+The limitations of conventional optimization have driven research into AI-driven solutions, with DRL emerging as a significant approach for dynamic resource allocation.
+DRL combines the function approximation capabilities of Deep Neural Network (DNN) with the trial-and-error framework of RL, enabling agents to learn optimal policies through interaction with their environment.
+Various DRL algorithms, including value-based methods like Deep Q-Network (DQN) and actor-critic methods such as Soft Actor-Critic, have been utilized in wireless resource management [26].
+In complex systems like O-RAN, Multi-Agent Reinforcement Learning (MARL) allows multiple agents to learn simultaneously, optimizing local and global objectives.
+FL has also gained importance as a distributed machine learning paradigm that facilitates collaborative training without centralizing data, thus preserving privacy.
+The combination of DRL and FL, termed Federated Deep Reinforcement Learning (FDRL), allows distributed agents to enhance DRL model building.
+They offer advantages such as adaptability, reduced reliance on precise network models, the ability to manage high-dimensional state spaces, and enhanced data privacy [27].
+However, DRL faces challenges such as sample inefficiency, long training times, and poor generalization in new scenarios.
+Additionally, designing practical reward functions is complex, and the ”black box” nature of deep networks raises interpretability concerns.
+Moreover, FL struggles with non-identically distributed data and communication overhead.
+DRL effectively learns reactive policies but does not model the generative processes or distributional characteristics of optimal solutions, highlighting the need to explore generative AI approaches to understand these underlying data distributions better.
+3.3 The Rise of Generative AI in Wireless Network Optimization
+A new frontier in AI-driven network optimization is emerging with the application of GAI.
+Unlike discriminative models that learn decision boundaries or RL agents that learn reactive policies, GAI models are designed to learn the underlying probability distribution of data.
+This allows them to generate new, synthetic data instances or solutions that exhibit similar characteristics to the training data [28].
+The unique capabilities of GAI offer several advantages for complex optimization problems, such as O-RAN resource allocation, including the ability to proactively generate novel and optimized solutions, augment limited training datasets, and robustly model the inherent uncertainty of wireless environments by capturing complex data distributions [29].
+This generative capability is realized through several prominent model architectures.
+VAEs are adept at learning compressed, meaningful representations of high-dimensional network data, making them suitable for both generating solutions and performing inference [1, 49].
+Generative Adversarial Networks uses a competitive training process and has been primarily explored for data augmentation, creating realistic synthetic channel data or traffic patterns to enhance the robustness of other AI models [35, 45].
+More recently, Diffusion Models have shown promise for generating high-fidelity samples and are being investigated for imitating expert allocation policies [30].
+The choice of model often depends on the specific task, whether it is direct solution generation, data augmentation, or learning latent representations for other processes.
+3.4 Identifying the Research Gap: Positioning the Proposed SS-VAE Framework
+Despite notable advancements, current resource allocation methods in sliced O-RAN environments still face significant challenges.
+Conventional optimization and heuristic approaches often lack the speed required for real-time control and struggle to adapt to dynamic network conditions.
+DRL provides adaptability but suffers from poor sample efficiency and limited generalization, as it typically learns reactive policies without fully modeling the optimal solution distribution.
+Advanced techniques, such as FL and Transfer Learning (TL), offer some benefits but do not adequately address the need to learn from limited labeled data while exploring the vast solution space of NP-hard allocation problems.
+There is a pressing need for approaches that can learn robust representations from scarce or mixed-quality data, model complex uncertainties, and explore the solution space in a generative manner to find near-optimal resource allocation strategies.
+Specifically, this is essential for coordinating multiple xApps within dynamic O-RAN environments.
+The challenge is heightened by the limited generalizability of existing models across diverse deployment conditions, including varying numbers of User Equipment (UE), fluctuating traffic patterns, and heterogeneous network configurations.
+This limits their scalability and effectiveness in real-world scenarios.
+To address these issues, we propose the Generative Semi-Supervised VAE-Contrastive Learning (SS-VAE) framework, which integrates key AI concepts to fill this gap.
+At its core, the VAE learns the underlying probability distributions of network states and optimal resource allocations, enabling proactive generation and refinement of allocation decisions.
+This generative capability is vital for navigating the complex solution landscape of joint resource allocation, moving beyond the reactive tendencies of DRL and the data demands of purely supervised learning.
+Table I summarizes the key characteristics, contributions, and limitations of selected related works alongside the proposed SS-VAE approach.
+TABLE I: Comparative Analysis of AI-Driven Resource Allocation Methods in O-RAN
+Reference Problem Domain Methodology Key Contribution(s) Limitation(s) Primary Differentiator vs. This Work
+SS-VAE Multi-resource (power, PRB, UE assoc.) allocation for eMBB/URLLC slicing in O-RAN.
+Semi-Supervised Generative AI (VAE with Contrastive Learning).
+Novel unified GAI framework.
+Data-efficient via semi-supervision.
+Generative exploration of solution space.
+Requires initial limited labels from a high-complexity solver.
+(Baseline for comparison)
 Li et al.
-[25] introduced a method that utilizes DRL for proactive resource allocation to mobile UE, considering both time and frequency domains while accounting for mobility.
+[31] Proactive resource allocation considering UE mobility in HetNets.
+DRL.
+Adapts proactively to user mobility.
+Standard DRL challenges (e.g., sample complexity, generalization).
+Learns a reactive policy, whereas our work is generative, learning the solution distribution.
 Wu et al.
-[26] proposed a dynamic resource allocation method for radio and computation resources through RAN slicing, including a defined learning scheme with constraints.
-Further, Ayala et al.
-[27] presented a method for the joint allocation of virtualized RAN’s radio and computation resources using a DRL approach based on an actor-critic network.
+[32] Joint radio and computation resource allocation for vehicular RAN slicing.
+DRL.
+Dynamically allocates both radio and compute resources.
+Relies on reactive DRL policy; generalization across diverse slice types.
+Reactive DRL policy vs.
+our generative exploration of the joint resource space.
 Wang et al.
-[28] introduced a DRL-based self-play approach that can be integrated into RAN architecture to efficiently manage RU and DU resources, leading to cost-effective RAN operations.
-In another approach, Ndikumana et al.
-[29] combined federated learning and Deep Q-Learning with a reward function to reduce offloading, fronthaul routing, and computation delays in O-RAN.
-Similarly, Samarakoon et al.
-[30] introduced a federated learning algorithm to optimize power transmission and radio resource allocation in vehicular networks.
-Additionally, Elsayed et al.
-[31] proposed a reinforcement learning-based method for joint power and radio resource allocation in 5G networks.
-III System Model and Problem Formulation
-In our study, we address the resource allocation problem using network slicing within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice.
-The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery.
-In contrast, eMBB slices focus on high data rates for mobile broadband.
-Both slices have different QoS requirements for network resource allocation [10].
-Slicing the RAN into separate components for URLLC and eMBB optimizes the network for different traffic types and services, improving the efficient and flexible utilization of network resources.
-Key resources in O-RAN allocation include transmission power and PRBs.
-Traffic demands, often quantified by data rate and bandwidth, reflect the volume of data requiring transmission through the network.
-QoS requirements define the desired network performance based on factors such as data rate.
-The system model for our resource allocation problem with network slicing is illustrated in Figure 1.
-Table I describes all variables and parameters used in this study.
+[33] Resource management for O-RAN RU and DU entities.
+Self-Play DRL.
+Achieves adaptive decision-making through a self-play paradigm.
+Scalability and generalization of self-play in a complex O-RAN environment.
+Game-theoretic DRL policy vs.
+our GAI approach to learning underlying data characteristics.
+Ndikumana et al.
+[34] Joint task offloading and fronthaul routing in O-RAN to reduce delay.
+FL + DRL.
+Ensures data privacy via FL while optimizing routing decisions.
+High complexity in coordinating FL and DRL; potential convergence issues with non-IID data.
+Augments a reactive DRL policy with FL, while our method is a unified generative model.
+Mhatre et al.
+[35] Efficient resource management in 6G networks.
+TL + DRL.
+Reduces DRL training overhead by transferring knowledge from related tasks.
+TL is task-dependent; still relies on a reactive DRL core.
+Aids a reactive policy with TL; our method aims for broader generalization via distributional learning.
+Chen & Heydari [36] Resource governance via adaptive adjustment of network topology.
+VAE + DRL.
+Optimizes network structure using an RL agent in a VAE-compressed latent space.
+High training complexity; uses VAE for state compression, not generation.
+VAE is used for state representation for an RL agent; our method is a unified GAI for direct solution generation.
+Qiao et al.
+[37] Multi-timescale resource allocation for O-RAN slicing (radio & computing).
+Parallel Hierarchical DRL.
+Decomposes the problem hierarchically to handle complexity and coupled constraints.
+High complexity of the hierarchical DRL architecture.
+Advanced DRL structure that still learns reactive policies; our method is a single, unified generative model.
+Lotfi et al.
+[26] Dynamic O-RAN slicing and resource management.
+LLM + MARL.
+Enriches RL state with semantic context from an LLM for improved planning.
+High computational cost of LLMs; relies on the quality of LLM output.
+Augments DRL state with LLM semantics; we learn robust features directly from raw data via contrastive learning.
+Salama et al.
+[27] Resource allocation in O-RAN specifically to support FL processes.
+FL + RL + Model-based.
+Optimizes O-RAN resources (e.g., power) to improve the efficiency of an ongoing FL task.
+Solves a different, specific problem (RA for FL), not a general network slicing problem.
+Different problem domains and uses a hybrid of non-generative methods.
+Qazzaz et al.
+[38] Dynamic PRB allocation in O-RAN based on traffic and QoS.
+Supervised ML (Random Forest Classifier).
+Provides fast decisions by selecting the best policy from a predefined, finite set of options.
+Cannot generate novel allocation policies; the quality of predefined policies bounds performance.
+A discriminative model that selects a policy; our model is generative and creates new allocation solutions.
+4 System Model
+This study addresses the resource allocation problem in the O-RAN architecture through network slicing, focusing on two primary slice types: URLLC and eMBB.
+The URLLC slice is designed for ultra-low latency and high-reliability communications, which are essential for critical applications.
+In contrast, the eMBB slice aims to deliver high data rates for mobile broadband services.
+Each slice type imposes distinct QoS requirements on resource allocation, as conceptually illustrated for the O-RAN functional split in Figure 1 [6].
+Based on Figure 1, the O-RU is shared across services.
+Within the O-DU, the High PHY layer is common, while the MAC and RLC layers are logically separated per slice.
+In the O-CU-UP, the SDAP and PDCP layers are also sliced to cater to differentiated service needs.
+Logically slicing the RAN into separate URLLC and eMBB service-aware components allows for optimized network operation tailored to diverse traffic characteristics, thereby enhancing overall resource efficiency and operational flexibility.
+Key resources considered for allocation in our O-RAN model include transmission power and PRBs.
+Traffic demands are typically characterized by data rate and bandwidth requirements, while QoS parameters specify the target performance metrics for each service.
+The overall system model is depicted in Figure 2.
+All variables and parameters used throughout this paper are defined and detailed in Table II.
 Figure 2: O-RAN network system model
-In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC.
-These slices are designed to meet the QoS requirements of User Equipments.
-Our analysis focuses on two key network functions, power control, and radio resource allocation, which are jointly considered.
-We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$.
-Each BS $b$ serves two sets of UEs: $\mathcal{U}_{e,b}$ for eMBB and $\mathcal{U}_{u,b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e,b}+U_{u,b}$.
-Each UE is exclusively associated with a single RU.
-The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [32].
-Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain.
-We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices.
-The total number of slices ($S$) will be $S=S_{e}+S_{u}$.
-The UE association is indicated by a binary variable $\alpha_{u,b,s}$, which takes a value of 1 if UE $u$ is connected to RU $b$ in slice $s$, and 0 if not.
-Since each UE should be connected to only one RU, the following constraint must be met:
-$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
-We define the binary variable $\beta_{u,b,m,s}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise.
-$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
-Similarly, each PRB $m$ of RU $b$ should be associated with only one UE, as defined below:
-$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
-We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u,b,m,s}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u,b,m,s}$.
-The rate achieved for a UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u,b,m,s}$, where $\eta_{u,b,m,s}$ represents the Signal to Interference & Noise Ratio (SINR).
-The calculation for $R_{u,b,m,s}$ is determined by:
-$\eta_{u,b,m,s}=\frac{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}{\sum_{j\neq b}^{N}\sum_{l\neq s}^{S}\sum_{i\neq u}^{U}\alpha_{i,j,l}h_{i,j,m,l}p_{i,j,m,l}+\sigma^{2}},$ (4)
-$R_{u,b,m,s}=\log_{2}(1+\eta_{u,b,m,s}),$ (5)
-$R_{u,s}=\sum_{b\in B}\sum_{m\in M}\alpha_{u,b,s}\beta_{u,b,m,s}R_{u,b,m,s},$ (6)
-$R_{u}$ $=\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\alpha_{u,b,m,s}\sum_{m\in\mathcal{M}}\beta_{u,b,m,s}R_{u,b,m,s}$
-$=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}\beta_{u,b,m}R_{u,b,m,s}\;\;\;\;\forall m\in\mathcal{M},\;s\in\mathcal{S}.$ (7)
-The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P_{b}$ and $C_{b}$, respectively.
-Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values.
-This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE.
-These values are determined as follows:
-$P_{b}=\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{u,b,m,s}}+\sigma_{1}^{2},$ (8)
-$C_{b}=\log_{2}(1+\frac{\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u,b,s}h_{u,b,m,s}p_{b,m,s}}}{\sigma_{1}^{2}})$
-$=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}),$ (9)
-The channel gain matrices in our model are generated using a Rayleigh fading process to simulate variable channel conditions typical in non-line-of-sight environments.
-The channel vector from the PRB $m$ in RU $b$ to the $u$ UE in the $s$ slice, denoted by $\mathbf{h}_{u,b,m,s}$, is modeled as $\mathbf{h}_{u,b,m,s}=\sqrt{\beta_{u,b,m,s}}\mathbf{g}_{u,b,m,s}$, where $\beta_{u,b,m,s}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u,b,m,s}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable.
-To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
-The total delay experienced by each UE is the sum of propagation and transmission delays.
-The overall propagation delay ($D^{pro}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{b,p}$) links.
-In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=\frac{L}{C}$, where $L$ is the link length and $C$ is the propagation speed.
-Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{b,t}$) links.
-Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u,s}}$, where $R_{u,s}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size.
-The total delay for UE $u$ in slice $s$ is $D_{u,s}=\frac{L}{C}+\frac{\mu}{R_{u,s}}$, and should be below a specific threshold ($D^{max}_{u,s}$), expressed as $D_{u,s}\leq D^{max}_{u,s}$.
-The objective function is designed to optimize the overall network utility, which involves maximizing the weighted sum of the data rates achieved by UEs while ensuring satisfying the QoS requirements.
-In this context, $w_{s}$ indicates the priority associated with individual service types.
-The intelligent RU is responsible for effectively managing the requirements of both network slices.
-Hence, we can mathematically express the problem in the following manner:
-$\max_{\mathbf{p,\alpha,\beta}}$ $\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}w_{s}R_{u,s}$ (10)
-subject to
-$P_{b}\leq P_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (11)
-$p_{u,b,m,s}\geq 0\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (12)
-$p_{u,b,m,s}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (13)
-$R_{u,s}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
-$C_{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (15)
-$D_{u,s}\leq D^{max}_{u,s}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S},$ (16)
-$\alpha_{u,b,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (17)
-$\beta_{u,b,m,s}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (18)
-$\sum_{b\in\mathcal{B}}\alpha_{u,b,s}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}$ (19)
-$\beta_{u,b,m,s}\leq\alpha_{u,b,s}\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},u\in\mathcal{U},s\in\mathcal{S}$ (20)
-$\sum_{u\in\mathcal{U}}\alpha_{u,b,s}\beta_{u,b,m,s}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (21)
-This formulation can be seen as a well-known NP-hard problem, a multi-dimensional knapsack problem.
-Since we can reduce it to our resource allocation problem in polynomial time, it follows that it is also NP-hard.
-Given the NP-hard nature of the resource allocation problem, finding an exact solution in polynomial time is computationally infeasible for large-scale networks.
-The system model is constrained by limited resources such as UE or RU and slice power and fronthaul capacity, preventing the aggregate throughput from exceeding its optimal value.
-This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely.
-As a result, the algorithm used to solve the system will converge to an optimal solution.
-If the objective function is a strictly ascending function concerning the number of iterations, it will converge to its global optimum.
-The algorithm will converge to a local optimum if it is a non-monotonically ascending function.
-This guarantees the stability and convergence of the system model within the feasible region.
 Notation Definition
-$b,\mathcal{B},B$ Index, set, and the number of RUs
+$b,\mathcal{B},B$ Index, set, and the number of Radio Units
 $u,\mathcal{U},U$ Index, set, and the number of UEs
@@
 $h,H$ Index, and channel gain matrix between RUs and UEs
-$\alpha_{u,b,s}$ UE $u$ association to RU $b$ in slice $s$
-$\beta_{u,b,m,s}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
-$p_{u,b,m,s}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
-$R_{u,b,m,s}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
-$R_{u,s}$ Rate achieved for UE $u$ in slice $s$
+$\alpha_{u_{s}}^{b}$ UE $u$ association to RU $b$ in slice $s$
+$\beta_{u_{s},m}^{b}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
+$p_{u_{s},m}^{b}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
+$R_{u_{s},m}^{b}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
+$R_{u_{s}}$ Rate achieved for UE $u$ in slice $s$
 $R_{u}$ Rate achieved for UE $u$
 $w_{s}$ Priority weights for each service type
-$P_{b}$ Transmitted power RU in the fronthaul link
-$C_{b}$ Data rate in the fronthaul link
+$P^{b}$ Transmitted power RU in the fronthaul link
+$C^{b}$ Data rate in the fronthaul link
 $\sigma_{1}^{2}$ Quantization Noise
-$D^{pro}$ Overall propagation delay
+$D^{\text{pro}}$ Overall propagation delay
 $D^{fr,p}$ Propagation delay in the fronthaul link
 $D^{mid,p}$ Propagation delay in the midhaul link
-$D^{b,p}$ Propagation delay in the backhaul link
+$D^{bc,p}$ Propagation delay in the backhaul link
 $L$ Length of the fiber link
@@
 $D^{mid,t}$ Transmission delay in the midhaul link
-$D^{b,t}$ Transmission delay in the backhaul link
+$D^{bc,t}$ Transmission delay in the backhaul link
 $R$ Data rate of the packet
 $\mu$ Mean packet size
-TABLE I: Notations with definitions for the system model parameters.
-IV Methodology
-Our research addresses the resource allocation problem within the O-RAN architecture by introducing a semi-supervised learning algorithm.
-To achieve this, we have compiled a comprehensive dataset that includes various resource allocation scenarios, covering different objective functions and constraint combinations.
-We have created labeled data (computed $\alpha$, $\beta$, and transmission powers) using exhaustive search algorithm (ESA), thoroughly exploring the feasible solution space to find the global optimum solution as input for our algorithm’s supervised learning part and as a benchmark for models evaluation.
-We use a deep neural network (DNN) model with a contrastive loss function to solve the resource allocation problem.
-This DNN model effectively manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions.
-By training our model, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem (QoS).
-This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
-We aim to evaluate the performance of our SS-VAE method using multiple regression-based performance metrics to ensure its robustness.
-We will also validate the effectiveness of our model by comparing its results with those of the ESA across various scenarios.
-While ESA is an algorithm for resource allocation decisions that find global optimum solutions, its complexity makes it impractical for real-world deployment.
-Therefore, we will also compare its performance with state-of-the-art methods known for their practical deployment capabilities, such as DRL.
-This thorough evaluation will demonstrate the reliability of our method across different scenarios.
-Figure 2 illustrates our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
+TABLE II: Notations with definitions for the system model parameters.
+In our study, we explore a deployment based on O-RAN, wherein each Base Station (BS), known as RU in O-RAN, supports two different slices: eMBB and URLLC.
+These slices are designed to meet the QoS requirements of UEs.
+Our analysis focuses on two key network functions, power control and radio resource allocation, which are jointly considered.
+We represent the set of all RUs as $\mathcal{B}$ and the set of UEs as $\mathcal{U}$.
+Each RU $b$ serves two sets of UEs: $\mathcal{U}_{e}^{b}$ for eMBB and $\mathcal{U}_{u}^{b}$ for URLLC, leading to a total of $U$ UEs, where $U=U_{e}^{b}+U_{u}^{b}$.
+Each UE is exclusively associated with a single RU.
+The temporal and spectral layout is divided into $\mathcal{M}$ segments known as PRBs, which are the smallest unit of resources for allocation [39].
+Each PRB covers a specific time slot in the temporal domain and a frequency range of f Hz in the spectral domain.
+We denote $S_{e}$ for eMBB slices and $S_{u}$ for URLLC slices.
+The UE association is indicated by a binary variable $\alpha_{u_{s}}^{b}$ taking a value of 1 if UE $u$ in slice $s$ is connected to RU $b$ and 0 otherwise.
+Since each UE should be connected to only one RU, the following constraint must be met:
+$\sum_{b\in\mathcal{B}}\alpha_{u_{s}}^{b}=1,\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (1)
+We define the binary variable $\beta_{u_{s},m}^{b}$ as the indicator for PRB allocation, taking the value of 1 when UE $u$ in slice $s$ is connected to PRB $m$ of RU $b$, and 0 otherwise. We have
+$\beta_{u_{s},m}^{b}\leq\alpha_{u_{s}}^{b},\;\;\;\;\forall b\in\mathcal{B},\;\;m\in\mathcal{M},\;\;u\in\mathcal{U},\;\;s\in\mathcal{S}.$ (2)
+Similarly, each PRB $m$ of RU $b$ is associated with only one UE, as
+$\sum_{u\in\mathcal{U}}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}\leq 1,\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},\;\;s\in\mathcal{S}.$ (3)
+We denote the transmission power of PRB $m$ in RU $b$ in slice $s$ as $p_{u_{s},m}^{b}$ for UE $u$, and the channel gain between RU $b$ and UE $u$ in slice $s$ as $h_{u_{s},m}^{b}$.
+The rate achieved for UE $u$ in slice $s$ by allocating PRB $m$ of RU $b$ is denoted by $R_{u_{s},m}^{b}$, where $\eta_{u_{s},m}^{b}$ represents the Signal to Interference & Noise Ratio (SINR).
+The calculation for $R_{u_{s},m}^{b}$ is determined by
+$\eta_{u_{s},m}^{b}=\frac{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{u_{s},m}^{b}}{\sum_{j\neq b}^{N}\sum_{l\neq s}^{S}\sum_{i\neq u}^{U}\alpha_{i_{l}}^{j}h_{i_{l},m}^{j}p_{i_{l},m}^{j}+\sigma^{2}},$ (4)
+$R_{u_{s},m}^{b}=\log_{2}(1+\eta_{u_{s},m}^{b}),$ (5)
+$R_{u_{s}}=\sum_{b\in B}\sum_{m\in M}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}R_{u_{s},m}^{b},$ (6)
+$R_{u}$ $=\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}\alpha_{u,b,m,s}\sum_{m\in\mathcal{M}}\beta_{u_{s},m}^{b}R_{u,b,m,s}$
+$=\sum_{u\in\mathcal{U}}\sum_{m\in\mathcal{M}}\beta_{u,m}^{b}R_{u_{s},m}^{b}\;\;\;\;\forall m\in\mathcal{M},\;s\in\mathcal{S}.$ (7)
+The total power transmitted by RU $b$ and the data rate for UE in the fronthaul link are represented as $P^{b}$ and $C^{b}$, respectively.
+Quantization noise ($\sigma_{1}^{2}$) can potentially impact the signal’s fidelity during transmission over the fronthaul, mainly when the original signal displays significant variations or spans a wide range of values.
+This noise may lead to inaccuracies in the transmitted signal, thereby influencing the QoS perceived by the UE.
+These values are determined as follows:
+$P^{b}=\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{u_{s},m}^{b}}+\sigma_{1}^{2},$ (8)
+$C^{b}=\log_{2}(1+\frac{\sum_{u=1}^{U}\sum_{m=1}^{M}\sum_{s=1}^{S}{\alpha_{u_{s}}^{b}h_{u_{s},m}^{b}p_{s,m}^{s}}}{\sigma_{1}^{2}})$
+$=\log_{2}(\frac{P_{b}}{\sigma_{1}^{2}}).$ (9)
+The channel characteristics in our model are simulated using a Rayleigh fading process, which represents the variable channel conditions typical of non-line-of-sight (NLOS) environments.
+The channel vector from PRB $m$ in RU $b$ to $u$ UE in $s$ slice, denoted by $\mathbf{h}_{u_{s},m}^{b}$, is modeled as $\mathbf{h}_{u_{s},m}^{b}=\sqrt{\beta_{u_{s},m}^{b}}\mathbf{g}_{u_{s},m}^{b}$, where $\beta_{u_{s},m}^{b}$ represents the large-scale fading coefficient, accounting for path loss and shadowing, and $\mathbf{g}_{u_{s},m}^{b}\sim\mathcal{N}(0,N_{0}I_{D_{s}})$ is the fast and flat fading effects modeled as a circularly symmetric complex Gaussian random variable.
+To reflect the impact of interference and uncertainties, the channel gain matrix $\mathbf{H}$ is augmented with Gaussian noise, expressed as $\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and adjustable variance $\sigma^{2}$.
+The total delay experienced by each UE is the sum of propagation and transmission delays.
+The overall propagation delay ($D^{\text{pro}}$) consists of the delays in the fronthaul ($D^{fr,p}$), midhaul ($D^{mid,p}$), and backhaul ($D^{bc,p}$) links.
+In each of these links, the propagation delay is determined as the time it takes for a signal to traverse the distance, calculated as $D=L/C$, where $L$ is the link length and $C$ is the propagation speed.
+Similarly, the total transmission delay ($D^{tr}$) is the sum of transmission delays in the fronthaul ($D^{fr,t}$), midhaul ($D^{mid,t}$), and backhaul ($D^{bc,t}$) links.
+Within each link, the transmission delay corresponds to the time needed to transmit all packets into the transmission medium, determined by $D=\frac{\mu}{R_{u_{s}}}$, where $R_{u_{s}}$ is the data rate for UE $u$ in slice $s$, and $\mu$ is the mean packet size.
+The total delay for UE $u$ in slice $s$ is $D_{u_{s}}=L/C+\mu/R_{u_{s}}$, and should be below a specific threshold ($D^{\max}_{u_{s}}$), expressed as $D_{u_{s}}\leq D^{\max}_{u_{s}}$.
+The objective of our resource allocation scheme is to optimize the overall network utility.
+This primarily involves maximizing the weighted sum of the data rates, $R_{u_{s}}$, achieved by all UEs, while simultaneously satisfying their respective QoS requirements, particularly the delay constraints.
+The weight $w_{s}$ reflects the priority or importance associated with service type $s$ (e.g., eMBB or URLLC).
+The intelligent resource allocation function, managed by xApps at the RIC and influencing RU behavior, is responsible for effectively managing these diverse requirements.
+Hence, we can mathematically express the optimization problem as follows:
+$\max_{\mathbf{p,\alpha,\beta}}$ $\sum_{s\in\mathcal{S}}\sum_{u\in\mathcal{U}}w_{s}R_{u_{s}}$ (10)
+subject to
+$P^{b}\leq P_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (11)
+$p_{u_{s},m}^{b}\geq 0\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (12)
+$p_{u_{s},m}^{b}\leq P_{s}^{max}\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S},u\in\mathcal{U}$ (13)
+$R_{u_{s}}\geq R_{s}^{min}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (14)
+$C^{b}\leq C_{b}^{max}\;\;\;\forall b\in\mathcal{B}$ (15)
+$D_{u_{s}}\leq D^{max}_{u_{s}}\;\;\;\forall u\in\mathcal{U},s\in\mathcal{S}$ (16)
+$\alpha_{u_{s}}^{b}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},s\in\mathcal{S}$ (17)
+$\beta_{u_{s},m}^{b}\in\{0,1\}\;\;\;\forall u\in\mathcal{U},b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (18)
+$\sum_{b\in\mathcal{B}}\alpha_{u_{s}}^{b}=1\;\;\;\forall u\in\mathcal{U},\;\;s\in\mathcal{S}$ (19)
+$\beta_{u_{s},m}^{b}\leq\alpha_{u_{s}}^{b}\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},u\in\mathcal{U},s\in\mathcal{S}$ (20)
+$\sum_{u\in\mathcal{U}}\alpha_{u_{s}}^{b}\beta_{u_{s},m}^{b}\leq 1\;\;\;\;\forall b\in\mathcal{B},m\in\mathcal{M},s\in\mathcal{S}$ (21)
+Equation (10) defines the objective of maximizing the total weighted data rate across all users and services.
+Constraints (11), (12), and (13) enforce limitations on the transmit power for each RU, ensure non-negative power allocation for each UE, and restrict the received power per UE, respectively.
+Constraint (14) guarantees the QoS by ensuring that each service meets its minimum required data rate.
+Furthermore, (15) and (16) account for the fronthaul capacity limitations of each RU and the delay constraints for each UE, respectively, by their QoS requirements.
+Constraints (17) and (18) define the binary nature of the decision variables $\alpha$ and $\beta$.
+Finally, constraints (19), (20), and (21) ensure that each UE is associated with exactly one RU and that Resource Block (RB) assignments are only made if the UE is associated with the corresponding RU while preventing multiple UEs from sharing the same RB on the same RU.
+This formulation can be seen as a well-known NP-hard problem, a multi-dimensional knapsack problem.
+Since we can reduce it to our resource allocation problem in polynomial time, it follows that it is also NP-hard.
+Given the NP-hard nature of the resource allocation problem, finding an exact solution in polynomial time is computationally infeasible for large-scale networks.
+The system model is constrained by limited resources, such as UE or RU, and slice power and fronthaul capacity, which prevents the aggregate throughput from exceeding its optimal value.
+This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely.
+As a result, the algorithm used to solve the system will converge to an optimal solution.
+If the objective function is a strictly increasing function concerning the number of iterations, it will converge to its global optimum.
+The algorithm will converge to a local optimum if the function is non-monotonically ascending.
+This guarantees the stability and convergence of the system model within the feasible region.
+5 Methodology
+To address the complex resource allocation problem in the O-RAN architecture, we introduce SS-VAE, a novel GAI framework rooted in semi-supervised learning.
+This method employs a specialized DNN architecture that intrinsically integrates a VAE —leveraging its ability to learn underlying data distributions—with a contrastive loss function designed to enhance representation quality.
+The SS-VAE is designed to comprehend complex relationships between input features and optimal resource allocation decisions (e.g., parameters $\alpha$, $\beta$, and transmission powers) that meet stringent QoS requirements.
+A comprehensive dataset underpins the development and initial supervised training of SS-VAE.
+This dataset features diverse O-RAN scenarios, with labeled data representing optimal allocation parameters generated via an Exhaustive Search Algorithm (ESA), which also serves as an optimal performance benchmark.
 Figure 3: Overview of the proposed methodology.
-IV-A Exhaustive Search Algorithm
-The ESA is a fundamental approach used to find the global optimal solution to resource allocation problems by evaluating all possible configurations and selecting the best one based on a predefined objective function [33].
-The ESA is a valuable benchmark solution that provides a ground truth against comparing results from other methods.
-Its advantages include simplicity of understanding and implementation and the assurance of yielding an optimal solution within a finite time frame.
-However, its guarantee of optimality comes at the cost of high computational effort, making it suitable primarily for small or well-defined problems where the solution space can be thoroughly explored.
-Nonetheless, ESA’s main limitation lies in its computational complexity, especially when addressing resource allocation problems with high dimensions and complex constraints.
-The computational cost increases exponentially as the problem size grows, making ESA impractical for large-scale resource allocation tasks [34].
-Despite this limitation, ESA is still widely used as a benchmark for comparing different optimization algorithms and as a reference for evaluating new optimization techniques.
-To evaluate the performance, accuracy, and reliability of the SS-VAE method, we will compare its results with those obtained by ESA.
-We will use the ESA to obtain a subset of solutions for the objective variables ($\alpha$, $\beta$, $p$) in the resource allocation optimization problem with network slicing.
-This labeled dataset will be used to train the supervised learning part of our SS-VAE method.
-The detailed steps of ESA are described in Algorithm 1.
+The robustness and efficacy of the SS-VAE framework will be rigorously evaluated using multiple regression-based metrics suitable for assessing the accuracy of the generated resource allocation parameters.
+Its performance will be benchmarked against both the globally optimal solutions derived from ESA and deployment-ready, state-of-the-art methods such as DRL agents.
+This comprehensive validation aims to demonstrate the reliability of SS-VAE across various operational scenarios, particularly highlighting its capability to approach optimal performance without the prohibitive computational cost associated with ESA.
+Figure 3 illustrates our proposed methodology, which encompasses the training pipeline for SS-VAE and a comparative evaluation against benchmark algorithms.
+5.1 Exhaustive Search Algorithm (ESA)
+We employ ESA method as a basis of comparison for our algorithm.
+The ESA is a fundamental method to find the global optimal solution for resource allocation problems by exhaustively evaluating all possible configurations based on a predefined objective function [40].
+It serves as a benchmark by providing a ground truth for comparing other methods.
+While it is simple to understand and guarantees optimality within a finite time, its high computational cost limits its use to small or well-defined problems [41].
+Nevertheless, it remains a widely used benchmark for evaluating and comparing optimization algorithms.
+We use the ESA to create a labeled dataset by obtaining a subset of optimal solutions for the objective variables (i.e., $\alpha$, $\beta$, $p$) in the resource allocation optimization problem.
+This dataset is subsequently used to train the supervised learning component of our SS-VAE method.
+The detailed steps of the ESA are described in Algorithm 1.
 Algorithm 1 ESA for Resource Allocation in O-RAN
@@
 6:for each possible combination of $(\alpha,\beta,p)$ do
-7:     Calculate $R_{u,s}$ using Eq. (III)
+7:     Calculate $R_{u,s}$ using Eq. (4)
 8:     if all constraints are satisfied then
@@
 17:return $best\_solution$, $best\_value$, $\mathcal{X}$
-IV-B Deep Reinforcement learning Algorithm
-As discussed in Section II, DRL algorithms effectively address resource allocation, proving their practicality in wireless networks and enhancing key metrics [35].
-In this study, we use DRL, where an agent allocates resources based on rewards linked to the objective in Problem 10.
-Hence, we model the problem as a Markov Decision Process (MDP) and use a DRL algorithm as a benchmark for evaluating our SS-VAE method.
-In the resource allocation problem (Equation 10), an O-RAN orchestrator is responsible for determining the PRB allocation decisions and identifying the corresponding RU for each UE within each slice.
-This problem can be represented as a MDP [35, 36, 37, 38].
+5.2 Deep Reinforcement learning Algorithm
+To evaluate the proposed SS-VAE methodology, we benchmark its performance against a DRL model, a paradigm increasingly recognized for its efficacy in navigating complex resource allocation tasks within contemporary wireless networks [42].
+The resource allocation challenge, as formulated in Problem 10, is cast as a Markov Decision Process (MDP).
+This formalization allows the O-RAN orchestrator, responsible for PRB assignments and RU associations per slice, to operate as a learning agent [42, 43, 44, 45].
+The MDP is characterized by:
   1. I.
@@
 Thus, we specify the reward function for our specific scenario as follows:
-    $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u,s}+\Theta_{const.}\;C_{u,b,m,s}+\Theta_{bias},$ (22)
+    $\mathfrak{R(s,a)}=\Theta_{r}\;R_{u_{s}}+\Theta_{const.}\;C_{u_{s},m}^{b}+\Theta_{bias},$ (22)
     where, $\Theta_{r}$, $\Theta_{const.}$, and $\Theta_{bias}$ represent the respective weights assigned to the objective function (Data Rate), constraints, and the bias value.
 The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP, that maximizes the cumulative reward obtained through dynamic learning from acquired data.
-The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state.
-Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task.
-Determining the value function in such scenarios presents a significant challenge.
-Two primary strategies are employed to address MDPs: model-based and model-free approaches.
-In uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
-Our resource allocation problem involves managing high-dimensional and complex state and action spaces due to continuous variables like transmission power and numerous UEs and PRB resources.
-Traditional Q-learning methods struggle in this context because of inefficient learning with rarely encountered state-action pairs [35, 38].
-To solve the problem of finding parameters such as transmission power ($p$), $\alpha$, and $\beta$, we use the Deep Q-Network (DQN) approach, which handles complex state-action spaces by approximating the $Q$ function with a deep neural network.
-DQN benefits from experience replay and target networks, enhancing learning stability and efficiency [38].
-Experience replay stores transition in memory ($\mathbb{M}$) and samples them to train the neural network ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$), reducing correlation in the data and improving gradient estimation.
-A target network stabilizes learning by updating the $Q$-network’s parameters with independent samples [38].
-Our DQN training employs the epsilon-greedy policy for action selection.
-This means that the agent chooses the action with the highest $Q$-value with a probability of $1-\epsilon$, and with a probability of $\epsilon$, it explores and selects a random action.
-The policy function is represented as:
+The inherent complexity of wireless resource allocation, characterized by high-dimensional state and action spaces and continuous variables (e.g., transmission power), renders traditional Q-learning approaches inefficient due to the curse of dimensionality and issues with generalizing from sparsely visited state-action pairs [42, 45].
+We employ a DQN framework to surmount these challenges.
+DQN leverages a DNN to approximate the action-value function (Q-function), enabling it to manage large-scale problems effectively.
+Crucially, DQN incorporates mechanisms such as experience replay—storing transitions ($\mathfrak{S}_{t},\mathfrak{A}_{t},\mathfrak{R}_{t},\mathfrak{S}_{t+1}$) in a replay memory ($\mathbb{M}$) to break data correlations—and target networks to stabilize the learning process and improve convergence.
+To ensure a robust exploration-exploitation balance during training, the DQN agent utilizes an epsilon-greedy strategy for action selection:
 $\pi(S)=\begin{cases}\text{argmax}_{\mathfrak{a}}Q(\mathfrak{S,a})&\text{with probability }1-\epsilon\\ \text{random action}&\text{with probability }\epsilon\end{cases}$
-It is worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
-IV-C Proposed Method (SS-VAE)
-Traditional resource allocation methods in wireless communication systems, such as heuristic and meta-heuristic algorithms, often rely on predefined rules or optimization algorithms that do not adapt well to changing network conditions.
-In recent times, DL techniques have been used to improve the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems [39, 40, 41, 42].
-These approaches are practical in learning complex, non-linear relationships between input features (e.g., channel matrices) and output decisions (e.g., power and resource allocation).
-The distinctive feature of DL-based resource allocation is its capacity to manage wireless communication systems’ dynamic and complex nature effectively.
-DL-based approaches can adjust to changes in network conditions, user demands, and interference patterns, providing more flexible and efficient resource management than traditional methods.
-Traditional methods often face challenges in real-time adaptation due to predefined rules, complexity in optimization, and a lack of real-time learning.
-Our SS-VAE method inputs the channel gain matrix into a DNN.
-This network learns optimal resource allocation decisions that satisfy network constraints and maximize the weighted aggregated data rate.
-We use a flattened vector, with dimensions $UB$, derived from the channel gain matrix as the input for the machine learning model.
-Our main contribution involves integrating a VAE with contrastive loss to solve the resource allocation problem.
-By adding contrastive loss into the VAE‘s loss function, we enhance the encoder’s ability to learn more refined representations, significantly improving the accuracy of resource allocation decisions.
-The SS-VAE method allows the encoder layers of the VAE to be explicitly utilized for the resource allocation task.
-In our problem, the connection between the xApps that handle power transmissions and PRB allocation emerges from their shared foundation within the VAE framework.
-The VAE operates as a shared space where both xApps come together, enabling a collaborative exploration and consideration of the relevant features for both power transmission and PRB allocations.
-The xApps access a common latent representation of the channel gain matrices through this shared space.
-This shared understanding facilitates a cohesive analysis where the relationships between power transmissions and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the key features that impact both domains.
-This integration of information helps make more comprehensive and informed decisions about resource allocation, as it ensures that both xApps use a unified understanding of the underlying data to optimize their respective objectives.
-This coordination allows for a synchronized approach to resource allocation within the O-RAN architecture.
-IV-C1 Variational Autoencoder
-The VAE is a type of DNN generative model used in unsupervised learning tasks.
-It is designed to generate new data samples that resemble a provided set of training data.
-One notable advantage of using a VAE is its ability to adeptly capture the intrinsic distribution of the channel gain matrix (input data), which is valuable when addressing the resource allocation problem [43].
-In the VAE part of our method, the input data (flattened channel gain matrix) is represented by $x$, while a lower-dimensional data representation (latent space) is denoted as $z$.
-The generative model based on the latent variable is represented as $p_{\theta}(x)$.
-The main objective of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$.
-This likelihood can be expressed mathematically as:
+This allows the agent to exploit its current knowledge by selecting the action with the highest estimated Q-value with probability ($1-\epsilon$) or to explore the action space randomly with probability ($\epsilon$).
+The comprehensive training algorithm for the DQN is detailed, and our specific implementation for this problem is encapsulated in Algorithm 3 [45].
+5.3 Semi-Supervised VAE (SS-VAE)
+To address the resource allocation problem, we introduce our SS-VAE method in this subsection.
+The model is designed to process a flattened channel gain matrix (dimension $U\times B$) through a DNN to find optimal resource allocation strategies.
+The central contribution of our method is the integration of a VAE with a contrastive loss, which improves the learning of refined representations..
+In our framework, xApps for power transmission and PRB allocation coordinate through a shared VAE foundation.
+This VAE acts as a common space, allowing xApps to collaboratively consider relevant features by accessing a shared latent representation of channel gain matrices.
+This shared understanding facilitates a cohesive analysis and simultaneous consideration of key features impacting both domains.
+This leads to more comprehensive and informed resource allocation decisions and enables synchronized O-RAN resource management.
+5.3.1 Variational Autoencoder
+The VAE, a DNN generative model for unsupervised learning, excels at capturing the intrinsic distribution of input data like the channel gain matrix, which is valuable for the resource allocation problem [46].
+It maps input data $x$ (flattened channel gain matrix) to a lower-dimensional latent space $z$, with $p_{\theta}(x)$ as the generative model.
+The VAE aims to maximize data likelihood given model parameters $\theta$, as expressed mathematically by:
 $\log p_{\theta}(x)=\mathbb{E}{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{KL}(q_{\phi}(z|x)|p(z)),$ (23)
-where $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable based on the input data, the term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable.
-The loss function (Equation 23) can be optimized by adjusting the parameters of the generative model ($\theta$) and the encoder network represented ($\phi$).
-This optimization process involves iteratively updating these parameters and computing the gradient of the objective with respect to them [43].
-The architecture of the VAE for addressing the resource allocation problem with network slicing is illustrated Figure 3.
-The VAE comprises an encoder and a decoder.
-The encoder compresses the channel gain matrix into a lower-dimensional latent code while the decoder reconstructs the channel gain matrix from this code.
-During training, the VAE minimizes reconstruction loss, improving the accuracy of predicted transmission power.
+where $q_{\phi}(z|x)$ is the encoder to model the posterior distribution of the latent variable based on the input data, and $D_{KL}(q_{\phi}(z|x)|p(z)),$ is the Kullback-Leibler (KL) divergence this posterior distribution and the prior distribution of the latent variable.
+Optimizing the loss (Equation 23) involves adjusting generative model ($\theta$) and encoder ($\phi$) parameters [46].
+As illustrated in Figure 4, the VAE ’s encoder compresses the channel gain matrix into a lower-dimensional latent code, and the decoder reconstructs it, minimizing reconstruction loss during training.
 Figure 4: Architecture of the VAE.
-Several key advantages can be observed when using the VAE for resource allocation.
-The model can represent complex distributions, capture the non-linear relationship between channel gains and resource allocation decisions, and provide a compressed latent space helpful for visualization, clustering, and anomaly detection tasks.
-Moreover, resource allocation problems are inherently non-convex and challenging for traditional optimization methods, where the VAEs can propose a more versatile and practical solution to addressing these complexities.
+Using VAEs for resource allocation offers advantages like representing complex distributions, capturing non-linear relationships, and providing a compressed latent space useful for other tasks.
+They offer a versatile solution to inherently non-convex and challenging resource allocation problems.
 The dataset used to train the SS-VAE model consists of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples.
@@
 Specifically, we have used power transmission and allocation decisions ($p,\alpha,\beta$) obtained by ESA as labels for training our SS-VAE model.
-To train the supervised learning component of our algorithm, we employ the ADAM optimizer [44].
-It facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples.
-The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
+The supervised learning component of our algorithm uses the ADAM [47] optimizer and an Mean Squared Error (MSE) loss function for the $Q_{L}$ labeled samples, defined as follows:
 $L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||encoder(VAE(H_{i}))-\Gamma_{i}||^{2}.$ (24)
-The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u,b,s}$ and $\beta_{u,b,m,s}$).
-Our model takes a flattened vector of dimension $UB$ as input, obtained from the channel gain matrix.
-IV-C2 Contrastive Loss
-Contrastive loss is a widely-used loss function in DL for various representation learning tasks, including image or text classification.
-Its fundamental aim is to construct a representation of data points that provokes similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17].
-Mathematically, the contrastive loss can be expressed as follows:
+The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u_{s}}^{b}$ and $\beta_{u_{s},m}^{b}$) from the flattened $UB$ dimensional input channel gain matrix.
+5.3.2 Contrastive Loss
+Contrastive loss, common in DL for representation learning, aims to map similar data points closer and dissimilar ones further apart in the representation space [13].
+Mathematically, it can be expressed as:
 $L=\frac{1}{2N}\sum_{i=1}^{N}(y_{i}\cdot d^{2}(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$
 $+(1-y_{i})\cdot\max(0,m-d^{2}(\mathbf{x}_{i},\mathbf{x}_{i}^{-}))),$ (25)
-where $N$ is the number of data points, $\mathbf{x}_{i}$ is the representation of the $i^{th}$ data point, $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ are the representations of the similar and dissimilar data points respectively, $y_{i}$ is a binary label indicating whether $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$ are similar or dissimilar, and $d(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$ is the distance function between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$.
-The parameter $m$ is a margin that separates the similar and dissimilar representations.
-To apply contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices.
-This can be achieved using a channel’s statistical model that reflects the relevant properties of the communication environment.
-By introducing randomness through a Rayleigh fading model, we can produce channel gain matrices with varying coefficients in magnitude and phase, thus creating either similar or dissimilar matrices.
-To generate a channel gain matrix using this approach, we create a matrix of complex Gaussian random variables with zero mean, where the variance is influenced by the distance between the transmitter and receiver while also accounting for interference as noise [45, 46].
-The variance can be adjusted to control the level of correlation between the channels, thereby generating similar or dissimilar channel gain matrices as needed.
-We can train our model with contrastive loss function once we have created similar and dissimilar channel gain matrices.
-The loss function can be formulated as:
+where $N$ is the number of data points, $\mathbf{x}_{i}$ is the $i^{th}$ data point, $\mathbf{x}_{i}^{+}$ and $\mathbf{x}_{i}^{-}$ are similar and dissimilar data points, $y_{i}$ is a binary label indicating whether $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$ are similar or dissimilar, $d(\mathbf{x}_{i},\mathbf{x}_{i}^{+})$ is the distance function between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+}$, and $m$ is a margin that separates the similar and dissimilar data points.
+To apply contrastive loss in our resource allocation problem, we generate similar and dissimilar channel gain matrices using a channel statistical model that captures the communication environment.
+Introducing randomness via a Rayleigh fading model produces matrices with varying magnitude and phase coefficients.
+Specifically, we create a complex Gaussian random matrix with zero mean, where the variance depends on transmitter-receiver distance and accounts for interference as noise [48, 49].
+Adjusting this variance controls channel correlation, creating similar or dissimilar matrices.
+With these, we train our model using the contrastive loss function formulated as follows:
 $L_{Contrastive}(\theta)=\frac{1}{\beta_{L}}$
@@
 where $\theta$ is the set of parameters in the VAE model, $\tau$ is a temperature hyperparameter, and $L(\theta)$ is the contrastive loss function.
-$\underline{H_{i}}$ and $\bar{H_{i}}$ are similar matrices to the main channel gain matrix ($H_{i}$) and are used in training the SS-VAE model [47].
-By minimizing this loss function, we can find a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices; therefore, it is used to solve the resource allocation problem.
-The connection between the VAE and the additional contrastive loss is essential for an optimal solution.
-The initial step establishes a strong base by using supervised learning to acquire a good representation of the resource allocation problem.
-The subsequent step further improves these representations, enhancing the model’s capacity to generalize and effectively handle new, unforeseen scenarios.
-IV-C3 Training the SS-VAE model
-The training phase of SS-VAE method consists of two parts.
-First, supervised learning involves training the VAE and using the VAE’s Encoder to predict power transmission.
-The second part is unsupervised learning, which uses contrastive loss to improve the generalization and robustness of our method.
-The latent space dimension represents the size of the hidden representation learned by the VAE [48].
-A larger latent space dimension can lead to a more expressive model but also increases the risk of overfitting.
-The learning rate determines the speed at which the model learns from the data.
-The batch size determines the number of samples used for each iteration of the training process.
-The number of epochs determines the number of times the model will be trained on the data.
-$L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [49].
-The model is developed using the PyTorch [50] software framework.
-The dataset’s validation and test sizes are $20\%$, allowing for a more thorough evaluation of the model’s performance.
-Hyperparameter tuning yields the value of hyperparameters, which can be found in Table IV.
-IV-C4 Performance Metrics
+$\underline{H_{i}}$ and $\bar{H_{i}}$ are matrices similar to the main channel gain matrix $H_{i}$, used for training the SS-VAE model [50].
+Minimizing this loss finds $\theta$ that enables the model to distinguish similar and dissimilar channel gain matrices, thereby addressing the resource allocation problem.
+The connection between the VAE and the contrastive loss is key for optimal performance.
+The first step uses supervised learning to learn a good representation of the problem, and the next step refines these representations, improving the model’s generalization to new, unseen scenarios.
+5.3.3 Training the SS-VAE model
+The SS-VAE training consists of two parts: supervised learning (i.e., training the VAE using its encoder for prediction) and unsupervised learning (i.e., utilizing contrastive loss to enhance generalization and robustness).
+Key hyperparameters include the latent space dimension (balancing expressiveness and overfitting) [51], learning rate, batch size, and the number of epochs.
+$L1$ and $L2$ regularization mitigates overfitting [52].
+The model uses PyTorch [53], with a 20% validation/test split.
+Tuned hyperparameters are in Table V.
+5.3.4 Performance Metrics
 […]
-V SIMULATION RESULTS
+6 SIMULATION RESULTS
 We analyzed the initial values in the following subsections to simulate the SS-VAE algorithm. Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN.
-V-A Initial points and performance of training the model
+6.1 Simulation Configuration
 This section presents the numerical results for the resource allocation problem to evaluate the performance of the SS-VAE model compared to ESA and DQN.
-In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna.
-We consider a network with two slices, eMBB and URLLC services.
-These UEs are randomly distributed within a single-cell environment, mostly 200 to 500 meters from the RU.
-The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table IV, Table IV, and Table IV.
-TABLE II: Resource Allocation Simulation Parameters
+In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna with 10 UEs for each service.
+We consider a network with two slices/services, eMBB and URLLC services.
+These UEs are randomly distributed within a single-cell environment, mostly 10 to 500 meters from the RU.
+The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table V, Table V, and Table V.
+For traffic generation, URLLC packets follow a constant bit rate (CBR) model with small fixed payloads of 32 bytes transmitted every 1 ms, resulting in a data rate of approximately 256 kbps per URLLC user.
+In contrast, eMBB traffic is modeled using a Poisson arrival process with packets sized at 1000 bytes to emulate realistic bursty user behavior such as web browsing and file downloads.
+For the channel modeling, we adopt the standard 3GPP-compliant formulation where the channel coefficient between user $u_{s}$ and RU $b$ on PRB $m$ is given by $\mathbf{h}_{u_{s},m}^{b}=\sqrt{\beta_{u_{s},m}^{b}},\mathbf{g}_{u_{s},m}^{b}$.
+The large-scale fading component $\beta_{u_{s},m}^{b}$ accounts for both path loss and shadowing and is computed as $\beta_{u_{s},m}^{b}=10^{-(\text{PL}_{\text{dB}}+S)/10}$, where the path loss in dB is modeled using the 3GPP Urban Macro (UMa) path loss model which is formulated as $\text{PL}_{\text{dB}}=13.54+39.08\log_{10}(d)+20\log_{10}(f_{c})$, where $d$ is the 2D distance between the UE and RU in meters, and $f_{c}$ is the carrier frequency in GHz.
+For $f_{c}=2$ GHz, this simplifies to $\text{PL}_{\text{dB}}=13.54+39.08\log_{10}(d)+6.02.$ Shadow fading $S$ follows a log-normal distribution $\mathcal{N}(0,8^{2})$.
+The small-scale fading component $\mathbf{g}_{u_{s},m}^{b}$ is modeled as a Rayleigh fading coefficient, i.e., $\mathbf{g}_{u_{s},m}^{b}\sim\mathcal{CN}(0,1)$, to represent non-line-of-sight propagation conditions.
+TABLE III: Resource Allocation Simulation Parameters
 Parameter Value
@@
 Maximum transmit power of each O-RU 40 dBm
-Maximum fronthaul capacity 46 bps
-Maximum data rate for eMBB 20 bps
-Maximum data rate for URLLC 2 bps
+Maximum fronthaul capacity 46 bps/Hz
+Maximum data rate for eMBB 10 bps/Hz
+Maximum data rate for URLLC 2 bps/Hz
 Maximum power for URLLC & eMBB 30 dBm
-TABLE III: DQN Training Parameters
+TABLE IV: DQN Training Parameters
 Number of episodes and steps 250, 50
-$\gamma$, $\epsilon_{finalvalue}$, $\epsilon_{decayfactor}$ 0.9, 0.1, 0.9995
+$\gamma$, $\epsilon_{\text{final value}}$, $\epsilon_{\text{decay factor}}$ 0.9, 0.1, 0.9995
 Replay memory size 400
-TABLE IV: SS-VAE Model Training Parameters
+TABLE V: SS-VAE Model Training Parameters
 Type Training Parameter Value
@@
 Train/Validation Split 0.2
+6.2 Performance of Model Training
 Figure 5: Loss values of training and validation dataset.
+This section describes the performance of the trained model.
 Figure 5 shows the training and validation loss values.
 The decreasing trend of loss values indicates that the SS-VAE method could accurately estimate the target values (power of transmission, association indicators).
-The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE’s to learn the underlying information.
-The performance metrics for the test data, as given in Table V, clearly indicate that SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability.
+The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE to learn the underlying information.
+The performance metrics for the test data, as given in Table VI, clearly indicate that the SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability.
 The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions.
@@
 Furthermore, the pre-training method involving contrastive loss has improved our model’s ability to learn the underlying features, reducing generalization errors.
-TABLE V: Performance metrics of the SS-VAE model
-Performance Metric MAE $R^{2}_{Score}$ Cosine Similarity
+TABLE VI: Performance metrics of the SS-VAE model
+Performance Metric MAE $R^{2}_{\text{Score}}$ Cosine Similarity
 Value 0.09207532 0.76795106 0.988023
-V-B Comparison of the algorithms
-We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 5.
-We analyze the aggregated throughput under varying numbers of labeled data points obtained by SS-VAE algorithm and the ESA.
+6.3 Comparison of the Algorithms
+(a) Aggregated throughput of SS-VAE and ESA.
+(b) Association error of SS-VAE and ESA.
+(c) Aggregated throughput vs. number of UEs.
+(d) Aggregated throughput vs. number of UEs ($P_{O\text{-}RU}$).
+(e) Aggregated throughput vs. O-RU power.
+(f) Aggregated throughput vs. Power of slice.
+Figure 6: Simulation results showing various performance metrics and algorithm comparisons for SS-VAE, ESA, and DQN models.
+We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 6a.
+We analyze the aggregated throughput under varying numbers of labeled data points obtained by the SS-VAE algorithm and the ESA.
 As depicted in Figure 6a, SS-VAE achieves a similar aggregated throughput compared to the ESA.
@@
 The difference in aggregated throughput between these methods is $2.1$ for $2000$ labeled samples, which will be decreased using more labeled samples.
-(a) Agg. throughput vs. Number of labeled data.
-(b) Association Error vs. Number of labeled data.
-Figure 5: Comparison of the algorithms (SS-VAE vs. ESA).
-We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$).
-Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled samples increases.
+We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u_{s}}^{b}$) and PRB allocation ($\beta_{u_{s},m}^{b}$).
+Figure 6b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them $(|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|)/\text{\# Number of }(\alpha\&\beta)$, as the number of labeled samples increases.
 The difference between them reaches its minimum at 2000 samples (4.7% error) and is expected to decrease significantly as more labeled samples are used.
@@
 Therefore, we will also compare it with the DQN algorithm, which has shown effective results in similar resource allocation problems within wireless networks [54, 42, 55].
-V-C Performance results
-(a) Aggr. throughput vs. number of UEs.
-(b) Aggr. throughput vs. number of UEs ($P_{O-RU}$).
-Figure 6: Aggregated throughput of the SS-VAE vs. Number of UEs.
-Figure 6 illustrates the aggregate weighted rate across different numbers of UEs as achieved by three algorithms: SS-VAE, ESA, and DQN.
-Figure 6a demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs.
+6.4 Performance Results
+Figure 6c demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs.
 The difference in performance between SS-VAE and ESA is consistent across all UEs.
-Moreover, SS-VAE outperforms DQN in aggregated throughput across all UE.
+Moreover, ESA achieves higher aggregate throughput than SS-VAE, outperforming DQN in total throughput across all UEs.
 In addition, the aggregated throughput increases initially as more UEs join the service for all three examined algorithms.
 However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints.
-Figure 6b shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs.
+Figure 6d shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs.
 In resonation with [17], increasing the maximum power of O-RU and the number of UEs leads to higher aggregated throughput.
 The highest aggregated throughput is observed when the maximum power of O-RU is at its highest.
-(a) Aggr. throughput vs. O-RU power.
-(b) Aggr. throughput vs. Power of slice.
-Figure 7: Aggregated throughput of SS-VAE vs. Power values.
-Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, about the maximum power of slice and RU.
-Figure 7a and Figure 7b both illustrate the aggregate weighted throughput achieved by SS-VAE and DQN under differetn maximum power levels for RU and slices, respectively, in eMBB and URLLC services.
-Both figures demonstrate a direct correlation between increased maximum power levels and higher weighted throughput.
+Figures 6e and 6f show the aggregated throughput for two service types—eMBB and URLLC—obtained using the SS-VAE and DQN algorithms.
+Specifically, Figure 6e illustrates the weighted aggregate throughput as a function of the maximum power of the RU, while Figure 6f presents the throughput variation concerning the maximum power allocated per slice.
+Both figures compare the performance of SS-VAE and DQN across different power constraints for these services.
+They demonstrate a direct correlation between increased maximum power levels and higher weighted throughput.
 Across all scenarios, the SS-VAE algorithm consistently outperforms the DQN algorithm, providing higher aggregate throughput for both service types, regardless of whether the power levels pertain to the RU or the individual slices.
@@
 Our SS-VAE method employs the strengths of supervised and unsupervised learning to provide a scalable and efficient solution for resource allocation in O-RAN, outperforming these benchmarks in real-world scenarios.
-This contrasts to DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency.
+This contrasts with DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency.
 Additionally, DQN often requires significant training time to converge to an optimal solution, especially in environments with high variability and complex constraints.
-It may also face challenges with exploration efficiency, often needing many iterations to fully explore the solution space, resulting in suboptimal performance in dynamic conditions.
-VI COMPUTATIONAL COMPLEXITY ANALYSIS
+It may also face challenges with exploration efficiency, often requiring many iterations to fully explore the solution space, which can result in suboptimal performance under dynamic conditions.
+7 COMPUTATIONAL COMPLEXITY ANALYSIS
 This section analyzes the computational costs of the ESA, DQN, and the SS-VAE method and discusses how each algorithm scales with increasing network size and complexity.
-VI-1 ESA
+7.0.1 ESA
 The ESA evaluates all possible combinations of resource allocations to find the global optimal solution.
 If there are $|U|$ UEs, $|B|$ RUs, $M$ PRBs, and $P_{\text{levels}}$ number of discrete power levels, the total cost is given by:
-Total $\text{Total Cost}=(|U|\times|B|)\times(M+1)!\times P_{\text{levels}}$
-Where $(M+1)!$ represents the worst case of assigning UE to PRB or not, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements.
+Total $\text{Total Cost}=(|B|^{|U|})\times\prod_{b}(\sum_{s}\sum_{u}\alpha_{u_{s}}^{b}+1)^{M}\times P_{\text{levels}}$
+Where $\prod_{b}(\sum_{s}\sum_{u}\alpha_{u_{s}}^{b}+1)^{M}$ represents the worst case number of possible PRB assignments to UEs without intra-cell interference in each RU $b$, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements.
 Despite its guarantee of finding the global optimal solution, it is impractical for real-time applications and large-scale networks.
-VI-2 DQN
+7.0.2 DQN
 The complexity of training a DQN is expressed as $O(E\cdot(T\cdot(|S|\cdot|A|+F)))$, where $E$ is the number of episodes, $T$ is the number of steps per episode, $|S|$ is the size of the state space, $|A|$ is the size of the action space, and $F$ denotes the complexity of the neural network forward pass and backpropagation.
@@
 This makes DQN less scalable for large, complex networks.
-VI-3 SS-VAE
+7.0.3 SS-VAE
 The complexity of training a VAE is $O(E\cdot(D\cdot L+L^{2}))$, where $E$ is the number of epochs, $D$ is the dimensionality of the input data, and $L$ is the number of latent dimensions.
@@
 While the training phase is computationally intensive, it is performed offline, and the inference phase is relatively lightweight, making it suitable for real-time applications in dynamic O-RAN environments.
-VI-4 Comparison and Implications
+7.0.4 Comparison and Implications
 The ESA guarantees finding the global optimal solution, but its computational infeasibility makes it impractical for large-scale networks.
@@
 It achieves near-optimal performance with significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems.
-VII CONCLUSION
-This paper presents a novel model for the downlink of the O-RAN system, designed to coordinate multiple independent xAPPs to address the resource allocation problem with network slicing in the O-RAN architecture.
-It focuses on optimizing the aggregate weighted throughput by managing RU associations, power allocation, and PRBs to support two central 5G services, eMBB, and URLLC.
-Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we proposed a semi-supervised learning approach to solve the resource allocation problem efficiently.
-Our algorithm begins by reformulating the problem as an optimization problem.
-The training phase involves supervised learning using a VAE to estimate power transmission and association indicators, followed by unsupervised learning using a contrastive loss approach to improve the model’s generalization and robustness.
-We compared the SS-VAE method with benchmark algorithms ESA and DQN in various scenarios and discussed their computational complexity.
-The results indicate that our method performs similarly to ESA while outperforming DQN.
-We also examined scenarios for eMBB and URLLC services, considering their specific QoS requirements.
-Through simulations, we demonstrated the efficiency of our model, making it a robust candidate for resource allocation in O-RAN systems.
-Given the dynamic and complex nature of O-RAN environments, the sample efficiency of our model allows it to adapt to changing conditions without extensive retraining, ensuring high performance with fewer labeled data samples and enabling faster deployment and response times.
-The SS-VAE model provides robustness, reliability, and generalization from limited data.
-In our future research, we aim to explore several areas.
-Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and study the impact of deployment costs associated with network function placement.
-Addressing UE mobility is crucial in practical scenarios.
-Furthermore, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, paving the way for new research opportunities in this dynamic field.
+8 Conclusions
+This paper introduced SS-VAE, a novel GAI framework for optimizing downlink O-RAN resource allocation and network slicing.
+SS-VAE coordinates xApps to manage RU associations, power, and PRBs, maximizing weighted throughput for eMBB and URLLC services.
+To tackle the NP-hard optimization challenge, SS-VAE employs a unified semi-supervised DNN.
+This architecture synergistically integrates a VAE—whose generative capabilities are honed using ESA-derived optimal parameters to determine allocation decisions—with a contrastive loss function that enhances representation learning, generalization, and robustness using both labeled and unlabeled data.
+Evaluations against the ESA and DQN benchmarks demonstrated SS-VAE’s ability to achieve near-optimal performance, comparable to ESA and significantly outperforming DQN while incurring the reduced computational cost.
+Our simulations confirmed the efficacy of SS-VAE in managing eMBB/URLLC QoS across diverse O-RAN scenarios.
+As a semi-supervised GAI framework, SS-VAE offers notable sample efficiency, robust generalization from limited data, and adaptability to dynamic conditions, marking it as a compelling solution for practical O-RAN deployments.
+Future research will address end-to-end UE delay, investigate deployment costs, and incorporate UE mobility.
+We also aim to explore enhanced GAI techniques and other advanced AI or hybrid optimization algorithms to improve further resource allocation strategies in dynamic O-RAN systems, thereby opening new avenues for research in this evolving field.